### v1

# F-95000 Cergy-Pontoise, France

Abstract
We present an approach to computing consistent answers to analytic queries in data warehouses operating under a star schema and possibly containing missing values and inconsistent data. Our approach is based on earlier work concerning consistent query answering for standard, non-analytic queries in multi-table databases. In that work we presented polynomial algorithms for computing either the exact consistent answer to a standard, non analytic query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
We extend this approach to computing exact consistent answers of analytic queries over star schemas, provided that the selection condition in the query involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute). The main contributions of this paper are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema under the above restrictions on the selection condition, and (b) showing that, under the same restrictions the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse.
Keywords: Star schema $\bullet$ analytic query $\bullet$ inconsistent data $\bullet$ consistent query answering
1 Introduction
The efficient processing of analytic queries is an important issue in databases and has motivated considerable research work since the last three decades. The main purpose of analytic queries is to extract relevant ‘statistics’ from huge volumes of data, resulting from the integration of heterogeneous databases and stored in what is called a data warehouse inmon96 ; Kimball . For efficiency reasons, the data stored in a data warehouse is generally organized according to a non-normalized schema, called a star schema. A star schema consists of two types of relation schemas (also called  tables): a fact table and a number of dimension tables. In this context, an analytic query can be seen as an SQL Group-by query involving some aggregate function such as $min$, $max$, $count$ or $sum$ operating over attributes in the fact table called measure attributes (or simply measures). Let us see an example to illustrate the concepts of star schema and analytic query.
Example 1
Consider a company structured into branches located all over the world and selling products of various types to customers. To analyze the efficiency of the company operations, one may for instance be interested in the quantities of products sold in each branch during the past year. In order to answer efficiently such a query, knowing that the data warehouse may contain billions of sales, the data are organized according to the following star schema:
  • Fact table. This table, denoted by $F$ is meant to store all sales by the company. In our example, $F$ is defined over attributes $K_{B}$, $K_{P}$, $K_{C}$, $K_{D}$ and $Qty$, with $K_{B}K_{P}K_{C}K_{D}$ being its (primary) key, which means that $F$ must satisfy the functional dependency $K_{B}K_{P}K_{C}K_{D}\to Qty$. In other words, there can’t be two distinct sales concerning the same branch, the same product, the same customer and the same date, associated with more than one quantity.
  • Dimension tables. There are four dimension tables, one for each of the attributes $K_{B},K_{P},K_{C},K_{D}$:
    • Branch, defined over the attributes $K_{B}$, $B\_Loc$, $B\_Ctry$ standing respectively for the branch identifier, the town in which the branch is located and the country which this town belongs to. The attribute $K_{B}$ is the (primary) key of Branch, meaning that the table Branch must satisfy the functional dependencies $K_{B}\to B\_Loc$ and $K_{B}\to B\_Ctry$.
    • Prod, defined over the attributes $K_{P}$, $P\_Type$, $Price$, where $K_{P}$ is the product identifier, and $P\_Type$ and $Price$ are respectively the type and the price of a product. The attribute $K_{P}$ is the (primary) key of Prod, meaning that the table Prod must satisfy the functional dependencies $K_{P}\to P\_Type$ and $K_{B}\to Price$.
    • Cust, defined over the attributes $K_{C}$, $C\_Name$, $C\_Addr$ standing respectively for customer identifier, name and address of the customer. The attribute $K_{C}$ is the (primary) key of the table Cust, meaning that Cust must satisfy the functional dependencies $K_{C}\to C\_Name$ and $K_{C}\to C\_Addr$.
    • Date, defined over the attributes $K_{D}$, $Month$, $Year$ standing respectively for the date identifier or key, the month and the year of the date. The attribute $K_{D}$ is the (primary) key of Date, meaning that Date must satisfy the functional dependencies $K_{D}\to Month$ and $K_{D}\to Year$.
Moreover, referential constraints are generally enforced in order to ensure that any key value occurring in $F$ also occurs in the corresponding dimension table. In the case of our example these constraints are expressed through the following inclusions: $\pi_{K_{B}}(F)\subseteq\pi_{K_{B}}({\sf Branch})$, $\pi_{K_{P}}(F)\subseteq\pi_{K_{P}}({\sf Prod})$, $\pi_{K_{C}}(F)\subseteq\pi_{K_{C}}({\sf Cust})$ and $\pi_{K_{D}}(F)\subseteq\pi_{K_{D}}({\sf Date})$.
In this setting a typical analytic query is to display the total quantity of all products sold in all branches during the year $2024$. This query can be expressed in SQL as follows:
select $K_{P}$, $sum(Qty)$ from $J$ where $Year=2024$ group by $K_{P}$
Here $J$ denotes the (lossless) join of all dimension tables with the fact table $F$ (although the join can by simplified by involving only $F$, Prod and Date). $\Box$
How to efficiently evaluate analytic queries against huge volumes of data has been widely investigated and lies outside the scope of the present paper; the reader is referred to Ullman regarding standard SQL query optimization and to BellatrecheGL05 regarding more specific optimization techniques for analytic queries.
Now, most approaches to optimize the evaluation of analytic queries assume that the functional dependencies and referential constraints are satisfied by the data warehouse. However, in practice, the situation is quite different as the data warehouse may contain inconsistencies and also missing data. For instance, in the above example, a customer may appear in the data warehouse with two distinct addresses (one in $Paris$ and one in $Athens$), thus violating the functional dependency $K_{C}\to C\_Addr$; or the price of a product may be missing in the table Prod. We draw attention on that, in the case of the above query, these ‘inconsistencies’ should not affect the computation of its answer, because the query does not refer to customer addresses, nor to product prices. Notice also that, if a product identifier occurs in the fact table $F$ but not in the dimension table Prod - thus violating the referential constraint $\pi_{K_{P}}(F)\subseteq\pi_{K_{P}}({\sf Prod})$, all sales involving this product can be processed when computing the answer to the above query. This is so because, when computing the answer to this query, the only needed attribute value among all attributes of the table Cust is the $K_{C}$-value of the tuple in $F$ being processed.
A more problematic situation is if the selection condition in the query is $Year=2024$ and $C\_Addr=Paris$. This is so because among all transactions regarding customers whose address may be Paris, some concern customers whose address may violate the dependency $K_{C}\to C\_Addr$ in the table Cust. Dealing with such inconsistencies, known as the problem of computing the consistent answer to an analytic query, is not trivial, and as argued in ArenasBCHRS03 ; FuxmanFM05 , techniques used for standard non analytic queries cannot be used for analytic queries.
To cope with inconsistencies and missing values in data warehouses, our approach is based on our earlier work LauS25 dealing with consistent query answering for standard, non analytic queries in multi-table databases. In that work, we presented polynomial algorithms for computing either the exact consistent answer to a standard non analytic query or bounds of the exact answer, depending on whether the query involves or not a selection condition. In the present paper, we show that in the case of a star schema, under the restrictions that the selection condition involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute), the exact consistent answer can be effectively computed. In the following section, we summarize briefly the main results of the approach in LauS25 and then focus on analytic queries over star schemas.
Considering queries, analytic or not, whose selection condition satisfies the two restrictions mentioned above, the main contribution of this paper is showing that: (a) computing the exact consistent answer to a usual projection-selection-join query over a star schema is in time polynomial in the size of the data warehouse (in contrast to LauS25 , where consistent answers to non analytic queries are approximated when involving a selection condition), and (b) the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse (with two exceptions where only approximations are given).
The paper is organized as follows: In Section 2 we recall the main features of our previous work in LauS25 , on which the present approach is based. In Section 3 we first recall the definition of a star schema and argue that the approach in LauS25 applies in this context. In Section 4 we investigate further the concept of repairs in the context of star schemas. Section 5 deals with consistent answers to queries in the case of standard projection-selection queries as well as in the case of analytic queries. In Section 6, we propose algorithms for efficiently computing the consistent answers, or in some few cases, an approximation of the consistent answers to analytic queries. In Section 7 we compare our approach to other approaches from the literature and in Section 8 we summarize the contents of the paper and suggest research directions for future work.
2 The Approach of LauS25
2.1 The Approach at a Glance
Traditionally, to verify the consistency of a multi-table database with respect to a set $FD$ of functional dependencies one applies the well-known Chase algorithm Ullman . The input of this algorithm is a table $T$ over the set $U$ of all attributes appearing in the database. $T$ has as many rows as there are tuples in the database and each tuple is placed on a separate row, eventually with missing values. The algorithm derives new tuples by applying the dependencies of $FD$ as long as no pair of tuples is in conflict with some dependency; and stops as soon as such a conflict is encountered. Let $Chase(T)$ denote the result upon termination.
We recall here that a dependency application, also known as the Lossless-Join rule is defined as follows FaginMU82 ; Ullman :
for all $t$ and $t^{\prime}$ in the current value of $Chase(T)$
if there exists $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are defined over $XA$ and $t.X=t^{\prime}.X$ then
if $t.A$ and $t^{\prime}.A$ are distinct domain values, then fail
else if $t.A=a$ and $t^{\prime}.A$ is null then assign $a$ to $t^{\prime}.A$
In this context, a tuple $t$ in the current value of $Chase(T)$ is said to be conflicting if the following holds: there is a tuple $t^{\prime}$ in the current value of $Chase(T)$ and a dependency $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are both defined over $XA$, $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$. A tuple $t$ is called non-conflicting if $t$ is not a conflicting tuple.
Now, if Chase is successful (i.e., no conflicts are encountered and no more new tuples can be derived) then the database is declared consistent else conflicting. If the database is consistent then processing of queries (whether standard queries or analytic queries) proceeds as usual, else the following question arises: can we still extract useful (i.e. non conflicting) information from the conflicting database? The work in LauS25 gives a positive answer to this question based on an extended version of the Chase algorithm, called the $m\_Chase$ algorithm to be presented formally in the following subsection. The input of this algorithm is no more $T$ but a table $\mathcal{T}$ containing all tuples that can be built up from constants in the active domains of the database. The set $\mathcal{T}$ is called the ‘Universe of discourse’ of the $m\_Chase$ algorithm and the notion of ‘conflicting tuple’ remains the same but now concerns all tuples of $\mathcal{T}$ and not only those of $T$.
It is shown in LauS25 that the tuples of $\mathcal{T}$ can be characterized in two orthogonal ways: a tuple of $\mathcal{T}$ can be either true or false, and it can be either conflicting or non-conflicting. This characterization can be intuitively described as follows. If Chase terminates successfully and $Chase(T)$ denotes the output table then a tuple $t$ of $\mathcal{T}$ is true if it appears in $Chase(T)$ and false otherwise (i.e. $t$ is false if it appears in $\mathcal{T}\setminus Chase(T)$).
However, if the Chase algorithm fails then we don’t know which tuples are true and which are non-conflicting. The $m\_Chase$ algorithm remedies this deficiency by modifying the Chase algorithm as follows: instead of stopping the application of functional dependencies on table $T$ when a conflict is encountered, the application continues (and the true tuples are stored) until no more tuples are found. In doing so, all true tuples and all conflicting tuples are computed - and therefore each tuple of $\mathcal{T}$ can be characterized as true/false and as conflicting/non-conflicting.
It follows from the above definition of conflicting tuple that if $t$ is conflicting then every true super-tuple of $t$ is also conflicting. Therefore the conflicting tuples can be retrieved as true super-tuples of true tuples of the form $xa$ over $XA$ such that: $(a)$ $X\to A$ is a dependency in $FD$ and $(b)$ $a$ and $a^{\prime}$ are in $adom(A)$ such that $a\neq a^{\prime}$ and $xa^{\prime}$ is true (here $adom(A)$ stands for ‘active domain’ of $A$). Then assuming that all true tuples and all conflicting tuples are known, we can define a tuple $t$ of $\mathcal{T}$ to be consistent if $t$ is true and non-conflicting. Note that every sub-tuple of a true tuple is true and that every sub-tuple of a consistent tuple is consistent. Finally, call ‘consistent subset’ of $\mathcal{T}$ any set $S$ of true tuples of $\mathcal{T}$ such that the set of all tuples inferred from $S$ using the functional dependencies contains no conflicting tuples in $\mathcal{S}$. Let us illustrate these concepts using the following example.
Example 2
Suppose $T=\{ab,bc,ac^{\prime}\}$ and $FD=\{A\to C,B\to C\}$. Then all tuples in $T$ are true and non conflicting (hence consistent), but the application of the functional dependencies on $T$ allows to infer the tuples $abc$ and $abc^{\prime}$, which are conflicting tuples of $\mathcal{T}$ inferred from $T$. In fact it can be seen that $(a)$ the true tuples of $\mathcal{T}$ are $abc$, $abc^{\prime}$ and all their sub-tuples, implying that any other tuple of $\mathcal{T}$ is false in $\mathcal{T}$, and $(b)$ the conflicting tuples in $\mathcal{T}$ are $abc$, $abc^{\prime}$, $ac$, $ac^{\prime}$, $bc$ and $bc^{\prime}$, implying that any other tuple in $\mathcal{T}$ is non conflicting in $\mathcal{T}$. In this example, the consistent tuples of $\mathcal{T}$ are $ab$, $a$, $b$, $c$ and $c^{\prime}$.
In this context, the set $R=\{abc,c^{\prime}\}$ is a consistent subset of $\mathcal{T}$. Indeed, since $\mathcal{R}$ is the set of all tuples built up using constants $a$, $b$, $c$ and $c^{\prime}$, we have $\mathcal{R}=\mathcal{T}$. Moreover the tuples in $R$ are tuples true in $\mathcal{T}$ and no conflicting tuples in $\mathcal{R}$ can be generated from the tuples in $R$. It is important to notice that, although $abc$ is conflicting in $\mathcal{T}$, $abc$ is not conflicting in $\mathcal{R}$ because $ac^{\prime}$, $bc^{\prime}$ and $abc^{\prime}$ are not true in $\mathcal{R}$. $\Box$
Finally, based on the concepts introduced so far, a repair of $T$ in LauS25 is defined to be a maximal and consistent subset of $\mathcal{T}$ containing a maximal and consistent set of tuples which are consistent in $\mathcal{T}$. In our previous example, the subset $R=\{abc,c^{\prime}\}$ of $\mathcal{T}$ is a repair of $T$ because $(a)$ as we have just seen, $R$ is a consistent subset of $\mathcal{T}$, $(b)$ $R$ is maximal because adding to $R$ a tuple true of $\mathcal{T}$ either does not bring any new true tuple in $\mathcal{R}$ (e.g., adding the tuple $ac$) or generates a conflicting tuple in $\mathcal{R}$ (e.g., adding the tuple $ac^{\prime}$), and $(c)$ all consistent tuples of $\mathcal{T}$ are true in $\mathcal{R}$. Note that similar arguments show that the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of $\mathcal{T}$, however, $S$ is not a repair of $T$, because $ab$ is a consistent tuple of $\mathcal{T}$ which is not true in $\mathcal{S}$. By the way, as we shall see in Section 4.1, our definition of repair is more restrictive than the usual definition ArenasBC99 ; Wijsen19 in which a repair is defined to be a maximal and consistent subset of $\mathcal{T}$. For example, the set $S=\{bc,ac^{\prime}\}$ is a repair of $T$ following ArenasBC99 ; Wijsen19 , but it is not a repair of $T$ following our approach.
2.2 The m-Chase Algorithm
Clearly to apply the $m\_Chase$-based approach described above, one has to answer the following questions:
  • Does the $m\_Chase$ algorithm terminate?
  • Is the result independent of the order in which the functional dependencies are applied?
  • Does the result contain all true tuples and all conflicting tuples that the dependencies can derive? In other words: which underlying semantics ensure that $m\_Chase$ algorithm is correct?
All these questions find positive answers in LauS25 , based on the set theoretic semantics introduced in CKS86 ; Spyratos87 , under the assumption that the set $FD$ is normalized. Following LauS25 , if $FD^{+}$ denotes the closure of $FD$ under the Armstrong’s axioms Armstrong , then $FD$ is said to be normalized if it contains all dependencies in $FD^{+}$ such that:
FD1:
every dependency in $FD$ is of the form $X\to A$ where $A$ is an attribute in $U$ not in $X$
FD2:
for every $X\to A$ in $FD$, there is no $Y\subset X$ such that $Y\to A$ is implied by $FD$ (i.e., such that $Y\to A$ is in $FD^{+}$)
As shown in LauS25 , every set $FD$ of functional dependencies can be put in an equivalent normalized form. Moreover, a set $FD$ of functional dependencies is said to be cyclic if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A$ is in $Y$ and $B$ in $X$. It is shown in LauS25 that cyclic sets of functional dependencies raise important difficulties when it comes to computing consistent answers. It is easy to see that the sets $FD$ considered in Example 1 and in Example 2 are both normalized and acyclic.
In this section, we recall from LauS25 the basic formalism on which the algorithm $m\_Chase$ relies, namely that of multi-valued tuple. A multi-valued tuple, or m-tuple, extends the notion of tuple in the sense that an m-tuple associates every attribute $A$ with a possibly empty subset of the active domain of $A$ as opposed to a single value from the active domain.
Definition 1
A multi-valued tuple $\sigma$ over universe $U$, or m-tuple, is a function from $U$ to the cross product ${\Large{\sf X}}_{A\in U}\mathcal{P}(adom(A))$, where $\mathcal{P}(adom(A))$ is the power set of $adom(A)$. The set of all attributes $A$ such that $\sigma(A)\neq\emptyset$, is called the schema of $\sigma$, denoted by $sch(\sigma)$. Given $\sigma$ and a subset $X$ of $sch(\sigma)$, the restriction of $\sigma$ to $X$, denoted $\sigma(X)$, is the m-tuple defined by $(\sigma(X))(A)=\sigma(A)$ for every $A$ in $X$ and $(\sigma(X))(A)=\emptyset$ for any $A$ not in $X$.
Given an m-tuple $\sigma$, the set ${\sf tuples}(\sigma)$ denotes the set of all tuples $t$ such that $sch(t)=sch(\sigma)$ and for every $A$ in $sch(t)$, $t.A$ belongs to $\sigma(A)$. $\Box$
[…]
We call m-table over $U$ any finite set of m-tuples over $U$. For all $\sigma$ and $\sigma^{\prime}$ in an m-table $\Sigma$, and $X\to A$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma^{\prime})$, the following rule called m-Chase rule generalizes the chase rule.
$\bullet$ m-Chase rule: Let $\sigma_{1}=\sigma\cup\sigma^{\prime}(A)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}\cup\sigma(A)$
Case of $\sigma_{1}\sqsubseteq\sigma^{\prime}_{1}$: replace $\sigma$ with $\sigma^{\prime}_{1}$, and remove $\sigma_{1}$
Case of $\sigma^{\prime}_{1}\sqsubseteq\sigma_{1}$: replace $\sigma^{\prime}$ with $\sigma_{1}$, and remove $\sigma^{\prime}_{1}$
Otherwise: replace $\sigma$ and $\sigma^{\prime}$ with $\sigma_{1}$ and $\sigma^{\prime}_{1}$, respectively.
As shown in Algorithm 1, our algorithm consists in applying the above m-Chase rule whenever ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ until no further transformation is possible. The output is an m-table denoted by $m\_Chase(T)$. It has been shown in LauS25 that this algorithm always terminates and that the partition semantics of tuples in $\mathcal{T}$ (as introduced in Spyratos87 and extended in JIIS ; LauS25 ), can be defined based on $m\_Chase(T)$ as follows:
Algorithm 1 The m-Chase Algorithm
0:  A table $T$ over $U$ and a normalized set $FD$ of functional dependencies over $U$.
0:  An m-table denoted by $m\_Chase(T)$.
1:  $\Sigma:=\{\sigma_{t}~{}|~{}t\in T\}$ // $\sigma_{t}$ is the m-tuple such that $\sigma_{t}(A)=\{t.A\}$ for $A\in sch(t)$
2:  $change:=true$
3:  while $change=true$ do
4:     $change:=false$
5:     for all $\sigma$ and $\sigma^{\prime}$ in $\Sigma$ do
6:        for all $X\to A$ in $FD$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma)$ do
7:           if ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ then
8:              apply the m-Chase rule to $\sigma$ and $\sigma^{\prime}$
9:              $change:=true$
10:  $m\_Chase(T):=\Sigma$
11:  return  $m\_Chase(T)$
Proposition 1
Let $T$ be a table over universe $U$ with $FD$ as set of functional dependencies. The following holds:
  1. 1.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf True}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ (i.e., $\sigma$ has nonempty components over attributes on which $t$ is defined),
    -- $t\sqsubseteq\sigma$.
  2. 2.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf Confl}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    -- there exists $X\to A$ in $FD$ such that $XA\subseteq sch(t)$ and $|{\sf tuples}(\sigma(A)|>1$.
  3. 3.
    A tuple $t$ in $\mathcal{T}$ is in ${\sf Cons}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
    -- for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $|{\sf tuples}(\sigma(A))|=1$.
  4. 4.
    For every $\sigma$ in $m\_Chase(T)$ and every $S\subseteq sch(\sigma)$, either ${\sf tuples}(\sigma(S))\subseteq{\sf Cons}(\mathcal{T})$ or ${\sf tuples}(\sigma(S))\subseteq{\sf Confl}(\mathcal{T})$. $\Box$
As shown in LauS25 , the computation of $m\_Chase(T)$ is in $\mathcal{O}(|m\_Chase(T)|^{3}.\delta^{2})$, where $\delta$ is the maximal cardinality of the components of m-tuples in $m\_Chase(T)$, which is precisely the maximum number of $A$-values associated with $X$-values when $X\to A$ is a functional dependency in $FD$. As Algorithm 1 shows that $|m\_Chase(T)|\leq|T|$, we state that the computation of $m\_Chase(T)$ is in $\mathcal{O}(|T|^{3}.\delta^{2})$, i.e., polynomial in the size of $T$.
To illustrate Algorithm 1 and Proposition 1, consider again the context of Example 2 where $U=\{A,B,C\}$, $FD=\{A\to C,B\to C\}$ and $T=\{ab,bc,ac^{\prime}\}$. Running Algorithm 1 yields the following steps:
  • The algorithm starts with the m-table $\Sigma=\{(a)(b),(b)(c),(a)(c^{\prime})\}$
  • Applying $B\to C$ to the first two m-tuples, we obtain $\Sigma=\{(a)(b)(c),(a)(c^{\prime})\}$
  • Applying now $A\to C$ to these two m-tuples, we obtain $\Sigma=\{(a)(b)(cc^{\prime})\}$.
Since no new m-tuple can be generated from $\Sigma$, $m\_Chase(T)=\{(a)(b)(cc^{\prime})\}$ is returned by Algorithm 1, and so, by Proposition 1, it follows that
  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of tuples in ${\sf tuples}((a)(b)(c,c^{\prime}))$, that is ${\sf True}(\mathcal{T})=\{abc,abc^{\prime},ab,$ $ac,ac^{\prime},bc,bc^{\prime},a,b,c,c^{\prime}\}$. In other words, there are no false tuples in this example.
  • ${\sf Confl}(\mathcal{T})=\{abc,abc^{\prime},ac,ac^{\prime},bc,bc^{\prime}\}$.
  • ${\sf Cons}(\mathcal{T})={\sf True}(\mathcal{T})\setminus{\sf Confl}(\mathcal{T})$, that is ${\sf Cons}(\mathcal{T})=\{ab,a,b,c,c^{\prime}\}$.
In the following section, we first define the notion of star schema and then we show that the results from LauS25 that have just been recalled apply in this context as well.
3 Star Schemas
3.1 The Context of our Approach
We first recall from the literature Ullman that a star schema, as considered in our approach, consists of the following tables and constraints:
  • $n$ dimension tables $D_{1},\ldots,D_{n}$. For $i=1,\ldots,n$, $D_{i}$ is defined over attributes $K_{i},A_{i}^{1},\ldots,A_{i}^{d_{i}}$. For $i=1,\ldots,n$, the schema of $D_{i}$ is denoted by $sch(D_{i})$, and the set $sch(D_{i})\setminus\{K_{i}\}$ is denoted by $sch^{*}(D_{i})$.
  • a fact table $F$ defined over $K_{1},\ldots,K_{n},M_{1},\ldots,M_{p}$. The attributes $M_{1},\ldots,M_{p}$ are called measures, and we denote by $\mathbf{M}$ the set of all measures, that is $\mathbf{M}=\{M_{1},\ldots,M_{p}\}$.
  • $FD=\bigcup_{i=1}^{i=n}\{K_{i}\to A_{i}^{j}~{}|~{}j=1,\ldots,d_{i}\}\cup\{K_{1}\ldots K_{n}\to M_{k}~{}|~{}k=1,\ldots,p\}$.
    In other words, for $i=1,\ldots,n$, $K_{i}$ is the key of $D_{i}$ and $K_{1}\ldots K_{n}$ is the key of $F$. We denote by $\mathbf{K}$ the set of all dimension keys that is $\mathbf{K}=\{K_{1},\ldots,K_{n}\}$.
It is easy to see that if $FD$ is defined as above, then for every non trivial functional dependency $X\to A$ in $FD^{+}$ we have: $X\cap\mathbf{K}\neq\emptyset$. More precisely, if $A$ is in $\mathbf{K}$ then $A$ must occur in $X$, in which case $X\to A$ is trivial (because $FD$ contains no dependency whose right hand-side is in $\mathbf{K}$), if $A$ is in $sch^{*}(D_{i})$ then $X$ must contain $K_{i}$ and if $A$ is in $\mathbf{M}$ then $X$ must contain $\mathbf{K}$. Thus, for every non trivial functional dependency $X\to A$ in $FD^{+}$, there exists $X_{0}\to A$ in $FD$ such that $X_{0}\subseteq X$.
Since the left hand-sides of the dependencies in $FD$ can not be reduced further, this means that $FD$ is normalized. On the other hand as the left hand-sides of functional dependencies in $FD$ are attributes in $\mathbf{K}$, not occurring in the right hand-sides of these dependencies, $FD$ is acyclic. As a consequence, all results in LauS25 apply in the context of star schemas.
In what follows, we call data warehouse a database whose schema is a star schema. Moreover, we use the terms ‘data warehouse’ and ‘table’ instead of ‘multi-relation database’ and ‘relation’, to better fit the usual terminology when dealing with data warehouses. In our approach, it is thus possible to deal with data warehouses in which some of the tables $D_{i}$ or $F$ have missing values for some of their attributes. However, in order to consider cases that make sense in practice, we restrict missing values in the warehouse tables as follows:
  1. 1.
    For every $i=1,\ldots,n$, every $t$ in $D_{i}$ is defined over the key attribute $K_{i}$ and over at least one non-key attribute in $sch^{*}(D_{i})$. We consider that storing a key value with no associated non-key value makes no sense.
  2. 2.
    For every $t$ in $F$, $t$ is defined over $\mathbf{K}$ and over at least one measure attribute in $\mathbf{M}$. We consider that storing a fact with no associated measure value makes no sense.
Example 3
We illustrate the above concepts using a toy example that will serve as a running example in the remainder of this paper. We consider two dimensions $D_{1}$ and $D_{2}$ such that $sch(D_{1})=K_{1}A_{1}^{1}A_{1}^{2}$ and $sch(D_{2})=K_{1}A_{2}^{1}A_{2}^{2}$. Moreover the fact table $F$ is such that $sch(F)=K_{1}K_{2}M_{1}$ meaning that we consider one measure attribute $M_{1}$. As specified above, we have $FD=\{K_{1}\to A_{1}^{1},K_{1}\to A_{1}^{2},K_{2}\to A_{2}^{1},K_{2}\to A_{2}^{2},K_{1}K_{2}\to M_{1}\}$. The content of the tables $D_{1}$, $D_{2}$ and $F$ is shown in the following Figure 1.
$D_{1}$ $K_{1}$ $A_{1}^{1}$ $A_{1}^{2}$
$k_{1}$ $a_{1}$ $a_{2}$
$k_{1}$ $a^{\prime}_{1}$ $a_{2}$
$k^{\prime}_{1}$ $a_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
$D_{2}$ $K_{2}$ $A_{2}^{1}$ $A_{2}^{2}$
$k_{2}$ $b_{1}$ $b_{2}$
$k^{\prime}_{2}$ $b_{1}$
$F$ $K_{1}$ $K_{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $m^{\prime}_{1}$
$k_{1}$ $k^{\prime}_{2}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $m_{1}$
Figure 1: The tables of the data warehouse in our running example
We observe that these tables are indeed those of a star schema and that they comply with the two restrictions above regarding missing values. Moreover, it should be emphasized that $D_{1}$ and $F$ do not satisfy $FD$. Indeed, the first two tuples in $D_{1}$ violate $K_{1}\to A_{1}^{1}$ and the second and third tuples in $F$ violate $K_{1}K_{2}\to M_{1}$. On the other hand $D_{2}$ satisfies its two associated functional dependencies $K_{2}\to A_{2}^{1}$ and $K_{2}\to A_{2}^{2}$.
We also stress that the key value $k^{\prime\prime}_{1}$ occurs in $D_{1}$ but not in $F$, whereas the key value $k^{\prime\prime}_{2}$ over $K_{2}$ occurs in $F$ but not in $D_{2}$. These two cases respectively illustrate that key values in a dimension table may not occur in the fact table and that the foreign key constraint between a dimension table and the fact table may not be satisfied (contrary to what is generally assumed in the literature). $\Box$
3.2 True, Consistent and Conflicting Tuples in a Star Schema
The following proposition states important properties regarding m-tuples of the table $m\_Chase(T)$ where $T$ is the table collecting all tuples in the data warehouse. In the remainder of this paper, we refer to such table as a star-table.
Proposition 2
Let $T$ be a star-table over universe $U$. The following hold:
  1. 1.
    For every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$. Consequently, if $\mathbf{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbf{K}))|=1$.
  2. 2.
    For every tuple $k$ over $\mathbf{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$.
  3. 3.
    Moreover, $m\_Chase(T)$ contains the following two kinds of m-tuples:
    1. (a)
      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
      -- $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=\{k_{i_{0}}\}$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$.
      -- for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~{}|~{}(\exists q\in D_{i})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$
    2. (b)
      $\sigma$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$, and
      -- for every $M_{j}\in\mathbf{M}$, $\sigma(M_{j})=\{m_{j}~{}|~{}(\exists t\in F)(t.\mathbf{K}=\sigma(\mathbf{K})\wedge t.M_{j}=m_{j})\}$
      -- for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$, $\sigma(A)=\{a~{}|~{}(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
Proof. 1. This result comes from the fact that, in order to generate multi-valued components of an attribute $A$, the algorithm $m\_Chase$ has to consider a functional dependency whose right hand-side is $A$. As in the case of a star-table, no dependency in $FD$ has its right hand-side in $\mathbf{K}$, the proof of this item is complete.
2. Let $\sigma_{1}$ and $\sigma_{2}$ be in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma_{i})$ for $i=1,2$ and $\sigma_{1}(\mathbf{K})=\sigma_{2}(\mathbf{K})$. Since $\sigma_{1}\neq\sigma_{2}$ (the algorithm eliminates duplicates) there exists $A$ not in $\mathbf{K}$ such that $\sigma_{1}(A)\neq\sigma_{2}(A)$. Since $A$ is not in $\mathbf{K}$, either there is $i_{0}$ in $\{1,\ldots,n\}$ such that $A\in sch^{*}(D_{i_{0}})$ or $A$ is in $\mathbf{M}$. In either case, $FD$ contains a dependency of the form $X\to A$ where $X=K_{i_{0}}$ in the former case, or $X=\mathbf{K}$ in the latter case. Since we have $\sigma_{1}(X)=\sigma_{2}(X)$, applying $m\_Chase$ to $m\_Chase(T)$ changes $\sigma_{1}(A)$ and $\sigma_{2}(A)$ into $\sigma_{1}(A)\cup\sigma_{2}(A)$. As by definition of $m\_Chase(T)$, we have $m\_Chase(m\_Chase(T))=m\_Chase(T)$, we obtain a contradiction which completes the proof.
3. We first notice that, as stated above, for every $K_{i}$ in $\mathbf{K}$ and every $\sigma$ in $m\_Chase(T)$ such that $K_{i}$ is in $sch(\sigma)$, we have $|{\sf tuples}(\sigma(K_{i}))|=1$. This is so because no functional dependency in $FD$ has a right hand-side in $\mathbf{K}$, which makes it impossible to generate conflicts on these attributes.
Moreover, the m-tuples of the first kind are the result of joining two rows $\rho_{1}$ and $\rho_{2}$ in the current table such that $sch(\rho_{1})=sch(\rho_{2})=sch(D_{i_{0}})$, $\rho_{1}(K_{i_{0}})=\rho_{2}(K_{i_{0}})$, and this key value does not occur in $F$ thus preventing these m-tuples to be combined with m-tuples over $sch(F)$. Similarly, the tuples of the second kind are obtained by joining every tuple $km$ in $F$ with tuples $t_{i}$ in $D_{i}$ such that $k.K_{i}=t_{i}.K_{i}$ ($i=1,\ldots,n$). We thus obtain that the set $\Sigma$ of all m-tuples as described above occur in $m\_Chase(T)$.
To end the proof we now argue that applying the $m\_Chase$ procedure to $\Sigma$ has no effect. This is so because for every $X\to A$ in $FD$ and all rows $\rho_{1}$ and $\rho_{2}$ in $\Sigma$ such that $\rho_{1}(X)\cap\rho_{2}(X)\neq\emptyset$, the $X$-values of these rows are reduced to one tuple and thus we have $\rho_{1}(X)=\rho_{2}(X)$. By definition of $\Sigma$, it follows that $\rho_{1}(A)=\rho_{2}(A)$, which completes the proof. $\Box$
$T$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$k_{1}$ $a_{1}$ $a_{2}$
$k_{1}$ $a^{\prime}_{1}$ $a_{2}$
$k^{\prime}_{1}$ $a_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
$k_{2}$ $b_{1}$ $b_{2}$
$k^{\prime}_{2}$ $b_{1}$
$k_{1}$ $k_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $m^{\prime}_{1}$
$k_{1}$ $k^{\prime}_{2}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $m_{1}$
$m\_Chase(T)$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$(k_{1})$ $(k_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(b_{2})$ $(m_{1})$
$(k_{1})$ $(k^{\prime}_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(m^{\prime}_{1}\,m^{\prime\prime}_{1})$
$(k^{\prime}_{1})$ $(k^{\prime\prime}_{2})$ $(a_{1})$ $(m_{1})$
$(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
Figure 2: The star-table $T$ and the m-table $m\_Chase(T)$ of Example 3
In the context of Example 3, the star-table $T$ and its associated m-table $m\_Chase(T)$ are shown in Figure 2. In particular, it can be seen that, in $m\_Chase(T)$, the last m-tuple satisfies item 3.a of Proposition 2 while the first three m-tuples satisfy item 3.b. As will be seen later, these three m-tuples are the ones to be relevant for analytic queries.
[…]
Using Proposition 1, the set ${\sf Confl}_{\min}(\mathcal{T})$ is characterized as follows: $t$ is in ${\sf Confl}_{\min}(\mathcal{T})$ if and only if one of the following two statements holds:
  • there exist $\sigma$ in $m\_Chase(T)$, $i_{0}$ in $\{1,\ldots,n\}$ and $A$ in $sch^{*}(D_{i_{0}})\cap sch(\sigma)\cap sch(t)$ such that $t=k_{i_{0}}a$ is in ${\sf tuples}(\sigma(K_{i_{0}}A))$ and $|{\sf tuples}(\sigma(A))|>1$
  • there exists $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)\cap sch(t)$, there exists $M_{i}$ in $\mathbf{M}\cap sch(\sigma)\cap sch(t)$ such that $t=km_{i}$ is in ${\sf tuples}(\sigma(\mathbf{K}M_{i}))$ and $|{\sf tuples}(\sigma(M_{i}))|>1$.
By complementation with respect to ${\sf True}(\mathcal{T})$, a tuple $t$ is in ${\sf Cons}(\mathcal{T})$ if and only if it has no sub-tuple satisfying one of the above statements.
Example 4
Applying Proposition 1 in the context of Example 3, for which Figure 2 displays the star-table $T$ and the m-table $m\_Chase(T)$, the sets ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$ are as follows:
  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of the tuples in ${\sf tuples}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. Thus, ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of:
    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
  • ${\sf Confl}(\mathcal{T})$ is the set of all true super-tuples of tuples in ${\sf Confl}_{\min}(\mathcal{T})=\{k_{1}a_{1}$, $k_{1}a^{\prime}_{1}$, $k_{1}k^{\prime}_{2}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}m^{\prime\prime}_{1}\}$. The maximal tuples in ${\sf Confl}(\mathcal{T})$ are:
    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$.
  • ${\sf Cons}(\mathcal{T})={\sf True}(\mathcal{T})\setminus{\sf Confl}(\mathcal{T})$. Thus ${\sf Cons}(\mathcal{T})$ is the set of all sub-tuples of:
    -- $k_{1}k_{2}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
    -- $k_{1}k^{\prime}_{2}a_{2}b_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$. $\Box$
4 Repairs and their Basic Properties
In this section we adapt the definition of repair given in LauS25 to the case of a star-table and then we further investigate repairs of star-tables.
4.1 Definition of Repairs
As explained in LauS25 , contrary to most approaches to consistent query answering in the literature, it is not appropriate to define a repair of $T$ as a maximal subset of ${\sf True}(\mathcal{T})$ satisfying $FD$. This is so because it is intuitively justified to define a repair $R$ of $T$ so as ${\sf True}(\mathcal{R})$ contains ${\sf Cons}(\mathcal{T})$. Unfortunately as shown earlier in Example 2, it may happen that a maximal consistent subset S of ${\sf True}(\mathcal{T})$ does not contain all consistent tuples. On the other hand, it has also been shown in LauS25 , that in case of cyclic sets of functional dependencies (that is if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A\in Y$ and $B\in X$ hold), the set ${\sf Cons}(\mathcal{T})$ may not satisfy $FD$.
As in the case of a star-table, the set $FD$ of functional dependencies is clearly acyclic, ${\sf Cons}(\mathcal{T})$ satisfies $FD$. We thus define repairs of star-tables as follows.
Definition 2
Let $T$ be a star-table over universe $U$. A repair $R$ of $T$ is a table over $U$ such that:
  1. 1.
    ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$
  2. 2.
    $R\models FD$
  3. 3.
    For every table $R^{\prime}$ satisfying 1, and 2 above, and such that ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}^{\prime})$.
The set of all repairs of $T$ is denoted by ${\sf Rep}(T)$. $\Box$
It has been shown in LauS25 that the following holds, based on Definition 2:
  • For every table $T$, ${\sf Rep}(T)\neq\emptyset$.
  • If $T\models FD$ then for every $R$ in ${\sf Rep}(T)$, ${\sf True}(\mathcal{R})={\sf True}(\mathcal{T})$. In this case $R$ and $T$ carry the same information, but the tables $R$ and $T$ might not be equal. For example for $FD=\emptyset$ and $T=\{abc,ac\}$, $R=\{abc\}$ is a repair of $T$.
Elaborating briefly on the remark in the second item above, we notice that two distinct tables defined over the same universe and the same set of functional dependencies can have the same sets of true tuples. Databases such as $T$ and $R$ above are said to be equivalent and when we refer to a table $T$ we in fact refer to any table $\widetilde{T}$ such that ${\sf True}(\mathcal{T})={\sf True}(\mathcal{\widetilde{T}})$.
The following basic theorem has been shown in LauS25 to hold whenever the set $FD$ is acyclic.
Theorem 4.1
Let $T$ be a table over universe $U$ and $FD$ an acyclic set of functional dependencies over $U$. Then:
${\sf Cons}(\mathcal{T})=\bigcap_{R\in{\sf Rep}(T)}{\sf True}(\mathcal{R}).$
$\Box$
4.2 Repairs of Star-Tables
In this section, we show that, in the case of star-tables, repairs satisfy important properties that don’t hold in general.
Another important specificity of star-tables is that our notion of repair coincides with that in the literature, applied to ${\sf True}(\mathcal{T})$. More precisely, we show that if $T$ is a star-table and $S$ is a maximal consistent subset of ${\sf True}(\mathcal{T})$, then $S$ is a repair of $T$ in the sense of Definition 2, implying in particular that ${\sf True}(\mathcal{S})$ contains ${\sf Cons}(\mathcal{T})$. As a consequence, in Definition 2, the first item can be more simply stated as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$. This result relies on the following preliminary two lemmas.
Lemma 1
Let $T$ be a star-table over universe $U$, $S$ a subset of ${\sf True}(\mathcal{T})$ and $t$ a tuple in ${\sf Cons}(\mathcal{T})$. If $S\models FD$ then $S\cup\{t\}\models FD$.
Proof. Since $S\models FD$, every $\sigma$ in the m-table $m\_Chase(S)$ is such that, for every $A$ in $sch(\sigma)$, $\sigma(A)=(a)$. Thus $m\_Chase(S)$ can be seen as a table that we denote by $S^{*}$. Moreover, denoting by $S_{t}$ the table $S\cup\{t\}$, we have $m\_Chase(S_{t})=m\_Chase(S^{*}\cup\{t\})$. Let $S^{*}_{t}=S^{*}\cup\{t\}$, and let us consider the computation of $m\_Chase(S^{*}_{t})$.
To this end, given $q_{1}$ and $q_{2}$ in $S^{*}_{t}$ and $X\to A$ in $FD$ such that $q_{1}.X=q_{2}.X$, the only possible cases are as follows:
-- If $q_{1}$ and $q_{2}$ are in $S^{*}_{t}$, then either $q_{1}$ and $q_{2}$ are not defined over $A$ or, they are both defined such that $q_{1}.A=q_{2}.A$. In this case $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are not defined over $A$, then $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are defined over $A$. Since $t\in{\sf Cons}(\mathcal{T})$, it is not possible that $q_{1}.A\neq t.A$. Thus in this case again, $m\_Chase$ does not change $S^{*}_{t}$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is defined over $A$ and $q_{2}$ is not defined over $A$. Then $m\_Chase$ changes $t$ into $ta$ where $a=q_{1}.A$.
-- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is not defined over $A$ and $q_{2}$ is defined over $A$. Then $m\_Chase$ changes $q_{1}$ into $qa$ where $a=t.A$.
Based on the previous cases, we denote by $\Sigma$ the table obtained from $S^{*}_{t}$ by the following transformations:
$(1)$ For every $q$ in $S^{*}$ and every $X\to A$ in $FD$ such that $q.X=t.X$, $q$ is not defined over $A$ and $t.A=a$, in $\Sigma$, $q.A$ is set to $a$.
$(2)$ For every $X\to A$ in $FD$ such that $q.X=t.X$, $q.A=a$ and $t$ is not defined over $A$, in $\Sigma$, $t.A$ is set to $a$.
Since in a star schema, for every attribute $A$ in $U$ there is at most one functional dependency whose right hand-side is $A$, the construction of $\Sigma$ cannot generate conflicts, thus entailing that $\Sigma$ contains no conflicts. We now show that $m\_Chase(\Sigma)=\Sigma$.
Indeed, let $q_{1}$ and $q_{2}$ be in $\Sigma$ and $X\to A$ be in $FD$. The only possibility for $m\_Chase$ to change $\Sigma$ is that $q_{1}.X=q_{2}.X$, $q_{1}.A=a$ and $q_{2}$ is not defined over $A$. As this does not happen in $S^{*}$ (because $m\_Chase(S^{*})=S^{*}$), and as this does not happen in $\Sigma$ (by definition of $\Sigma$), we obtain that $m\_Chase(\Sigma)=\Sigma$. It follows that $m\_Chase(S^{*}_{t})=\Sigma$, and since we have seen that no conflicts occur in $\Sigma$, $S^{*}_{t}\models FD$ holds. The proof is therefore complete. $\Box$
The following lemma is a consequence of Lemma 1.
Lemma 2
Let $T$ be a star-table over universe $U$, $S$ a subset of ${\sf True}(\mathcal{T})$. If $S\models FD$ then $S\cup{\sf Cons}(\mathcal{T})\models FD$.
Proof. If $N$ denotes the number of tuples in ${\sf Cons}(\mathcal{T})$, applying Lemma 1 successively to each tuple of ${\sf Cons}(\mathcal{T})$ generates a sequence of sets $S_{i}$ ($i=1,\ldots,N$) such that $S_{i}\models FD$ and ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{S}_{N})$. Since ${\sf True}(\mathcal{S}_{N})={\sf True}(S\cup{\sf Cons}(\mathcal{T}))$, the proof is complete. $\Box$
We now state the expected result that every maximal and consistent subset of ${\sf True}(\mathcal{T})$ defines a repair of $T$.
Proposition 3
Let $T$ be a star-table over universe $U$, and $S$ a subset of ${\sf True}(\mathcal{T})$ such that $S\models FD$ and there is no subset $S^{\prime}$ of ${\sf True}(\mathcal{T})$ such that $S^{\prime}\models FD$ and ${\sf True}(\mathcal{S})$ is a strict subset of ${\sf True}(\mathcal{S}^{\prime})$. Then $S$ is in ${\sf Rep}(T)$.
Proof. By Definition 2, in order to show that $S\in{\sf Rep}(T)$, we only have to prove that ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{S})$. If such is not the case, by Lemma 2, ${\sf True}(\mathcal{S})\cup{\sf Cons}(\mathcal{T})\models FD$. We thus obtain a consistent strict super-set of ${\sf True}(\mathcal{S})$, which is a contradiction with the last statement in the proposition. The proof is therefore complete. $\Box$
As a consequence of Proposition 3, we show that in the case of a star-table, given a true tuple $t$, there exists a repair in which this tuple is true, and that moreover, if this tuple is conflicting, there exists a repair in which this tuple is not true. It should be noticed here that, the second result is shown in LauS25 , whereas the first one is not.
Proposition 4
Let $T$ be a star-table over universe $U$. Then the following statements hold:
$\bullet$ For every $t$ in ${\sf True}(\mathcal{T})$ there exists $R$ in ${\sf Rep}(T)$ such that $t$ is in ${\sf True}(\mathcal{R})$.
$\bullet$ For every $t$ in ${\sf Confl}(\mathcal{T})$ there exists $R$ in ${\sf Rep}(T)$ such that $t$ is not in ${\sf True}(\mathcal{R})$.
Proof. To show the first statement, based on the fact that for every tuple $t$ in $\mathcal{T}$, $\{t\}\models FD$ trivially holds, as ${\sf True}(\mathcal{T})$ is finite, there exists a maximal subset $S$ of ${\sf True}(\mathcal{T})$ such that $t\in S$ and $S\models FD$. By Proposition 3, $S$ is in ${\sf Rep}(T)$. Thus the proof of the first statement of the proposition is complete. The second statement being the subject of Proposition 4 of LauS25 , the proof of the proposition is complete. $\Box$
As a last property of repairs of star-tables, we show below that if tuples are conflicting, then for every repair $R$, one of the tuples involved in the conflict is true in $R$.
Proposition 5
Let $T$ be a star-table over universe $U$. For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{k}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$. Then, for every $R$ in ${\sf Rep}(T)$, there exists $i_{0}$ in $\{1,\ldots,k\}$ such that $xa_{i_{0}}$ is in ${\sf True}(\mathcal{R})$.
Proof. It is easy to show that the proposition holds for $k=1$, since in this case $xa_{1}$ is in ${\sf Cons}(\mathcal{T})$, and thus in ${\sf True}(\mathcal{R})$ for every $R$ in ${\sf Rep}(T)$. We now assume that $k>1$, which implies that for every $\sigma$ in $m\_Chase(T)$ such that $XA\subseteq sch(\sigma)$ and $x$ is in ${\sf tuples}(\sigma(X))$, ${\sf tuples}(\sigma(XA))=(xa_{1},\ldots,xa_{k})$ with $k>1$. Moreover, given the form of functional dependencies in $FD$, either $(a)$ $X=\mathbf{K}$ and $A$ is in $\mathbf{M}$, or $(b)$ there exists $i_{0}$ in $\{1,\ldots,n\}$ such that $X=K_{i_{0}}$ and $A$ is in $sch^{*}(D_{i_{0}})$.
Let $R$ be in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no $xa_{l}$ from $\{xa_{1},\ldots,xa_{k}\}$, and let $l_{0}$ in $\{1,\ldots,k\}$. Denoting by $R^{\prime}$ the table $R\cup\{xa_{l_{0}}\}$ we show that in either case $(a)$ or $(b)$ above, $R^{\prime}\models FD$. To this end, we first notice that, since for every $k$ over $\mathbf{K}$ and for any $k_{i}$ over $K_{i}$, $k$ and $k_{i}$ are in ${\sf Cons}(\mathcal{T})$, the tuple $x$ is in ${\sf True}(\mathcal{R})$. Hence, ${\sf True}(\mathcal{R})$ contains tuples whose $X$-value is $x$ and for every such tuple, the $A$-value is missing. The cases $(a)$ and $(b)$ are then handled as follows when computing $m\_Chase(R^{\prime})$: every $t$ in ${\sf True}(\mathcal{R}^{\prime})$ such that $t.X=x$ is considered and its $A$-value is set to $a_{l_{0}}$. Denoting by $\Sigma$ the resulting m-table and let $X^{\prime}\to A^{\prime}$ be a dependency in $FD$, and let $\sigma_{1}$ and $\sigma_{2}$ be two m-tuples in $\Sigma$ such that $\sigma_{1}(X^{\prime})=\sigma_{2}(X^{\prime})$ where $X^{\prime}$ is either $\mathbf{K}$ or $K_{j}$.
-- If $X^{\prime}=X$ then by definition of $\Sigma$, $\sigma_{1}(A)=\sigma_{2}(A)=(a_{l_{0}})$. Thus $\Sigma$ is not changed.
-- If $X^{\prime}=K_{j}$, $X=K_{i}$ and $i\neq j$, $\sigma_{1}$ and $\sigma_{2}$ are both in $R^{\prime}$ and their $A^{\prime}$-components have not been changed when building $\Sigma$ (because $K_{j}\neq K_{i}$ implies $A^{\prime}\neq A$). Since $m\_Chase({\sf True}(\mathcal{R}^{\prime}))={\sf True}(\mathcal{R}^{\prime})$ and $R^{\prime}\models FD$, either $A^{\prime}$ is in $sch(\sigma_{1})\cap sch(\sigma_{2})$ and $\sigma_{1}(A^{\prime})=\sigma_{2}(A^{\prime})$, or $A^{\prime}$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$. In either case, $\Sigma$ is not changed.
-- If $X^{\prime}=K_{j}$ and $X=\mathbf{K}$, since as above we have $A^{\prime}\neq A$, $m\_Chase$ does not change $\Sigma$.
-- If $X^{\prime}=\mathbf{K}$ and $X=K_{i}$, then in ${\sf True}(\mathcal{R}^{\prime})$, we have $\sigma_{1}(X)=\sigma_{2}(X)=x$, and $A$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$. By definition of $\Sigma$, $\sigma_{1}(A)$ and $\sigma_{2}(A)$ are set to $(a_{l_{0}})$, and then here again, $m\_Chase$ does not change $\Sigma$.
We thus obtain that $\Sigma=m\_Chase(\mathcal{R}^{\prime})$ and that $R^{\prime}\models FD$, because $R\models FD$ and no conflicts occur in $\Sigma$. Moreover, as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}^{\prime})$. It follows from Definition 2 that $R$ can not be in ${\sf Rep}(T)$ since the inclusion ${\sf True}(\mathcal{R})\subset{\sf True}(\mathcal{R}^{\prime})$ is strict (because $xa_{l_{0}}$ is in ${\sf True}(\mathcal{R}^{\prime})$ but not in ${\sf True}(\mathcal{R})$). As this is a contradiction with our hypothesis that $R$ is in ${\sf Rep}(T)$, the proof is complete. $\Box$
We now show how to characterize all repairs of a star-table $T$, based on $m\_Chase(T)$. To this end, we define the following process (P):
Step P1. For every $i$ in $\{1,\ldots,n\}$ and every $A_{i}^{j}$ in $sch^{*}(D_{i})$, if $k_{i}$ is a $K_{i}$-value occurring in $m\_Chase(T)$, we choose one $A_{i}^{j}$-value in the set $\sigma(A_{i}^{j})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$. Denoting by $\varphi_{i}^{j}(k_{i})$ this value, we notice that $\varphi_{i}^{j}(k_{i})$ is well defined because, thanks to $K_{i}\to A_{i}^{j}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$ have the same value over $A_{i}^{j}$.
Step P2. For every $l$ in $\{1,\ldots,p\}$, if $k$ is a $\mathbf{K}$-value occurring in $m\_Chase(T)$, we choose one $M_{l}$-value in the set $\sigma(M_{l})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$. Denoting by $\varphi_{l}(k)$ this value, we notice that, as above, $\varphi_{l}(k)$ is well defined because, thanks to $\mathbf{K}\to M_{l}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$ have the same value over $M_{l}$.
Step P3. Denoting by $\varphi$ the set of all $\varphi_{i}^{j}(k_{i})$ and all $\varphi_{l}(k)$ defined above, for every $\sigma$ in $m\_Chase(T)$, let $t_{\varphi}(\sigma)$ be the tuple such that $sch(t_{\varphi}(\sigma))=sch(\sigma)$, defined by:
  • For every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$, then $t_{\varphi}(\sigma).K_{i}=k_{i}$ and for every $j$ such that $A_{i}^{j}\in sch(\sigma)$, $t_{\varphi}(\sigma).A_{i}^{j}=\varphi_{i}^{j}(k_{i})$.
  • If $\mathbf{K}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$, for every $l=1,\ldots,p$ such that $M_{l}\in sch(\sigma)$, let $t_{\varphi}(\sigma).M_{l}=\varphi_{l}(k)$.
Denoting by $R_{\varphi}$ the table $R_{\varphi}=\{t_{\varphi}(\sigma)~{}|~{}\sigma\in m\_Chase(T)\}\cup{\sf Cons}(\mathcal{T})$, the following proposition states that $R_{\varphi}$ is a repair and that all repairs are obtained through such process.
Proposition 6
Let $T$ be a star-table over $U$. $R$ is a repair of $T$ if and only if there is $\varphi$ as defined above such that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$.
Proof. See Appendix A. $\Box$
We illustrate repairs of star-tables in the context of our running Example 3 as follows.
$R_{1}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k_{1}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $a_{1}$ $m_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
$R_{2}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k_{1}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $a_{1}$ $m_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
$R_{3}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k_{1}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $a_{1}$ $m_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
$R_{4}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$ $m_{1}$
$k_{1}$ $k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k_{1}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a^{\prime}_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime\prime}_{1}$
$k^{\prime}_{2}$ $a_{1}$ $a_{2}$ $b_{1}$ $m^{\prime}_{1}$
$k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $a_{1}$ $m_{1}$
$k^{\prime\prime}_{1}$ $a^{\prime}_{1}$ $a^{\prime}_{2}$
Figure 3: The repairs of the star-table $T$ of Example 3
Example 5
In the context of Example 3, $T$ has four repairs as shown in Figure 3. It can be seen from the m-table $m\_Chase(T)$ in Figure 2 that, in each of these four tables:
  • the first two rows are generated using the first m-tuple in $m\_Chase(T)$, namely
    $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$,
  • the next five rows are generated using the second m-tuple in $m\_Chase(T)$, namely
    $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$,
  • the last two rows are generated using respectively the last two m-tuples in $m\_Chase(T)$, namely $(k^{\prime}_{1})(k^{\prime\prime}_{2})(a_{1})(m_{1})$ and $(k^{\prime\prime}_{1})(a^{\prime}_{1})(a^{\prime}_{2})$.
To illustrate how process (P) works, we apply it to generate the repair $R_{1}$ shown in Figure 3.
  • Step P1. Regarding attributes in $sch^{*}(D_{1})$, for $A_{1}^{1}$, the three $K_{1}$-values $k_{1}$, $k^{\prime}_{1}$ and $k^{\prime\prime}_{1}$ are respectively associated with $a_{1}$, $a_{1}$ and $a^{\prime}_{1}$, while for $A_{1}^{2}$, the two $K_{1}$-values $k_{1}$ and $k^{\prime\prime}_{1}$ are associated with $a_{2}$ and $a^{\prime}_{2}$, respectively (no $A_{1}^{2}$-value is associated with $k^{\prime}_{1}$ in $T$).
    Regarding attributes in $sch^{*}(D_{2})$, for $A_{2}^{1}$, the two $K_{2}$-values $k_{2}$ and $k^{\prime}_{2}$ are both associated with $b_{1}$ (no $A_{2}^{1}$-value is associated with $k^{\prime\prime}_{2}$ in $T$), and for $A_{2}^{2}$, the only $K_{2}$-values $k^{\prime}_{2}$ is associated with $b_{2}$ (no $A_{2}^{2}$-value is associated with $k^{\prime}_{2}$ or with $k^{\prime\prime}_{2}$ in $T$).
  • Step P2. The only measure attribute is here $M_{1}$ and the $\mathbf{K}$-values to be considered are $k_{1}k_{2}$, $k_{1}k^{\prime}_{2}$ and $k^{\prime}_{1}k^{\prime\prime}_{2}$ which are respectively associated with $M_{1}$-values $m_{1}$, $m^{\prime\prime}_{1}$ and $m_{1}$.
  • Step P3. Considering every m-tuple $\sigma$ of the table $m\_Chase(T)$ shown in Figure 2, we obtain the following four tuples: $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$ and $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
We obtain that $R_{\varphi}$ contains the four tuples above along with the tuples in ${\sf Cons}(\mathcal{T})$ that have been characterized in Example 3. Moreover, inspecting thoroughly the table $R_{1}$ and the set $R_{\varphi}$ shows that the two sets yield the same set of true tuples, corresponding to the same repair of $T$. $\Box$
5 Consistent Query Answering
Hereafter, given a star-table $T$, we consider different kinds of queries on $T$, all based on the SQL syntax:
  1. 1.
    Standard queries or simply queries, to be defined in the next section. These queries are simple projection-selection queries of the form select $X$ from $T$ where $\Gamma$, where $X$ is a list of attributes and $\Gamma$ an optional selection condition.
  2. 2.
    Analytic queries with no group-by statement, defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$, where $X$ and $\Gamma$ are as above, and where $aggr$ is an aggregation operator and $M_{i}$ is a measure attribute.
  3. 3.
    Analytic queries with a group-by statement, also defined in Section 5.3. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above.
  4. 4.
    Analytic queries with a group-by-having statement, dealt with in Section 6.4. These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above, and where $\Upsilon$ is a boolean expression involving aggregates.
5.1 Standard Queries
As mentioned just above, we use SQL as the query language and, as we query a single table $T$, the (standard) queries $Q$ that we consider have one of the following two forms:
$Q:$ select $X$ from $T$  or   $Q:$ select $X$ from $T$ where $\Gamma$
In either of these forms, $X$ is an attribute list seen as a relation schema, and in the second form the where clause specifies a selection condition $\Gamma$. As in SQL the where clause in a query is optional, the generic form of a query $Q$ is denoted by
$Q:$ select $X$ from $T$ [where $\Gamma$]
The set of all attributes occurring in $\Gamma$ is called the schema of $\Gamma$, denoted by $sch(\Gamma)$; and the attribute set $X\cup sch(\Gamma)$ is called the schema of $Q$, denoted by $sch(Q)$.
A selection condition $\Gamma$ is a well-formed formula involving the usual connectors $\neg$, $\vee$ and $\wedge$ and built up from atomic Boolean comparisons of one of the following forms: $A\,\theta\,a$ or $A\,\theta\,A^{\prime}$, where $\theta$ is a comparison predicate, $A$ and $A^{\prime}$ are attributes in $U$ whose domain elements are comparable through $\theta$, and $a$ is in $dom(A)$.
Given a selection condition $\Gamma$, we denote by $Sat(\Gamma)$ the set of all tuples in $\mathcal{T}(sch(\Gamma))$ satisfying $\Gamma$, as defined below:
$\bullet$ if $\Gamma$ is of the the form $A\,\theta\,a$, $Sat(\Gamma)=\{t\in\mathcal{T}(sch(\Gamma))~{}|~{}t.A\,\theta\,a\}$,
$\bullet$ if $\Gamma$ is of the the form $A\,\theta\,B$, $Sat(\Gamma)=\{t\in\mathcal{T}(sch(\Gamma))~{}|~{}t.A\,\theta\,t.B\}$,
$\bullet$ if $\Gamma$ is of the form $\Gamma_{1}\vee\Gamma_{2}$, $Sat(\Gamma)=Sat(\Gamma_{1})\cup Sat(\Gamma_{2})$,
$\bullet$ if $\Gamma$ is of the form $\Gamma_{1}\wedge\Gamma_{2}$, $Sat(\Gamma)=Sat(\Gamma_{1})\cap Sat(\Gamma_{2})$,
$\bullet$ if $\Gamma$ is of the form $\neg\Gamma_{1}$, $Sat(\Gamma)=\mathcal{T}(sch(\Gamma))\setminus Sat(\Gamma_{1})$.
As usual in the literature ArenasBC99 , we define the consistent answer of such a query $Q$ as the intersection of the answers to the query in every repair. The formal definition follows.
Definition 3
Let $T$ be a table over universe $U$ and $FD$ an acyclic set of functional dependencies over $U$. Given the query $Q:$ select $X$ from $T$ [where $\Gamma$], the consistent answer to $Q$ in $T$, denoted by ${\sf C\_ans}(Q)$, is defined by:
${\sf C\_ans}(Q)=\bigcap_{R\in{\sf Rep}(T)}{\sf Ans}\left(Q^{[R]}\right)$
where $Q^{[R]}$ is the query select $X$ from $R$ [where $\Gamma$], and ${\sf Ans}(Q^{[R]})$ is the answer to $Q^{[R]}$ in $R$. Formally, ${\sf Ans}(Q^{[R]})$ is the set of all tuples $x$ over $X$ such that there exits $\gamma$ in $Sat(\Gamma)$ such that $x\gamma$ is a tuple over $sch(Q)$ that belongs to ${\sf True}(\mathcal{R})$. $\Box$
It is important to notice that, as a consequence of Theorem 4.1, given a query $Q$, we have ${\sf C\_ans}(Q)\subseteq{\sf Cons}(\mathcal{T})$. This is so because every $x$ in ${\sf C\_ans}(Q)$ is true in every repair of $T$. We also point out that if $Q$ involves no selection condition then Theorem 4.1 allows for characterizing ${\sf C\_ans}(Q)$ as the set of all tuples in ${\sf Cons}(\mathcal{T})$ whose schema is $X$.
We however recall that the issue of easily characterizing ${\sf C\_ans}(Q)$ in the case of a query $Q$ involving a selection condition remains open. We recall in this respect from LauS25 , that ${\sf C\_ans}(Q)$ can be bounded as follows.
Proposition 7
Let $T$ be a table over universe $U$ and its associated acyclic set of functional dependencies $FD$. Given a query $Q:{\tt select}$ $X$ where $\Gamma$, let:
  • ${\sf C\_ans}^{-}(Q)=\{x\in\mathcal{T}(X)~{}|~{}(\exists t\in\mathcal{T}(sch(Q))(t\in{\sf Cons}(\mathcal{T})\wedge t.X=x\wedge t.sch(\Gamma)\in Sat(\Gamma)\}$
  • ${\sf C\_ans}^{+}(Q)=\{x\in\mathcal{T}(X)~{}|~{}(\exists t\in\mathcal{T}(sch(Q))(t\in{\sf True}(\mathcal{T})\wedge t.X=x\,\wedge$
    $x\in{\sf Cons}(\mathcal{T})\wedge t.sch(\Gamma)\in Sat(\Gamma)\}$
Then, ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ are respectively a lower bound and an upper bound of ${\sf C\_ans}(Q)$, i.e., the following inclusions hold: ${\sf C\_ans}^{-}(Q)\subseteq{\sf C\_ans}(Q)\subseteq{\sf C\_ans}^{+}(Q)$. $\Box$
Moreover, based on Proposition 1, it is shown in LauS25 that the two bounds ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ can be easily computed by means of one scan of the m-table $m\_Chase(T)$. It turns out that approximating the consistent answer to any query in our approach is polynomial. For example, in the context of Example 3,
  • For $Q_{1}:$ select $K_{1}$, $K_{2}$, $M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), we have ${\sf C\_ans}^{-}(Q_{1})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$ and ${\sf C\_ans}^{+}(Q_{1})=\{k_{1}k_{2}m_{1},k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$. On the other hand, it can be seen from Figure 3 that ${\sf C\_ans}(Q_{1})={\sf C\_ans}^{-}(Q_{1})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$.
  • For $Q_{2}:$ select $K_{1}$, $K_{2}$, $A_{2}^{1}$, $M_{1}$, from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$), then we have ${\sf C\_ans}^{-}(Q_{2})=\emptyset$ and ${\sf C\_ans}^{+}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. On the other hand, it can be seen from Figure 3 that ${\sf C\_ans}(Q_{2})={\sf C\_ans}^{+}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$.
5.2 Consistent Query Answering and Star Schemas
As mentioned in our introductory section, we can compute efficiently exact consistent answers (instead of bounds of exact consistent anwers as in LauS25 ) if the selection condition meets two requirements. The first requirement is that satisfaction of the selection condition can be tested attribute by attribute, as stated in the following definition.
Definition 4
A selection condition $\Gamma$ is said to be independent if $\Gamma$ is equivalent to the conjunction $\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ where for every $i=1,\ldots,k$, $\Gamma(A_{i})$ is a selection condition involving only attribute $A_{i}$. $\Box$
By Definition 4, if $\Gamma=\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ is independent, then $sch(\Gamma)=A_{1}\ldots A_{k}$ and a tuple $\gamma$ over $sch(\Gamma)$ is in $Sat(\Gamma)$ if and only if for every $i=1,\ldots,k$, $\gamma.A_{i}$ is in $Sat(\Gamma(A_{i}))$.
For example, $\Gamma=(A_{1}\leq 0)\vee((A_{1}\geq 4)\wedge((A_{2}=a))$ can be written as $((A_{1}\leq 0)\vee(A_{1}\geq 4))\wedge((A_{1}\leq 0)\vee(A_{2}=a))$. Thus, $\Gamma$ is not independent. On the other hand, $\Gamma^{\prime}=((A_{1}\leq 4)\vee(A_{1}\geq 0))\wedge(A_{2}=a)$ is obviously independent with $\Gamma^{\prime}(A_{1})=(A_{1}\geq 4)\vee(A_{1}\leq 0)$ and $\Gamma^{\prime}(A_{2})=(A_{2}=a)$, and we have $Sat(\Gamma^{\prime})=\{(a_{1},a_{2})\in adom(A_{1})\times adom(A_{2})~{}|~{}(a_{1}\leq 0\vee a_{1}\geq 4)\wedge(a_{2}=a)\}$.
We emphasize that, as shown by the example just above, independent selection conditions may involve disjunctions. Therefore, and contrary to most existing approaches to consistent query answering ArenasBC99 ; Wijsen09 , our approach does not restrict selection conditions to be conjunctive.
Another restriction that we consider on queries is that key attributes in $\mathbf{K}$ are not allowed in selection conditions. We point out that this syntactical restriction on queries does not rule out queries of practical interest, because in most relevant queries, selections are expressed using meaningful attributes rather than numerical abstract key-values. Notice in this respect that key attributes are mainly used in selection conditions to express join conditions, which are not explicit in our chase-based approach.
  • From now on, it is assumed, even when not explicitly mentioned, that selection conditions are independent and involve no key attributes from $\mathbf{K}$.
Actually, what the above assumption implies is that, in a star-table, given a query $Q:{\tt select}$ $X$ from $T$ where $\Gamma$, for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$ holds, we have $Y\subseteq X\setminus sch(\Gamma)$.
The following proposition states one of the main contributions of the paper, namely that, under the restriction above, the consistent answer to any query having an independent selection condition can be easily computed.
Proposition 8
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
  1. 1.
    $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$,
  2. 2.
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
  3. 3.
    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
Proof. See Appendix B. $\Box$
An important consequence of Proposition 8 is that the data complexity of consistent query answering to projection-selection-join queries on star schemes is polynomial, under the restriction that the selection condition is independent and involves no keys. The following corollary shows that the third item in Proposition 8 can be more simply stated when $\mathbf{K}$ is a subset of $X$.
Corollary 1
Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. If $\mathbf{K}\subseteq X$, ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
  1. 1.
    $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$,
  2. 2.
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
  3. 3.
    ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
Proof. As for every $A$ in $U$, there exists $Y\to A$ in $FD$ with $Y\subseteq\mathbf{K}$, if $\mathbf{K}\subseteq X$, for every $B$ in $sch(\Gamma)$ there exists $Y\to B$ in $FD$ with $Y\subseteq X$. Therefore, in this case, item 3 in Proposition 8 can be stated as: for every $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B)$. Since $\Gamma$ is independent, as mentioned right after Definition 4, this is equivalent ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. The proof is therefore complete. $\Box$
In the following example, we illustrate Proposition 8 and Corollary 1, and we show the impact of our two restrictions on the form of the selection conditions.
Example 6
In the context of Example 3, we consider the star-table $T$, and its associated m-table $m\_Chase(T)$ as shown in Figure 2, let $Q_{2}$ be the query defined by:
$Q_{2}:$ select $K_{1}$, $K_{2}$, $A_{2}^{1}$, $M_{1}$, from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$).
As already noticed after stating Proposition 7, we have ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. On the other hand, the selection condition in this query obviously satisfies our restrictions, since it is independent and involves no key attributes. Applying Proposition 8 or Corollary 1 (because $K_{1}K_{2}\subseteq sch(Q_{2})$) yields the following:
  • The m-tuples $\sigma$ in $m\_Chase(T)$ such that $sch(Q_{i})\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$, are $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ and $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$.
  • Among these m-tuples, only $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ is such that for every $Y\to B$ in $FD$ such that $YB\subseteq X_{i}$, $|{\sf tuples}(\sigma(B))|=1$.
  • This m-tuple satisfies ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
We therefore obtain the expected consistent answer, namely ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. Now, let $Q_{3}$ be defined by:
$Q_{3}:$ select $K_{1},K_{2},A^{2}_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ and $M_{1}=m_{1}$) or ($A_{1}^{1}=a^{\prime}_{1}$ and $M_{1}=m^{\prime\prime}_{1}$).
In this case, the selection condition $\Gamma_{3}$ is clearly not independent, and it can be seen from Figure 3 that ${\sf C\_ans}(Q_{3})$ is in fact empty, because:
-- in $R_{2}$, $k_{1}k_{2}a_{2}$ is only associated with $a^{\prime}_{1}m_{1}$ which does not belong to $Sat(\Gamma_{3})$,
-- in $R_{1}$, $k_{1}k^{\prime}_{2}a_{2}$ is only associated with $a_{1}m^{\prime}_{1}$ which does not belong to $Sat(\Gamma_{3})$.
On the other hand, considering the third item in Proposition 8, we emphasize the following: $k_{1}k^{\prime}_{2}a_{2}$ is discarded from the consistent answer because $a_{1}m^{\prime}_{1}$ is in ${\sf tuples}(\sigma_{2}(A_{1}^{1}M_{1}))$ but not in $Sat(\Gamma_{3})$. However, $\sigma_{1}$ can be seen as satisfying the third item in the proposition for the following reasons: the dependencies to consider are $K_{1}\to A_{1}^{1}$ and $K_{1}K_{2}\to M_{1}$, and ${\sf tuples}(\sigma_{1}(A^{1}_{1}))$, respectively ${\sf tuples}(\sigma_{1}(M_{1}))$, is a sub-set of the set of all $A_{1}^{1}$-values, respectively all $M_{1}$-values, occurring in $Sat(\Gamma_{3})$. A characterization of the consistent answer to a query involving a non independent selection condition is currently unknown to the authors.
Now, to illustrate our restriction on the key-attributes with respect to selection conditions, let $Q_{4}:$ select $A_{1}$ from $T$ where ($K_{1}=k_{1}$). It should be clear that the selection condition of $Q_{4}$ is independent, and that ${\sf C\_ans}(Q_{4})=\emptyset$ (because, by Figure 3, neither $a_{1}$ nor $a^{\prime}_{1}$ is associated with $k_{1}$ is every repair). On the other hand, it is easy to see that the conditions in Proposition 8 or Corollary 1 are satisfied by the first two m-tuples of $m\_Chase(T)$ (see Figure 2). A characterization of the consistent answer to a query with a selection condition involving a key attribute is currently unknown to the authors. $\Box$
In the next two sections, we consider analytic queries and see how their consistent answer can be defined and effectively computed, relying on Corollary 1.
5.3 Analytic Queries and their Consistent Answers
In the literature Ullman , an analytic query is a query involving an aggregate function among $count$, $min$, $max$, $sum$, and possibly a group-by clause. Formally, given a data warehouse over a star schema, an analytic query $\mathcal{AQ}$ has one of the following two forms:
  • $\mathcal{AQ}:{\tt select}$ $aggr(M_{i})$ from $\varphi$ where $\Gamma$, or
  • $\mathcal{AQ}:{\tt select}$ $X$, $aggr(M_{i})$ from $\varphi$ where $\Gamma$ group by $X$
where $\varphi$ is the join of the fact table $F$ with all dimension tables $D_{1},\ldots,D_{n}$, $X$ is a relation schema, and for $j=1,\ldots,p,$ $aggr$ is an aggregate function such as $count$, $min$, $max$, $sum$ and $M_{i}$ is a measure attribute.
In all traditional approaches to data warehouses, it is generally assumed that the tables in the data warehouse have no missing values and that the key-foreign-key constraints between $F$ and all dimension tables $D_{i}$ are satisfied. In this context, the answer to $\mathcal{AQ}$ is as follows:
  • If $\mathcal{AQ}$ involves no group-by clause, ${\sf Ans}(\mathcal{AQ})$ is the value of the aggregation evaluated over the set of all tuples in $\varphi$ satisfying the condition $\Gamma$. We notice that the aggregate may involve no attribute, when expressed as $count(*)$.
  • If $\mathcal{AQ}$ involves a statement group-by $X$, ${\sf Ans}(\mathcal{AQ})$ is the set of all pairs $\langle x,v_{x}\rangle$ where $x$ is such that there exists a tuple in $\varphi$ satisfying $\Gamma$ and whose $X$-value is $x$, and where $v_{x}$ is the value of the aggregation evaluated over all tuples in $\varphi$ satisfying $\Gamma$ and whose $X$-value is $x$.
Now, if the underlying data warehouse does not satisfy all functional dependencies, this traditional semantics of answer to analytic queries has to be revisited. When no dependency of the form $\mathbf{K}\to M_{i}$ is considered, this has been done in ArenasBCHRS03 , in the case of queries with no group-by clause and in FuxmanFM05 when a group-by clause occurs in the query. As shown in DixitK19 , in either case, the ‘repair semantics’ of the consistent answer to an analytic query consists intuitively in producing an interval in which aggregate values fall, when answering the query in any repair. The reader is referred to the forthcoming Section 7 for a more precise relationship between these approaches and our work.
In our approach, we follow the same line regarding analytic queries and their consistent answers, and we emphasize that, as in the previous section, all selection conditions are assumed to satisfy the restrictions that are independent and that they involve no attribute from $\mathbf{K}$.
Definition 5
Let $T$ be a star-table over universe $U$. We call analytic query with or without group-by clause, a query of the generic form:
$\mathcal{AQ}:~{}$select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$]
where the group-by clause may be omitted, in which case $X$ is not in the select clause.
The consistent answer to analytic query $\mathcal{AQ}$, denoted by ${\sf C\_ans}(\mathcal{AQ})$, is defined as follows:
(A) If $\mathcal{AQ}$ involves no group-by clause, then:
  • If there exists $R$ in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})={\tt NULL}$ if $aggr\neq count$, and ${\sf C\_ans}(\mathcal{AQ})=0$ if $aggr$ is $count$.
  • If for every $R$ in ${\sf Rep}(T)$, ${\sf True}(\mathcal{R})$ contains at least one tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$ such that:
    • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})=d$,
    • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})=glb$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})=lub$.
(B) If $\mathcal{AQ}$ involves a group-by clause, ${\sf C\_ans}(\mathcal{AQ})$ is a set of pairs of the form $\langle x,[glb,lub]\rangle$ such that:
  • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})$ contains $(x\,d)$,
  • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})$ respectively contain $(x\,glb)$ and $(x\,lub)$. $\Box$
The following remarks are in order regarding Definition 5:
  • In case (A), a NULL answer means that the consistent answer to the query $Q:~{}{\tt select}$ $\mathbf{K}$ from $T~{}{\tt where}~{}\Gamma$ is empty. In this case the aggregate cannot be computed if different than $count$, and if the aggregate is $count$, then the expected answer is 00. This explains why we return NULL in the former case.
  • In case (B), all $X$-values $x$ occurring in ${\sf C\_ans}(\mathcal{AQ})$ must be consistent, because the first condition requires that $x$ be true in every repair of $T$. On the other hand, if $x$ is consistent and if there is a repair $R$ such that ${\sf True}(\mathcal{R})$ contains no tuple of the form $x\,\gamma$ where $\gamma$ is in $Sat(\Gamma)$, there is no need to use a NULL because is this case, no pair involving $x$ will appear in ${\sf C\_ans}(\mathcal{AQ})$. Consequently, if no $X$-value fits the conditions in the definition, ${\sf C\_ans}(\mathcal{AQ})$ is simply set to $\emptyset$.
Additionally to consistent answers as defined above, we argue that our approach allows for another kind of consistent answers to analytic queries, relying on the tuples in ${\sf Cons}(\mathcal{T})$, or by Theorem 4.1 on tuples true in every repair. To define this kind of consistent answer, a given analytic query $\mathcal{AQ}:~{}$select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] is associated with a query involving no aggregate, defined by
$Q_{\mathcal{AQ}}:$ select $\mathbf{K},$[$X$]$,M_{i}$ from $T$ where $\Gamma$.
Then, relying on the consistent answer to $Q_{\mathcal{AQ}}$, we define the strongly consistent answer to $\mathcal{AQ}$, denoted by ${\sf C\_ans}^{*}(\mathcal{AQ})$, as the answer to the following analytic query:
select [$X$]$,aggr(M_{i})$ from ${\sf C\_ans}(Q_{\mathcal{AQ}})$ [group-by $X$].
We emphasize that, based on Corollary 1, since the select clause of $Q_{\mathcal{AQ}}$ contains $\mathbf{K}$, ${\sf C\_ans}(Q_{\mathcal{AQ}})$ can be effectively computed, assuming that $m\_Chase(T)$ is available. The important issue of efficiently computing answers to analytic queries is addressed in the next section.
Example 7
In the context of Example 3, $\mathcal{AQ}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$) is an example of analytic query with no group-by clause and involving the aggregate $\max$. The expected result is the interval containing all possible sums of $M_{1}$-values among all tuples defined over $K_{1}K_{2}$ and associated with $A_{1}^{1}$-value $a_{1}$ in all repairs. Referring to Figure 3, we have:
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{1}]})=m_{1}+m^{\prime}_{1}+m_{1}$, and ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{2}]})=m_{1}$,
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{3}]})=m_{1}+m^{\prime\prime}_{1}+m_{1}$, and ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{4}]})=m_{1}$.
Assuming that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive numbers and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{1})=[glb,lub]$ and where $glb=m_{1}$ and $lub=2.m_{1}+m^{\prime\prime}_{1}$.
On the other hand, we have $Q_{\mathcal{AQ}_{1}}:~{}$select $\mathbf{K},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), and it can be seen from Figure 3, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{1}})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})=m_{1}$.
To illustrate that if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=m$, then it does not always hold that $glb\leq m\leq lub$, consider the aggregate $count(*)$ instead of $sum(M_{1})$. Indeed, in this case, ${\sf C\_ans}(\mathcal{AQ})=[3,3]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=1$.
As another example of analytic query with no group-by clause and slightly different from $\mathcal{AQ}_{1}$, consider $\mathcal{AQ}^{\prime}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a^{\prime}_{1}$). Referring to Figure 3, we have in this case:
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{1}^{\prime[R_{1}]})={\tt NULL}$, and ${\sf Ans}(\mathcal{AQ}_{1}^{\prime[R_{2}]})=m_{1}$,
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{1}^{\prime[R_{3}]})={\tt NULL}$, and ${\sf Ans}(\mathcal{AQ}_{1}^{\prime[R_{4}]})=m_{1}$.
The presence of NULL for $R_{1}$ and $R_{3}$ above is due to the fact that these repairs contain no tuples satisfying the selection condition in $\mathcal{AQ}^{\prime}_{1}$. In this case, by Definition 5(A), the expected consistent answer is NULL.
The query $\mathcal{AQ}_{2}:{\tt select}$ $A_{2}^{1}$, $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$) group-by $A_{2}^{1}$ is an analytic query involving a group-by clause. Since in this example ${\sf True}(\mathcal{T})$ contains only one $A_{2}^{1}$-value, namely $b_{1}$, at most one pair $\langle b_{1},[glb,lub]\rangle$ is expected in ${\sf C\_ans}(\mathcal{AQ}_{2})$. Referring again to Figure 3, we have:
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{2}^{[R_{1}]})={\sf Ans}(\mathcal{AQ}_{2}^{[R_{2}]})=\{(b_{1},m_{1}+m^{\prime}_{1})\}$, and
$\bullet$ ${\sf Ans}(\mathcal{AQ}_{2}^{[R_{3}]})={\sf Ans}(\mathcal{AQ}_{2}^{[R_{4}]})=\{(b_{1},m_{1}+m^{\prime\prime}_{1})\}$.
Assuming as above that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{2})=\{\langle b_{1},[glb,lub]\rangle\}$ where $glb=m_{1}+m^{\prime}_{1}$ and $lub=m_{1}+m^{\prime\prime}_{1}$.
On the other hand, we have $Q_{\mathcal{AQ}_{2}}:~{}$select $K,A_{1}^{2},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$). It can be seen from Figure 3, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{2}})=\{k_{1}k_{2}b_{1}m_{1}\}$. Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{2})=\{(b_{1},m_{1})\}$.
This example illustrates that if $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$ and $x\,m$ is in ${\sf C\_ans}^{*}(\mathcal{AQ})$, then it does not always hold that $glb\leq m\leq lub$. Indeed, for $m_{1}=1$, $m^{\prime}_{1}=2$ and $m^{\prime\prime}_{1}=3$, we have $glb=3$, $lub=4$ showing that for $x=b_{1}$, $m\not\in[3,4]$. $\Box$
6 Computing Consistent Answers to Analytic Queries
In this section, we give algorithms for computing consistent answers to analytic queries, first when no group-by clause is present, then when a group-by clause is present. We also consider in this section the case of analytic queries involving a group-by-having clause, whereas the introduction of the clause distinct is the subject of the last sub-section. We emphasize again that the selection conditions in queries are independent and involve no key attributes.
6.1 Basic Result
Before presenting our algorithms, we state important properties of repairs and analytic queries. To this end, we introduce the following additional notation. Given a star-table $T$ and an analytic query $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$], let:
$\bullet$ $\Sigma(\mathcal{AQ})=\{\sigma\in m\_Chase(T)~{}|~{}(\mathbf{K}\cup M_{i}\cup sch(\Gamma)\subseteq sch(\sigma))\wedge(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset)\}$
$\bullet$ $\Sigma^{+}(\mathcal{AQ})=\{\sigma\in\Sigma(\mathcal{AQ})~{}|~{}\sigma(sch(\Gamma))\subseteq Sat(\Gamma)\}$.
Proposition 9
Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] an analytic query where $\Gamma$ is independent, i.e., $\Gamma=\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ where $sch(\Gamma)=A_{1}\ldots A_{k}$. If $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that
  1. 1.
    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1})$ such that $t.sch(\Gamma)\not\in Sat(\Gamma)$,
  2. 2.
    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2})$ such that $t.sch(\Gamma)\in Sat(\Gamma)$.
Proof. The proof relies on the fact that for every repair $R$ of $T$ and every $\sigma$ in $\Sigma(\mathcal{AQ})$, ${\sf True}(\mathcal{R})$ contains exactly one tuple $t$ such that $t\in{\sf tuples}(\sigma)$ and $t.K\in\sigma(\mathbf{K})$ (the existence is a consequence of Proposition 4, unicity follows from Proposition 2(1) because $|\sigma(\mathbf{K})|=1$, $\mathbf{K}$ is a key of $R$ and $R\models FD$). As a consequence of Proposition 6, the existence of repairs $R_{1}$ and $R_{2}$ is explicitly shown using the process (P).
1. Regarding $R_{1}$, we build $\varphi$ as follows: For every $\sigma$ in $\Sigma(\mathcal{AQ})$:
Step P1. For every $A_{p}^{k}$ in $sch^{*}(D_{p})\cap sch(\sigma)$ such that $R_{1}$ contains no tuple $k_{p}a_{p}^{k}$ from ${\sf tuples}(\sigma(K_{p}A_{p}^{k}))$
$(a)$ if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then choose an $A_{p}^{k}$-value in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\setminus Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. If ${\sf tuples}(\sigma(A_{p}^{k}))\subseteq Sat(\Gamma(A_{k}^{p}))$, choose any $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. It is important to notice that in this case, there always exists at least one attribute $A_{p}^{k}$ such that ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, because otherwise $\sigma$ would be in $\Sigma^{+}(\mathcal{AQ})$.
Step P2. If $M_{i}\not\in sch(\Gamma)$ or if $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma^{+}(\mathcal{AQ})$, choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k_{p})=m$. Otherwise (i.e., $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$), choose an $M_{i}$-value $m$ in $\sigma(M_{i})\setminus Sat(\Gamma(M_{i}))$ for $\varphi_{i}(K_{p})$.
Once all m-tuples in $\Sigma(\mathcal{AQ})$ have been considered, the two steps above are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values as done in the generic description of process (P). Consequently, the corresponding set $\varphi$ in process (P) is properly defined.
Step P3. Build up the tuples $t_{\varphi}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. We notice that
$(a)$ if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\in Sat(\Gamma)$,
$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\not\in Sat(\Gamma)$.
By Proposition 6, we therefore obtain a repair $R_{1}$ satisfying item 1 in the proposition.
2. Regarding the item 2, the repair $R_{2}$ is obtained as above, by changing $(b)$ in Step P1 and Step P2 as follows:
Step P1. $(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\cap Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$. It is important to notice that in this case, every attribute $A_{p}^{k}$ in $sch(\Gamma)$ is such that ${\sf tuples}(\sigma(A_{p}^{k}))\cap Sat(\Gamma(A_{k}^{p}))\neq\emptyset$, because otherwise ${\sf tuples}(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ would not be possible.
Step P2. Choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k)=m$ where $k=\sigma(\mathbf{K})$.
Then, as above, these two steps are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values, and Step P3 is considered. By Proposition 6, we obtain a repair $R_{2}$ that satisfies item 2 in the proposition. The proof is therefore complete. $\Box$
The following example shows that Proposition 9 does not always hold for non independent selection conditions.
Example 8
Let $U=\{K_{1},K_{2},A_{1},A_{2},M_{1}\}$ and $FD=\{K_{1}\to A_{1},K_{2}\to A_{2},K_{1}K_{2}\to M_{1}\}$, and the following tables
$D_{1}$ $K_{1}$ $A_{1}$ $k_{1}$ $10$ $k^{\prime}_{1}$ $-5$ $D_{2}$ $K_{2}$ $A_{2}$ $k_{2}$ $20$ $k^{\prime}_{2}$ $-1$ $k^{\prime}_{2}$ $30$ $F$ $K_{1}$ $K_{2}$ $M_{1}$ $k_{1}$ $k_{2}$ $-10$ $k_{1}$ $k^{\prime}_{2}$ $2$ $k^{\prime}1$ $k^{\prime}_{2}$ $-100$
Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
[…]
and $T$ has two repairs $R_{1}$ and $R_{2}$. Moreover, the tuples in ${\sf True}(\mathcal{R}_{i})$ ($i=1,2$) whose schemas contain $\mathbf{K}=K_{1}K_{2}$, respectively denoted $R^{\mathbf{K}}_{1}$ and $R^{\mathbf{K}}_{2}$, are:
$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,-1,2),(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)\}$
$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,30,2),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$.
For $\Gamma=(M_{1}\leq A_{1})\wedge(M_{1}\geq 0\Rightarrow A_{2}<0)\wedge(M_{1}<-50\Rightarrow A_{2}>20)$ and $\mathcal{AQ}$ an analytic query involving $\Gamma$, we have $\Sigma(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ and $\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1}\}$. More precisely:
  1. 1.
    ${\sf tuples}(\sigma_{1})=\{(k_{1},k_{2},10,20,-10)\}$ where $(10,20,-10)\in Sat(\Gamma)$. Thus, ${\sf tuples}(\sigma_{1})\subseteq Sat(\Gamma)$ holds.
  2. 2.
    ${\sf tuples}(\sigma_{2})=\{(k_{1},k^{\prime}_{2},10,-1,2),(k_{1},k^{\prime}_{2},10,30,2)\}$, where $(10,-1,2)\in Sat(\Gamma)$ and $(10,30,2)\not\in Sat(\Gamma)$. Thus ${\sf tuples}(\sigma_{2})\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{2})\subseteq Sat(\Gamma)$.
  3. 3.
    ${\sf tuples}(\sigma_{3})=\{(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$, where $(-5,30,-100)\in Sat(\Gamma)$ and $(-5,-1,-100)\not\in Sat(\Gamma)$. Thus ${\sf tuples}(\sigma_{3})\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{3})\subseteq Sat(\Gamma)$.
On the other hand, it should be clear that $\Gamma$ is not independent, and that neither ${\sf True}(\mathcal{R}_{1})$ nor ${\sf True}(\mathcal{R}_{2})$ contains the two tuples $(k_{1},k^{\prime}_{2},10,30,2)$ and $(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)$, simply because they form a set that does not satisfy $K_{2}\to A_{2}$. Thus Proposition 9(1) does not hold in this example. Similarly, Proposition 9(2) is not satisfied either because neither ${\sf True}(\mathcal{R}_{1})$ nor ${\sf True}(\mathcal{R}_{2})$ contains the two tuples $(k_{1},k^{\prime}_{2},10,-1,2)$ and $(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)$, again because of $K_{2}\to A_{2}$ in $FD$. $\Box$
In the following two sections 6.2 and 6.3, we successively introduce our algorithms for computing consistent answers to analytic queries, depending on whether they involve or not a group-by clause. In each case, the main algorithm relies on a procedure, called Compute_aggregate and shown in Algorithm 2, whose role is to scan (either entirely or partially) the m-table $m\_Chase(T)$ and return values that will appear in the answers.
6.2 Analytic Queries with no Group-by Clause
If the query involves no group-by clause, as shown in Algorithm 3, the procedure Compute_aggregate of Algorithm 2 is called to scan the whole m-table $m\_Chase(T)$. When running this call, line 1 in Algorithm 3, the main loop lines 6-32 of Algorithm 2 scans $m\_Chase(T)$ and computes values appearing in the answers. The following proposition shows that Algorithm 3 is correct, except when $aggr=sum$. In other words, if $aggr\neq sum$ Algorithm 3 returns ${\sf C\_ans}(\mathcal{AQ})$. Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 3 in any case.
Algorithm 2 Procedure Compute_Aggregate
0:  An m-table $S\subseteq m\_Chase(T)$ and $\mathcal{AQ}:~{}$select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$]
0:
$proc\_change$ // a boolean variable
$proc\_min$ // meant to be equal to the lower bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
$proc\_max$ // meant to be equal to the upper bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
$proc\_ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$ if $proc\_change$ has value true
1:  if $aggr\neq count$ then
[…]
3:  if $aggr=count$ then
4:      $proc\_min:=0$ ; $proc\_max:=0$ ; $proc\_ans^{*}:=0$
5:  $proc\_change:={\tt false}$
6:  for all $\sigma$ in $S$ do
7:      if $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ then
8:           if $aggr\neq count$ then
9:               if $M_{i}\in sch(\Gamma)$ then
10:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
11:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
12:               else
13:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i})))$
14:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i})))$
15:           if ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$ then
16:               $proc\_change:={\tt true}$
17:               $proc\_min:=aggr(min_{\sigma},proc\_min)$
18:               $proc\_max:=aggr(max_{\sigma},proc\_max)$
19:               if ${\sf tuples}(\sigma(M_{i}))=\{m_{0}\}$ then
20:                    $proc\_ans^{*}:=aggr(proc\_ans^{*},m_{0})$
21:           else
22:               if $aggr=min$ then
23:                    $proc\_min:=\min(min_{\sigma},proc\_min)$
24:               if $aggr=max$ then
25:                    $proc\_max:=\max(max_{\sigma},proc\_max)$
26:               if $aggr=count$ then
27:                    $proc\_max:=proc\_max+1$
28:               if $aggr=sum$ then
29:                    if $min_{\sigma}<0$ then
30:                        $proc\_min:=proc\_min+min_{\sigma}$
31:                    if $max_{\sigma}>0$ then
32:                        $proc\_max:=proc\_max+max_{\sigma}$
33:  return  $proc\_change$, $proc\_min$, $proc\_max$, $proc\_ans^{*}$
Proposition 10
Let $T$ be a star-table over universe $U$ and let $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$ be an analytic query with no group-by clause. Then we have:
  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
    Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$.
  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. See Appendix C. $\Box$
Algorithm 3 Consistent answer to analytic queries with no group-by clause
0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$
0:
$[min\_ans,max\_ans]$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
$ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$
1:  Call procedure Compute_aggregate with input parameters $m\_Chase(T)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
2:  if $change\_min\_max={\tt true}$ then
3:      return  $([min\_ans,max\_ans]$, $ans^{*}$)
4:  else
5:      if $aggr\neq count$ then
6:           return  $({\tt NULL},{\tt NULL})$
7:      if $aggr=count$ then
8:           return  ($[0,0]$, 00)
In the following example we show that Algorithm 3 may fail to return exact values for $glb$ and $lub$ when the aggregate operator is $sum$, operating on positive values and negative values.
Example 9
As in Example 8, we consider $U=\{K_{1},K_{2},A_{1},A_{2},M_{1}\}$ and $FD=\{K_{1}\to A_{1},K_{2}\to A_{2},K_{1}K_{2}\to M_{1}\}$, but here, with the following tables
$D_{1}$ $K_{1}$ $A_{1}$
$k_{1}$ $10$
$k^{\prime}_{1}$ $-15$
$k^{\prime}_{1}$ $20$
$D_{2}$ $K_{2}$ $A_{2}$
$k_{2}$ $2$
$k^{\prime}_{2}$ 00
$k^{\prime}_{2}$ $3$
$F$ $K_{1}$ $K_{2}$ $M_{1}$
$k_{1}$ $k_{2}$ $30$
$k^{\prime}_{1}$ $k^{\prime}_{2}$ $-10$
$k_{1}$ $k^{\prime}_{2}$ $100$
Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
[…]
and $T$ has four repairs denoted by $R_{i}$ for $i=1,\ldots,4$, whose sets $R^{\mathbf{K}}_{i}$ of tuples over $U$ are defined by:
$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$
$R^{\mathbf{K}}_{3}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
$R^{\mathbf{K}}_{4}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$.
For $\Gamma=(A_{1}>0)\wedge(A_{2}>0)$ and $\mathcal{AQ}:$ select $aggr(M_{1})$ from $T$ where $\Gamma$, we have $\Sigma(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ and $\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1}\}$, because ${\sf tuples}(\sigma_{1}(A_{1}A_{2}))\subseteq Sat(\Gamma)$ holds, ${\sf tuples}(\sigma_{2}(A_{1}A_{2}))\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{2}(A_{1}A_{2}))\subseteq Sat(\Gamma)$, and ${\sf tuples}(\sigma_{3}(A_{1}A_{2}))\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{3}(A_{1}A_{2}))\subseteq Sat(\Gamma)$. On the other hand, $\Gamma$ is clearly independent.
  • If $aggr=min$ then, for every $i=1,\ldots,3$, ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=30$ and ${\sf Ans}(\mathcal{AQ}^{[R_{4}]})=-10$. Thus, ${\sf C\_ans}(\mathcal{AQ})=[-10,30]$.
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\min(\{30,\min\{-10,30\}\})=-10$ when processing $\sigma_{2}$, and then to $\min(\{-10,\min\{100,-10\}\})=-10$ when processing $\sigma_{3}$. Similarly, $max\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\max(\{30,\min\{-10,30\}\})=30$ when processing $\sigma_{2}$, and then to $\max(\{30,\min\{100,30\}\})=30$ when processing $\sigma_{3}$. Hence, Algorithm 3 returns $[-10,30]$ as expected.
  • Similarly, if $aggr=max$, we have ${\sf C\_ans}(\mathcal{AQ})=[30,100]$, which is also returned by Algorithm 3.
  • If $aggr=count$ then we have ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})=1$, ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})=2$, ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})=1$ and ${\sf Ans}(\mathcal{AQ}^{[R_{4}]})=3$. Thus, ${\sf C\_ans}(\mathcal{AQ})=[1,3]$.
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $1$ when processing $\sigma_{1}$ and then unchanged when processing $\sigma_{2}$ and $\sigma_{3}$. Moreover, $max\_ans$ is increased by $1$ for each m-tuple $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus Algorithm 3 returns $[1,3]$ as expected.
  • If $aggr=sum$ then we have ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})=30$, ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})=130$, ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})=30$ and ${\sf Ans}(\mathcal{AQ}^{[R_{4}]})=120$. Thus, ${\sf C\_ans}(\mathcal{AQ})=[30,130]$.
    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is successively set to $30$, $(30-10)=20$ when processing $\sigma_{1}$ and then $\sigma_{2}$, and unchanged when processing $\sigma_{3}$. On the other hand, $max\_ans$ is successively set to $30$ when processing $\sigma_{1}$, unchanged when processing $\sigma_{2}$ and set to $(30+100)=130$ when processing $\sigma_{3}$. Thus the call of the procedure Compute_aggregate in Algorithm 3 returns $[20,130]$.
  • On the other hand, assuming that the second tuple in $F$ is $(k^{\prime}_{1}k^{\prime}_{2},10)$ (instead of $(k^{\prime}_{1}k^{\prime}_{2},-10)$), then ${\sf C\_ans}(\mathcal{AQ})=[30,140]$, because ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})$ is $30$ if $i=1$ or $i=3$, $130$ if $i=2$ and $140$ if $i=4$. It can be seen that when running the call of the procedure Compute_aggregate in Algorithm 3, we obtain $min\_ans=30$ and $max\_ans=140$. $\Box$
Algorithm 4 Consistent answer to analytic queries with a group-by clause
0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$
0:
$Cons\_Ans$: a set of pairs $\langle x,[min\_ans,max\_ans]\rangle$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
$Ans^{*}$: a set of pairs $(x,ans^{*})$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$
1:  $Cons\_Ans:=\emptyset$ ; $Ans^{*}:=\emptyset$ ; $Temp:=\emptyset$
2:  for all $\sigma$ in $m\_Chase(T)$ do
3:      if $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ then
4:           if $(Y\to B$ in $FD$ and $YB\subseteq X\Rightarrow|{\sf tuples}(\sigma(B))|=1)$ and $({\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$then
5:               $Temp:=Temp\cup\{\sigma\}$
6:  for all $X$-value $x$ occurring in $Temp$ do
7:      $Temp(x):=\{\sigma\in Temp~{}|~{}x\in{\sf tuples}(\sigma(X))\}$
8:      Call procedure Compute_aggregate with input parameters $Temp(x)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
9:      if $change\_min\_max={\tt true}$ then
10:           $Cons\_Ans:=Cons\_Ans\cup\{\langle x,[min\_ans,max\_ans]\rangle\}$
11:           $Ans^{*}:=Ans^{*}\cup\{(x,ans^{*})\}$
12:  return  $(Cons\_Ans,Ans^{*}$)
6.3 Analytic Queries with a Group-by Clause
As in the case of analytic queries with no group-by clause, when the query involves a group-by clause, the computation of the consistent answers also involves a call of the procedure Compute_aggregate, as shown line 8 of Algorithm 4. This algorithm works as follows: a first scan of $m\_Chase(T)$, lines 2-5 retrieves in the set $Temp$ all relevant m-tuples $\sigma$, namely, such that $sch(\sigma)$ contains all attributes occurring in $\mathcal{AQ}$ and such that the $X$-values are consistent and may be associated with a tuple over $sch(\Gamma)$ satisfying the condition $\Gamma$. Then, a subsequent loop lines 6-11 operates a call of the procedure Compute_aggregate for each $X$-value $x$ occurring in the m-tuples of the set $Temp$. For each such call, Algorithm 2 scans the subset $Temp(x)$ of $Temp$ and computes the corresponding aggregate values appearing in the answers associated with $x$.
The following proposition shows that Algorithm 4 is correct, except when $aggr=sum$. In other words, if $aggr\neq sum$ Algorithm 4 returns ${\sf C\_ans}(\mathcal{AQ})$. Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 4 in any case.
Proposition 11
Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}$ an analytic query with a group-by clause.
  • If $aggr=min$, $aggr=max$ or $aggr=count$, then $Cons\_Ans$ as returned by Algorithm 4 is equal to ${\sf C\_ans}(\mathcal{AQ})$.
  • If $aggr=sum$, then $\langle x,[min\_ans,max\_ans]\rangle$ is in $Cons\_Ans$ as returned by Algorithm 4 if and only if ${\sf C\_ans}(\mathcal{AQ})$ contains $\langle x,[lub,glb]\rangle$ such that $min\_ans\leq glb$ and $max\_ans\geq lub$.
    Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then $Cons\_Ans={\sf C\_ans}(\mathcal{AQ})$.
  • For every aggregate function and every selection condition $\Gamma$, $Ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. First, the loop lines 2-5 scans $m\_Chase(T)$ and collects in the set $Temp$ the only m-tuples necessary to compute the consistent answer. Indeed, the collected m-tuples are all m-tuples $\sigma$ such that: they are defined over a super-set of $\mathbf{K}\cup sch(\sigma)$, their $X$-component contains consistent tuples, and their m-tuples over $sch(\Gamma)$ contain at least one tuple of $Sat(\Gamma)$. This is so because any m-tuple in $m\_Chase(T)$ not satisfying the above conditions can not contribute to the consistent answer. Then, in the next steps for every collected $X$-value, its associated aggregate value is evaluated as done in Algorithm 3. As a consequence of Proposition 10, for $x$ in $Temp$, the loop lines 6-11 in Algorithm 4 generates correct answers. The proof is therefore complete. $\Box$
We emphasize that Algorithm 3 and Algorithm 4 show that the consistent answers to a given analytic query $\mathcal{AQ}$, with or without a group-by clause, are computed in linear time in the size of $m\_Chase(T)$. Recalling that the computation of $m\_Chase(T)$ has been shown in LauS25 to be polynomial in time with respect to the size of $T$, it turns out that ${\sf C\_ans}(\mathcal{AQ})$ and ${\sf C\_ans}^{*}(\mathcal{AQ})$ can be computed in polynomial time with respect to the size of $T$.
6.4 Analytic Queries with a group-by-having Clause
Another important feature of our approach is that analytic queries with a group-by-having clause can be handled in our framework. This is so because, intuitively a having clause specifies through a boolean expression, which groupings should be in the answer. For example, in the context of Example 3, let $\mathcal{AQ}_{3}$ be the following query:
$\mathcal{AQ}_{3}:{\tt select}$ $A_{2}^{1}$, $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$) group by $A_{2}^{1}$ having ($max(M_{1})<10$)
In this case, $\mathcal{AQ}_{3}$ can be associated with the analytic query $\mathcal{AQ}^{\prime}_{3}$ defined by:
$\mathcal{AQ}^{\prime}_{3}:{\tt select}$ $A_{2}^{1}$, $sum(M_{1})$, $max(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$) group by $A_{2}^{1}$
whose consistent answer ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})$ is computed as described above. Then, considering the triples of the form $\langle a_{2},[glb_{1},lub_{1}],[glb_{2},lub_{2}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})$, all pairs $\langle a_{2},[glb_{1},lub_{1}]\rangle$ such that $lub_{2}<10$ are inserted in ${\sf C\_ans}(\mathcal{AQ}_{3})$. In other words, we have:
${\sf C\_ans}(\mathcal{AQ}_{3})=\{\langle a_{2},[glb_{1},lub_{1}]\rangle~{}|~{}\langle a_{2},[glb_{1},lub_{1}],[glb_{2},lub_{2}]\rangle\in{\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})\wedge lub_{2}<10\}$.
More precisely, applying Proposition 11 to the star-table $T$ in Figure 2, we obtain the following:
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=15$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,20],[5,15]\rangle\}$ and so, ${\sf C\_ans}(\mathcal{AQ}_{3})$ is empty. Notice incidentally that it is easily seen from Figure 3 that there exist repairs $R$ in ${\sf Rep}(T)$ for which ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})=\{\langle a_{2},20,15\rangle\}$, implying that ${\sf Ans}(\mathcal{AQ}_{3}^{[R]})=\emptyset$.
  • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=8$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,13],[5,8]\rangle\}$ and so, $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$. In this case, it is easily seen from Figure 3 that for every repair $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})$ contains a pair $\langle a_{2},\mu,\nu\rangle$ where $10\leq\mu\leq 13$ and $\nu<10$, implying that $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
In the light of this example, we argue that our approach can deal with analytic queries involving a having clause with condition $\Upsilon$, under the following restrictions:
  1. 1.
[…]
  2. 2.
    For every aggregate term in $\mathcal{AQ}$ of the form $sum(M_{i})$, then all $M_{i}$-values are positive.
Calling $aggr(M_{i})$ an aggregate term, the first item above implies that $\Upsilon$ can be written as $\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$ where for every $p=1,\ldots,h$, $\Upsilon_{p}$ is the conjunction of all atoms in $\Upsilon$ involving the same aggregate term. Moreover, for every aggregate term $\lambda_{p}$ occurring in $\Upsilon$, given a number $\alpha$, let $\Upsilon_{p}^{[\alpha]}$ be the expression obtained by substituting in $\Upsilon_{p}$ every occurrence of $\lambda_{p}$ with $\alpha$. Then, it turns out that the set $Sat(\Upsilon_{p})=\{\alpha~{}|~{}\Upsilon_{p}^{[\alpha]}$ evaluates to true}}\}} is an interval.
Based on this important remark, the following proposition characterizes the consistent answers of analytic queries involving a group-by-having clause.
Proposition 12
Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}$ an analytic query with a group-by-having clause defined by
$\mathcal{AQ}:{\tt select}~{}X$, $aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$
such that the above restrictions are met. If $\Upsilon=\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$, let $\mathcal{AQ}^{\prime}$ be the following analytic query with no having clause:
$\mathcal{AQ}^{\prime}:{\tt select}~{}X$, $aggr(M_{i})$, $aggr_{1}(M_{i_{1}})$, $\ldots$ , $aggr_{h}(M_{i_{h}})$ from $T$ where $\Gamma$ group by $X$
Then, $\langle x,[glb,lub]\rangle$ belongs to ${\sf C\_ans}(\mathcal{AQ})$ if and only if there exists $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$.
Proof. Let $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$. Given $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{p}\rangle$ such that $\mu\in[glb,lub]$ and $\mu_{p}\in[glb_{p},lub_{p}]$ for $p=1,\ldots,h$ hold. Hence, for $p=1,\ldots,h$, $\mu_{p}$ is in $Sat(\Upsilon_{p})$, implying that $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$.
Conversely, let $\langle x,[glb,lub]\rangle$ be in ${\sf C\_ans}(\mathcal{AQ})$ and let $p$ such that $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ does not hold. Since $Sat(\Upsilon_{p})$ is an interval, this implies that $glb_{p}\not\in Sat(\Upsilon_{p})$ or that $lub_{p}\not\in Sat(\Upsilon_{p})$. Since there exists a repair $R_{min}$, respectively a repair $R_{max}$, such that $aggr_{p}(M_{i_{p}})=glb_{p}$, respectively $aggr_{p}(M_{i_{p}})=lub_{p}$, it turns out that there exists at least one repair $R$ such that $ans(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{h}\rangle$ such that $\mu_{p}\not\in Sat(\Upsilon_{p})$. Thus $\langle x,\mu\rangle$ is not in $ans(\mathcal{AQ}^{[R]})$, which implies that $\langle x,[glb,lub]\rangle$ is not in ${\sf C\_ans}(\mathcal{AQ})$; a contradiction which completes the proof. $\Box$
It should be clear from Proposition 11 and Proposition 12 that, under the restrictions earlier stated, the consistent answer of an analytic query involving a group-by-having clause can be computed in polynomial time with respect to the size of $T$.
To end this section, we illustrate why in our approach, the sets $Sat(\Upsilon_{p})$ ($p=1,\ldots,h$) must be intervals. In the context of our earlier example, consider the query $\mathcal{AQ}_{3}$ with a having clause defined by $\Upsilon={\tt(}max(M_{1})<50~{}{\tt or}~{}max(M_{1})>100{\tt)}$, instead of $max(M_{1})<10$. Notice that in this case, the above restrictions are not satisfied, because $\Upsilon$ is not a conjunction, and $Sat(\Upsilon)$ is clearly not an interval.
Let $m_{1}=20$, $m^{\prime}_{1}=10$ and $m^{\prime\prime}_{1}=150$ and let $\langle a_{2},[30,170],[20,150]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})$, implying that the values of $max(M_{1})$ range between $20$ and $150$, when varying the repair. Although there are repairs in which the maximum is $20$ or $150$, thus satisfying the having condition, it is not possible to conclude that for every repair, the maximum $M_{1}$-value satisfies this condition: it could for example be that for a repair, $max(M_{1})$ is equal to $80$. It is thus impossible to certify that $\langle a_{2},[glb,lub]\rangle$ is in the consistent answer of $\mathcal{AQ}_{3}$.
6.5 Handling distinct Clauses
In this section, we show how to handle the distinct clause associated with the operator $count$. Recalling that $count({\tt distinct}\,M_{i})$ counts each $M_{i}$-value only once, independently from the number of its occurrences, we consider analytic queries $\mathcal{AQ}$ with no group-by clause, knowing that the presence of a group-by clause does not raise particular difficulties. Given such a query $\mathcal{AQ}:$ select $count({\tt distinct}~{}M_{i})$ from $T$ where $\Gamma$, we let ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$.
Similarly to the general case of the aggregate $sum$, an effective way of computing ${\sf C\_ans}(\mathcal{AQ})$ is currently unknown to the authors. This is so because, according to Algorithm 2, in order to determine $lub$ (respectively $glb$) without counting the same value more than once, every choice of one $M_{i}$-value in each set ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$ (respectively in $\Sigma^{+}(\mathcal{AQ})$) must be analyzed separately. As such a processing clearly requires to compute all possible choices, it is not polynomial, and so not tractable.
Consequently, instead of computing the exact values of $lub$ and $glb$, we propose a tractable approximation of ${\sf C\_ans}(\mathcal{AQ})$. To this end, we first notice that the values to be counted for determining $lub$ are obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma(\mathcal{AQ})$. Therefore, the set resulting from such a choice can not contain $(a)$ more distinct values than there are m-tuples in $\Sigma(\mathcal{AQ})$ and $(b)$ more distinct values than there are distinct values in the union of all sets ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$. This can be formally stated as $(a)$ $lub\leq|\Sigma(\mathcal{AQ})|$ and $(b)$ $lub\leq|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$, which we write as
$lub\leq\inf\left(|\Sigma(\mathcal{AQ})|,|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|\right)$.
If $count\_max$ denotes the infimum defined above, we emphasize that Algorithm 2 can be modified so as to compute $count\_max$ in polynomial time. Indeed, it is easy to see that $|\Sigma(\mathcal{AQ})|$ is equal to the value of $max\_ans$ returned by Algorithm 3 in the case of $count$. However, computing the second cardinality requires extra storage and extra computation to ensure that every value in the union is counted only once.
Regarding the approximation of $glb$, we recall from Algorithm 2 that the values to be counted for determining $glb$ are distinct values obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Thus, if $\Sigma^{+}(\mathcal{AQ})=\emptyset$, $glb=0$, and otherwise, a trivial lower bound for $glb$ is $1$, because all sets ${\sf tuples}(\sigma(M_{i}))$ for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, are nonempty.
To find a more accurate value of this lower bound, we notice that any choice of one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, concerns at least as many distinct $M_{i}$-values as there are in the union of all ${\sf tuples}(\sigma(M_{i}))$ of cardinality $1$, for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$. Denoting by $\Sigma^{+}_{1}(\mathcal{AQ})$ the set of these m-tuples, that is $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma\in\Sigma^{+}(\mathcal{AQ})~{}|~{}|{\sf tuples}(\sigma(M_{i}))|=1\}$, whenever $\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, we have:
$glb\geq|\bigcup_{\sigma\in\Sigma^{+}_{1}(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$.
Denoting by $count\_min$ the cardinality shown above, we notice that, as for $count\_max$, Algorithm 3 can be modified so as to compute $count\_min$ in polynomial time.
Another important remark regarding distinct is that if for every $\sigma$ in $m\_Chase(T)$, we have $|{\sf tuples}(\sigma(M_{i}))|=1$ (i.e., the dependency $\mathbf{K}\to M_{i}$ is satisfied), then $\Sigma^{+}_{1}(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$, and:
$\bullet$ $count\_max=lub=|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$     $(1)$
$\bullet$ $count\_min=glb=|\bigcup_{\sigma\in\Sigma^{+}(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$     $(2)$
The above remark applies to the query $\mathcal{AQ}$ of Example 9 if $aggr=count({\tt distinct}\,M_{1})$, because ${\sf tuples}(\sigma_{i}(M_{1}))$ are singletons for every $i=1,2,3$. More precisely, $|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|=|\Sigma(\mathcal{AQ})|=3$ (because the values involved are pairwise distinct), and $|\Sigma^{+}(\mathcal{AQ})|=1$, showing that Algorithm 3 would return the exact value of ${\sf C\_ans}(\mathcal{AQ})$, that is the interval $[1,3]$.
Example 10
In the context of Example 3, let
$\mathcal{AQ}:$ select $count({\tt distinct}\,M_{1})$ from $T$
be an analytic query involving a distinct clause, but no selection condition. In this case, we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$ and it can be seen from Figure 2 that the only m-tuples $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ are the first three m-tuples in $m\_Chase(T)$ that we denote by $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$. Thus, $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$ are equal to $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ that are such that ${\sf tuples}(\sigma_{1}(M_{1}))={\sf tuples}(\sigma_{3}(M_{1}))=\{m_{1}\}$ and ${\sf tuples}(\sigma_{2}(M_{1}))=\{m^{\prime}_{1},m^{\prime\prime}_{1}\}$.
Referring to the four repairs of $m\_Chase(T)$ shown in Figure 3, we have ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=2$ for $i=1,\ldots,4$, implying that ${\sf C\_ans}(\mathcal{AQ})=2$. On the other hand, applying the approximations shown above, we have $|\Sigma(\mathcal{AQ})|=|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=3$ and thus, $count\_max=3$. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain $count\_min=1$, producing the interval $[1,3]$, which approximates, but is not equal to ${\sf C\_ans}(\mathcal{AQ})$.
Changing $m^{\prime}_{1}$ to $m_{1}$ in $\sigma_{2}$ yields that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})={\sf Ans}(\mathcal{AQ}^{[R_{2}]})=1$ and ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})={\sf Ans}(\mathcal{AQ}^{[R_{4}]})=2$. Thus ${\sf C\_ans}(\mathcal{AQ})=[1,2]$. On the other hand, applying the approximations as above, we have $|\Sigma(\mathcal{AQ})|=3$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=2$ and so, $count\_max=2$. Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$. Hence, we obtain again $count\_min=1$, producing the interval $[1,2]$, which is now equal to ${\sf C\_ans}(\mathcal{AQ})$. $\Box$
7 Related Work
The problem of consistent query answering in inconsistent databases has motivated considerable research efforts since its introduction in ArenasBC99 . These efforts were first focused on conjunctive and self-join free queries involving no aggregate operators. The case of queries with aggregate operators and possibly with group-by clauses has been introduced in ArenasBCHRS03 and then further investigated in FuxmanFM05 and in KhalfiouiW23 ; KhalfiouiW24 . Focusing on relations whose constraints are unique key-constraints, these approaches rely on earlier work (see Wijsen19 ) characterizing classes of queries that can be rewritten in polynomial time, into SQL queries whose evaluation yields the consistent answer as defined in ArenasBCHRS03 .
In parallel, the issue of consistent conjunctive query answering to queries involving aggregate operators has been addressed in DixitK19 , based on a radically different technique, mainly that of reducing the problem to the well-known SAT problem. Here again, the constraints of interest are key-constraints associated with relations, as in FuxmanFM05 . Comparing our approach with the work in DixitK19 or in FuxmanFM05 , the main differences are the following:
  1. 1.
    The functional dependencies in DixitK19 or in FuxmanFM05 do not exactly match those usually considered in star schemas. Indeed, the dependencies $\mathbf{K}\to M_{i}$ are not in the scope of these approaches, contrary to the present approach.
  2. 2.
    In our approach, we allow missing values in the input tables, which is not the case in DixitK19 or in FuxmanFM05 .
  3. 3.
    Whereas in these approaches conjunctive queries are considered, in our approach, we allow specific disjunctions that we call independent selection conditions. We recall that independent selection conditions generalize conjunctive selection conditions.
  4. 4.
    Dealing with aggregate operators $sum$ and $count({\tt distinct}\,.)$ in our approach may lead to approximate query answers, which is not the case in DixitK19 and in FuxmanFM05 .
Therefore, it turns out that the two approaches are hardly comparable. However, we argue in this section that all results stated in the previous sections also hold when the functional dependencies $\mathbf{K}\to M_{i}$ are not considered. Moreover, in this case, when the aggregate is $count({\tt distinct}\,.)$ on a measure attribute, the consistent answer can be effectively evaluated, instead of being approximated as earlier explained. More precisely, we show that our approach works in the context of DixitK19 or of FuxmanFM05 , assuming that the key of the fact table $F$ consists of all its attributes (namely all key attributes along with all measure attributes).
Indeed, if $FD$ contains no functional dependency of the form $\mathbf{K}\to M_{i}$ then for every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is a singleton. Immediate consequences are that, in this context:
  1. 1.
    $FD$ is normalized and acyclic, and so, the results stated in LauS25 hold.
  2. 2.
    Proposition 2(2) is not true: m-tuples in $m\_Chase(T)$ may be distinct and have the same $\mathbf{K}$-value.
  3. 3.
    Step P2 in process (P). is irrelevant and so, should be ignored. Moreover, the second item of Step P3 should be removed because for every $\sigma$ such that $\mathbf{K}M_{i}\subseteq sch(\sigma)$ for $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is reduced to one tuple, and this tuple is consistent. However, it should be emphasized that Proposition 6 still holds in this setting because adapting the proof amounts to discard all arguments related to inconsistencies of tuples involving $\mathbf{K}$-values associated with measure values.
  4. 4.
    The second item about the characterization of tuples in ${\sf Confl}_{\min}(\mathcal{T})$ that follows Proposition 1 should be ignored because $|{\sf tuples}(\sigma(M_{i})|>1$ cannot happen. On the other hand, it turns out that no tuple in ${\sf Confl}_{\min}(\mathcal{T})$ can involve a measure value, because it is easy to see that if $t=t^{\prime}m$ is in ${\sf Confl}(\mathcal{T})$ and $m$ is in $adom(M_{i})$, then $t^{\prime}$ is in ${\sf Confl}(\mathcal{T})$.
  5. 5.
    Despite the first paragraph in the proof of Proposition 9 (mentioning that this proof relies on the fact that $m\_Chase(T)$ cannot contain distinct m-tuples with the same $\mathbf{K}$-value), the proposition still holds. Indeed, as Proposition 6 still applies, for any of the two items in Proposition 9, the modified process (P) generates a repair which satisfies the conditions in the proposition.
  6. 6.
    Dealing with aggregate operator $count({\tt distinct}\,.)$ in this context yields the exact consistent answer, instead of an approximation. This is so because in this case all $\sigma(M_{i})$ contain one value, in which case the equalities (1) and (2) in the previous sub-section hold.
An important consequence is that Algorithm 2, and thus Algorithm 3 and Algorithm 4, work in this context, showing that consistent answers to analytic queries can be computed in polynomial time, even when no functional dependency of the form $\mathbf{K}\to M_{i}$ is considered. It can thus be stated that our approach extends approaches from the literature because we allow missing values in the input tables and we can handle some disjunctive selection conditions (namely independent selection conditions involving no key attributes), at the cost of approximate consistent answers when the aggregate operator is $sum$, operating on positive or negative numbers.
8 Concluding Remarks
Data analytics requires the collection and integration of data coming from different sources and stored in what is called a data warehouse that is a database operating under a specific type of schema called a star schema. As a consequence, more often than not, such integrated data contain inconsistencies and missing values. In this paper, we have seen an approach to efficiently compute consistent answers to analytic queries in data warehouses. Extending earlier work concerning the consistent query answering of standard, non-analytic queries in multi-table databases, we presented in this paper an approach to computing exact consistent answers of analytic queries over star schemas under two assumptions: (a) the selection condition in the query involves no keys and (b) the selection condition is independent. Our main contributions are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema, and (b) showing that the exact consistent answer to an analytic query over a star schema can be computed also in time polynomial in the size of the data warehouse.
Our current work follows two main lines of research. First, the implementation of our approach is an important issue under consideration. We emphasize in this respect that since the functional dependencies in a star schema have a simple form, the generic algorithm $m\_Chase$ as presented here can be optimized. Moreover, as $m\_Chase(T)$ has to be ‘re-computed’ after each update, the question of incrementally implementing the changes in $m\_Chase(T)$ is crucial regarding efficiency. We are currently investigating these issues.
A second line of research is investigating the introduction of hierarchies over dimensions. Indeed, hierarchies play a central role in data warehouse querying, because they allow for data analyses at different ‘levels’ (for instance at the level of states instead of cities in a location dimension). From a practical point of view, introducing hierarchies means introducing new functional dependencies among dimensional attributes. So, if these hierarchies have inconsistencies and/or missing values, the results on the computation of consistent answers presented here have to be revisited. Moreover, the case of queries involving the operators roll-up or drill-down (that allow for ‘navigating’ within hierarchies) has to be investigated in this context.
Declarations
The two authors contributed to the study, conception and design. Both read and approved the manuscript.
No funds, grants, or other support was received for conducting this study.
Appendix A Proof of Proposition 6
Proposition 6  Let $T$ be a star-table over $U$. $R$ is a repair of $T$ if and only if there is a $\varphi$ as defined above such that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$.
Proof. We first notice that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{T})$ holds because all tuples in $R_{\varphi}$ are in ${\sf True}(\mathcal{T})$. To show that $R_{\varphi}\models FD$, we denote by $T_{\varphi}$ the table $\{t_{\varphi}(\sigma)~{}|~{}\sigma\in m\_Chase(T)\}$, and we recall that $R_{\varphi}=T_{\varphi}\cup{\sf Cons}(\mathcal{T})$. We first prove that $m\_Chase(T_{\varphi})=T_{\varphi}$.
Indeed, let $X\to A$ in $FD$, $t_{\varphi}(\sigma)$ and $t_{\varphi}(\sigma^{\prime})$ in $T_{\varphi}$ such that $t_{\varphi}(\sigma).X=t_{\varphi}(\sigma^{\prime}).X$. If $\sigma$ and $\sigma^{\prime}$ are defined over $A$, then, by construction of $T_{\varphi}$, we have $t_{\varphi}(\sigma).A=t_{\varphi}(\sigma^{\prime}).A$, in which case $m\_Chase$ does not change $T_{\varphi}$. The case whereby $\sigma$ is defined over $A$ and $\sigma^{\prime}$ is not defined over $A$ is not possible in $m\_Chase(T)$, and thus it is not possible that $t_{\varphi}(\sigma)$ is defined over $A$ while $t_{\varphi}(\sigma^{\prime})$ is not. Therefore, we have $m\_Chase(T_{\varphi})=T_{\varphi}$, and this implies that $T_{\varphi}\models FD$, because no conflicts can occur in $T_{\varphi}$.
Given an m-table $\Sigma$ over universe $U$, we denote by $\tau(\Sigma)$ the set of all tuples occurring in $\Sigma$. More formally: $\tau(\Sigma)=\{q\in\mathcal{T}~{}|~{}(\exists\sigma\in\Sigma)(\exists t\in{\sf tuples}(\sigma))(q\sqsubseteq t)\}$.
Recalling from LauS25 that, when $FD$ is acyclic, we also have $m\_Chase({\sf Cons}(\mathcal{T}))={\sf Cons}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})\models FD$. We prove by induction that, at each step $k$ of the computation of $m\_Chase(R_{\varphi})$ the obtained m-table $\Sigma^{k}$ is such that $(1)$ $\Sigma^{k}$ contains no conflict and $(2)$ $\tau(\Sigma^{k})=\tau(R_{\varphi})$.
$\bullet$ For $k=0$, i.e., $\Sigma^{0}=R_{\varphi}$, $(2)$ is obvious. As for $(1)$, assume that $R_{\varphi}$ contains $t$ and $t^{\prime}$ such that for $X\to A$ in $FD$ we have $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$. In this case, as $t$ and $t^{\prime}$ cannot be both in $T_{\varphi}$ or both in ${\sf Cons}(\mathcal{T})$, we consider that $t\in T_{\varphi}$ and that $t^{\prime}\in{\sf Cons}(\mathcal{T})$. As $t$ is in ${\sf True}(\mathcal{T})$, this implies that $t^{\prime}$ is in ${\sf Confl}(\mathcal{T})$, which is a contradiction. We therefore obtain that $t.A=t^{\prime}A$, and thus, $R_{\varphi}$ contains no conflicts.
$\bullet$ Assuming now that the result holds for some $k>0$, we prove that such is also the case for $k+1$. As $\Sigma^{k}$ contains no conflicts, its rows can be seen as tuples. So consider $t$ and $t^{\prime}$ in $\Sigma^{k}$ and $X\to A$ in $FD$ such that $t.X=t^{\prime}.X$ and $t$ and $t^{\prime}$ are defined over $A$. As $\Sigma^{k}$ contains no conflict, we have $t.A=t^{\prime}.A$ and so, $\Sigma^{k+1}=\Sigma^{k}$. Thus $(1)$ and $(2)$ are trivially satisfied.
Consider now $t$ and $t^{\prime}$ in $\Sigma^{k}$ and $X\to A$ in $FD$ such that $t.X=t^{\prime}.X$ and $t.A=a$ and $A\not\in sch(t^{\prime})$. In this case, the tuple $t^{\prime}$ is set to $t^{\prime}_{A}$ such that $sch(t^{\prime}_{A})=sch(t)\cup\{A\}$, $t^{\prime}_{A}.A=a$ and $t^{\prime}_{A}.sch(t)=t$. Thus contrary to the previous case, $\Sigma^{k+1}=\Sigma^{k}$ does not hold. We however notice that this transformation generates not conflicts, and thus that $(1)$ is satisfied by $\Sigma^{k+1}$. We now argue that $t^{\prime}_{A}$ is a tuple in $\tau(R_{\varphi})$, which combined with the inductive hypothesis that $\tau(\Sigma^{k})=\tau(R_{\varphi})$, implies that $\tau(\Sigma^{k+1})=\tau(R_{\varphi})$. As $t$ and $t^{\prime}$ are in $\Sigma^{k}$, they both belong to $\tau(R_{\varphi})$, and so there are $q$ and $q^{\prime}$ in $R_{\varphi}$ such that $t\sqsubseteq q$ and $t^{\prime}\sqsubseteq q^{\prime}$. To show that $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$, we successively consider the following cases:
-- If $q$ and $q^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$): as these sets are ‘closed’ under m_Chase, $t$ and $t^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$), and so $t^{\prime}_{A}$ is in $T_{\varphi}$ (respectively in ${\sf Cons}(\mathcal{T})$), because $t^{\prime}_{A}$ is obtained through the m-Chase process.
-- If $q$ is in ${\sf Cons}(\mathcal{T})$ and $q^{\prime}$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$: in this case $xa$ is in ${\sf Cons}(\mathcal{T})$ (because $xa\sqsubseteq q$). If $\sigma^{\prime}$ is the m-tuple in $m\_Chase(T)$ such that $q^{\prime}=t_{\varphi}(\sigma^{\prime})$, $XA\subseteq sch(\sigma^{\prime})$ and $xa$ is in ${\sf tuples}(\sigma^{\prime}(XA))$. Hence $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$, and thus $t^{\prime}_{A}$ is in ${\sf tuples}(R_{\varphi})$.
-- If $t$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$ and $t^{\prime}$ in ${\sf Cons}(\mathcal{T})$: in this case, assume first that $xa$ is in ${\sf Cons}(\mathcal{T})$. Then, $t^{\prime}_{A}$ is in ${\sf Cons}(\mathcal{T})$, because for every $Y\to B$ in $FD$ other than $X\to A$, if $YB\subseteq sch(t^{\prime}_{A})$ then $YB\subseteq sch(t^{\prime})$.
Assuming now that $xa$ is in ${\sf Confl}(\mathcal{T})$, we have $a=\varphi(x)$ and so, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime}_{A})$, $t^{\prime}_{A}.B=\varphi(t^{\prime}_{A}.Y)$ (because as $t^{\prime}$ is in ${\sf Cons}(\mathcal{T})$, we also have that for every $Y\to B$ such that $YB\subseteq sch(t^{\prime})$, $t^{\prime}.B=\varphi(t^{\prime}.Y)$). Hence, there exists $\sigma^{\prime}$ in $m\_Chase(T)$ such that $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$. Therefore $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$.
As a consequence, $\Sigma^{k+1}$ satisfies $(1)$ and $(2)$, and thus, so does $m\_Chase(R_{\varphi})$, meaning that ${\sf True}(R_{\varphi})=\tau(R_{\varphi})$ and $R_{\varphi}\models FD$. By Definition 2, in order to show that $R_{\varphi}$ is in ${\sf Rep}(T)$, we have to show that ${\sf True}(\mathcal{R}_{\varphi})$ is maximal.
To this end, let $S$ be such that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{S})\subseteq{\sf True}(\mathcal{T})$ and $S\models FD$, and we consider a tuple $q$ in ${\sf True}(\mathcal{S})\setminus{\sf True}(\mathcal{R}_{\varphi})$. Then, as $q$ is in ${\sf True}(\mathcal{T})$, there exist $\sigma$ in $m\_Chase(T)$ and $t$ in ${\sf tuples}(\sigma)$ such that $q\sqsubseteq t$. Since ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}_{\varphi})$, it follows that $q$ is in ${\sf Confl}(\mathcal{T})$, implying that there exists $X\to A$ in $FD$ such that $XA\subseteq sch(q)$ and ${\sf True}(\mathcal{T})$ contains $xa^{\prime}$ such that $x=q.X$ and $a^{\prime}\neq q.A$. By construction of $R_{\varphi}$, ${\sf True}(\mathcal{R}_{\varphi})$ contains a tuple $t^{\prime}$ from ${\sf tuples}(\sigma)$, and so, we have $t.X=q.X=t^{\prime}.X$ (because for every $X\to A\in FD$, $|{\sf tuples}(\sigma(X))|=1$) and $t.A\neq t^{\prime}.A$. As $t.A=q.A$ and $t^{\prime}$ is in ${\sf True}(\mathcal{S})$, this implies that $S\not\models FD$, which is a contradiction. This part of the proof is therefore complete.
Conversely, let $R$ be in ${\sf Rep}(T)$. By Proposition 5, for every $X\to A$ in $FD$, for a given $X$-value $x$ in ${\sf True}(\mathcal{T})$, ${\sf True}(\mathcal{R})$ contains one tuple among all tuples of the form $xa_{i}$ from ${\sf True}(\mathcal{T})$ where $a_{i}$ is in $adom(A)$. According to the steps of Process (P), these tuples define a repair, denoted by $R_{\varphi}$, and we show that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$. Since $R$ and $R_{\varphi}$ are in ${\sf Rep}(T)$, ${\sf Cons}(\mathcal{T})$ is a subset of ${\sf True}(\mathcal{R})$ and of ${\sf True}(\mathcal{R}_{\varphi})$. Now, let $t$ be in ${\sf Confl}(\mathcal{T})$. Then, for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $t.XA$ is in ${\sf True}(\mathcal{R})$ if and only if $t.XA$ is in ${\sf True}(\mathcal{R}_{\varphi})$ because $t.X$ is in ${\sf Cons}(\mathcal{T})$, associated with the same $A$-value in $R$ and $R_{\varphi}$. It follows that when $t$ is in ${\sf Confl}(\mathcal{T})$, $t$ is in ${\sf True}(\mathcal{R})$ if and only if $t$ is in ${\sf True}(\mathcal{R}_{\varphi})$. Hence, we obtain that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$, which completes the proof. $\Box$
Appendix B Proof of Proposition 8
Proposition 8  Let $T$ be a star-table over universe $U$, and $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ a query, such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition. Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
  1. 1.
    $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$,
  2. 2.
    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
  3. 3.
    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
Proof. Assume first that $\sigma$ in $m\_Chase(T)$ satisfies the items of the proposition. To show that $x$ is ${\sf C\_ans}(Q)$, we consider a repair $R$ in ${\sf Rep}(T)$ and we show that ${\sf True}(\mathcal{R})$ contains a tuple $x\gamma$ where $\gamma$ is in $Sat(\Gamma)$.
To this end, we first notice that by Proposition 1, the first two items imply that $x$ is in ${\sf Cons}(\mathcal{T})$ and thus, by Theorem 4.1, $x$ is in ${\sf True}(\mathcal{R})$. Moreover, since $\Gamma$ is independent, item 1 implies that for every $B$ in $sch(\Gamma)$, there exits $b$ in $\sigma(B)$ such that $b$ is in $Sat(\Gamma(B))$. If $S_{1}$ is the set of all $B$ in $sch(\Gamma)$ such that there exists $Y\to B$ in $FD$ where $YB\subseteq sch(Q)$, let $S_{2}=sch(\Gamma)\setminus S_{1}$ and $\gamma_{2}$ a tuple in ${\sf tuples}(\sigma(S_{2}))$ such that for every $B$ in $S_{2}$, $\gamma_{2}.B\in Sat(\Gamma(B))$ (such a tuple exists because by item 1, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$). We show that $x\gamma_{2}$, which belongs to ${\sf tuples}(\sigma(XS_{2}))$, is in ${\sf True}(\mathcal{R})$. Indeed, by definition of $S_{2}$, if $Y\to B$ is in $FD$ and such that $YB\subseteq XS_{2}$, then $B\not\in S_{2}$. Thus, as $\mathbf{K}\cap S_{2}=\emptyset$ (because $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $S_{2}\subseteq sch(\Gamma)$), $YB\subseteq XS_{2}$ holds if and only if $YB\subseteq X$. Consequently $x\gamma_{2}$ is in ${\sf Cons}(\mathcal{T})$ because so is $x$, showing by Theorem 4.1 that $x\gamma_{2}$ belongs to ${\sf True}(\mathcal{R})$.
Now let $B$ be in $S_{1}$ and $Y\to B$ the corresponding dependency in $FD$ such that $Y\subseteq X$. By Proposition 6, $R$ can be defined using process (P) by which a tuple $t_{\varphi}(\sigma)$ from ${\sf tuples}(\sigma)$ is set to be in ${\sf True}(\mathcal{R})$. Since $sch(t_{\varphi}(\sigma))=sch(\sigma)$ and $sch(Q)\subseteq sch(\sigma)$, we write $t_{\varphi}(\sigma)$ as $x_{\sigma}\gamma_{\sigma}q$ where $x_{\sigma}=t_{\varphi}(\sigma).X$, $\gamma_{\sigma}=t_{\varphi}(\sigma).sch(\Gamma)$, and $q=t_{\varphi}(\sigma).(sch(\sigma)\setminus sch(Q))$. In this setting, we have $t_{\varphi}(\sigma).Y=t.Y$ (because, by Proposition 2(1), $\sigma(Y)$ contains one tuple), and thus, by the Lossless-join rule (which applies here because $R\models FD$) applied to $t_{\varphi}(\sigma)$, $x\gamma_{2}$ and $Y\to B$, we obtain that $x\gamma_{2}b$ is in ${\sf True}(\mathcal{R})$. Reproducing this reasoning for every $B$ in $S_{1}$, we obtain a tuple $x\gamma_{2}\gamma_{1}$ in $sch(Q)$ such that for every $B$ in $S_{1}$, $x\gamma_{2}\gamma_{1}.B$ is in $\sigma(B)$ and thus, by item 3, in $Sat(\Gamma(B))$. It follows that $x\gamma_{1}\gamma_{2}$ is in ${\sf True}(\mathcal{R})$ and that $\gamma_{1}\gamma_{2}$ is in $Sat(\Gamma)$. Thus $x$ is in ${\sf Ans}(Q^{[R]})$, and since this reasoning holds for every repair $R$ of $T$, we obtain that $x$ is in ${\sf C\_ans}(Q)$.
Conversely, let $x$ be in ${\sf C\_ans}(Q)$. By Proposition 7, $x$ belongs to ${\sf C\_ans}^{+}(Q)$, meaning that there exists $t$ in ${\sf True}(\mathcal{T})$ such that $sch(Q)\subseteq sch(t)$, $t.X=x$, $x\in{\sf Cons}(\mathcal{T})$ and $t.sch(\Gamma)\in Sat(\Gamma)$. As for every $t\in{\sf True}(\mathcal{T})$, there exists $\sigma$ in $m\_Chase(T)$ such that $t$ is in ${\sf tuples}(\sigma(sch(t)))$, there must exist $\sigma$ in $m\_Chase(T)$ such that $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ and for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$ (because $x$ is in ${\sf Cons}(\mathcal{T})$). In other words, there exists $\sigma$ in $m\_Chase(T)$ satisfying the first two items of the proposition.
Denoting by $\Sigma(Q)$ the set of all these m-tuples, let $\sigma$ be in $\Sigma(Q)$ for which item 3 is not satisfied. Then, let $Y\to B$ in $FD$ such that $Y\subseteq X$, $B\in sch(\Gamma)$ and $b$ in ${\sf tuples}(\sigma(B))\setminus Sat(\Gamma(B))$. By item 1, ${\sf tuples}(\sigma(sch(Q)))$ contains $t$ such that $t.X=x$ and $t.B=b$. As $t$ is in ${\sf True}(\mathcal{T})$, by Proposition 4, there exists a repair $R$ such that $t$ is in ${\sf True}(\mathcal{R})$. Then, for every $x\gamma$ such that $\gamma$ is in $Sat(\Gamma)$, we have $t.Y=x.Y$ and $t.B\neq\gamma.B$ (since $\gamma.B\in Sat(\Gamma(B))$ and $t.B\not\in Sat(\Gamma(B))$). Consequently, $t$ and $x\gamma$ can not both belong to ${\sf True}(\mathcal{R})$ (because $R\models FD$). Therefore, if $\Sigma(Q)$ contains one m-tuple not satisfying item 3, then $x$ cannot belong to ${\sf C\_ans}(Q)$. We moreover notice that if one $\sigma$ in $\Sigma(Q)$ does not satisfy item 3 then for any other $\sigma^{\prime}$ in $\Sigma(Q)$, we have $\sigma^{\prime}(Y)=\sigma(Y)$ and thus $\sigma^{\prime}(B)=\sigma(B)$. It thus follows that $\sigma^{\prime}$ does not satisfy item 3 either. As a consequence, we obtain that if $x$ is in ${\sf C\_ans}(Q)$ then $m\_Chase(T)$ contains an m-tuple satisfying all three items in the proposition. The proof is therefore complete. $\Box$
Appendix C Proof of Proposition 10
Proposition 10  Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$ an analytic query with no group-by clause.
  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
    Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$.
  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
Proof. We separately consider two cases, depending on whether the test line 2 in Algorithm 3 succeeds or not. If the test fails, i.e., if $change\_min\_max$ as returned by the call of Compute_aggregate has value false, this means that $m\_Chase(T)$ contains no $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is empty. Thus, the test fails if and only if there exists a repair $R$ of $T$ for which ${\sf Ans}(Q^{[R]})=\emptyset$. In this case ${\sf C\_ans}(\mathcal{AQ})$ is NULL, which is expressed by the fact that the values of $min\_ans$, $max\_ans$ and $ans^{*}$ returned by the call of Compute_aggregate are respectively -dummy, +dummy and -dummy. Hence, in this case, Algorithm 3 provides the correct answer.
Suppose now that the test line 2 succeeds, i.e., that the value of $change\_min\_max$ returned by the call of Compute_aggregate is true. The statement line 16 of Algorithm 2 shows that $m\_Chase(T)$ contains at least one $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$. By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is not empty. Thus, the test succeeds if and only if for every repair $R$ of $T$, ${\sf Ans}(Q^{[R]})\neq\emptyset$, in which case, $min\_ans$, $max\_ans$ and $ans^{*}$ are proper values, either values from $adom(M_{i})$ if $aggr\neq count$, or positive integers if $aggr=count$. In this case, it can be seen that the m-tuples $\sigma$ in $m\_Chase(T)$ that contribute to the construction of the interval are such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$. This is so, because of our assumption on missing values in the fact table $F$ (every tuple with an $M_{i}$-value must be defined over all key attributes); and because any $\sigma$ not defined over some attributes in $X\cup sch(\Gamma)$ or such that ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)=\emptyset$ cannot contribute in the consistent answer to $\mathcal{AQ}$. This explains why, when the test line 7 of Algorithm 2 fails, no action is taken.
On the other hand, the m-tuples $\sigma$ such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ are precisely those in $\Sigma(\mathcal{AQ})$, and recalling that $\Sigma^{+}(\mathcal{AQ})$ is the set of m-tuples in $\Sigma(\mathcal{AQ})$ such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$, we have the following for every $\sigma$ in $\Sigma(\mathcal{AQ})$:
$(1)$ By Proposition 4, for every $t\in{\sf tuples}(\sigma)$, there exists a repair $R$ such that $t\in{\sf True}(\mathcal{R})$. Moreover, $t$ is the unique tuple over $sch(\sigma)$ in ${\sf True}(\mathcal{R})$ having $t.K$ as a $\mathbf{K}$-value. Notice that since $t.sch(\Gamma)$ may not be in $Sat(\Gamma)$, it is possible that $t.M_{i}$ does not contribute in the computation of the aggregate in $R$.
$(2)$ If $\sigma$ is in $\Sigma^{+}(\mathcal{AQ})$, by Corollary 1, every repair $R$ is such that ${\sf True}(\mathcal{R})$ contains a tuple $t$ in ${\sf tuples}(\sigma)$. In this case, since $t.sch(\Gamma)$ is in $Sat(\Gamma)$, $t.M_{i}$ does contribute in the computation of the aggregate in $R$.
Before showing that the values returned by Algorithm 3 are as stated in Definition 5, we mention that the aggregate operators $min$, $max$ and $sum$ are not defined when their argument is empty, which we write as $aggr(\emptyset)={\tt NULL}$. Otherwise, if $v_{1}$, $v_{2}$ and $v_{3}$ are values to which $aggr$ applies, then
Commutativity: $aggr(\{v_{1},v_{2}\})=aggr(\{v_{2},v_{1}\})$.
Associativity: $aggr(\{v_{1},aggr(\{v_{2},v_{3}\})\})=aggr(\{aggr(\{v_{1},v_{2}\},v_{3}\})=aggr(\{v_{1},v_{2},v_{3}\})$.
Monotonicity: If $aggr\neq count$ and $v_{2}\leq v_{3}$ then $aggr(\{v_{1},v_{2}\})\leq aggr(\{v_{1},v_{3}\})$.
The first two properties show that aggregate values do not depend on the order the elementary values are considered and how they are grouped during the computation. Moreover, the third property shows that, if $aggr\neq count$, the higher the values, the higher the aggregate values. In our context, recalling that elementary values are values in $\sigma(M_{i})$ for $\sigma$ in $\Sigma(\mathcal{AQ})$, this last property shows that, when the aggregate is different than $count$, for a fixed $\sigma$, the least, respectively the highest, aggregate value is obtained by considering the least, respectively the highest, possible $M_{i}$-value. These values, respectively denoted by $min_{\sigma}$ and $max_{\sigma}$ are computed lines 8-14 of Algorithm 2.
Since the second property shows that aggregates can be seen as operating over a set of values we recall the following standard properties that will be used in this proof. Let $S_{1}$ and $S_{2}$ be sets of values to which $aggr$ applies, and such that $S_{1}\subseteq S_{2}$, then: $\min(S_{1})\geq\min(S_{2})$, $\max(S_{1})\leq\max(S_{2})$, $count(S_{1})\geq count(S_{2})$, and if all values in $S_{1}$ or $S_{2}$ are positive, $sum(S_{1})\leq sum(S_{2})$. We notice that regarding the last property, if values in $S_{1}$ or $S_{2}$ can be positive or negative, no generic comparison can be stated.
We now explain, for each aggregate, how the bounds of the interval of the consistent answer can be obtained, and see why Algorithm 3 returns (or not) these values. First, as for every $\sigma\in\Sigma^{+}(\mathcal{AQ})$, every repair contains exactly one tuple in ${\sf tuples}(\sigma)$, the $M_{i}$-values in $\sigma(M_{i})$ if $M_{i}\not\in sch(\Gamma)$, or in $\sigma(M_{i})\cap Sat(sch(\Gamma))$ otherwise, contribute in the computation of the aggregate. Moreover, since by monotonicity, $min_{\sigma}$ contributes in the computation of $min\_ans$ and $max_{\sigma}$ contributes in the computation of $max\_ans$, we have the following, given that $\Sigma^{+}(\mathcal{AQ})\subseteq\Sigma(\mathcal{AQ})$:
  • For $aggr=min$, $min\_ans=\min\{min_{\sigma}~{}|~{}\sigma\in\Sigma(\mathcal{AQ})\}$ and $max\_ans=\min\{max_{\sigma}~{}|~{}\sigma\in\Sigma^{+}(\mathcal{AQ})\}$. These values are computed respectively in lines 23 and 17 of Algorithm 2.
  • For $aggr=max$, $min\_ans=\max\{min_{\sigma}~{}|~{}\sigma\in\Sigma^{+}(\mathcal{AQ})\}$ and $max\_ans=\max\{max_{\sigma}~{}|~{}\sigma\in\Sigma(\mathcal{AQ})\}$. These values are computed respectively in lines 17 and 25 of Algorithm 2.
  • For $aggr=count$, $min\_ans$, respectively $max\_ans$ is the cardinality of $\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed respectively in lines 17 and 27 of Algorithm 2.
  • For $aggr=sum$, $min\_ans$, respectively $max\_ans$, is the minimal, respectively maximal, sum that can be obtained by adding one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$ and then possibly one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$. These values are computed in lines 17 and 30, respectively in lines 18 and 32 of Algorithm 2. Notice that if all $M_{i}$-values are positive, knowing that adding 00 is neutral, then the test line 29 always fails and thus, $proc\_min$ is the sum of all minimal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$. Similarly, in this case, the test line 31 always succeeds, except for 00, and thus, $proc\_max$ is the sum of all maximal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma(\mathcal{AQ})$.
To complete the proof that the returned values $min\_ans$ and $max\_ans$ are as stated in the proposition, we have to show that ${\sf Rep}(T)$ contains repairs $R_{min}$ and $R_{max}$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{min}]})=min\_ans$ and ${\sf Ans}(\mathcal{AQ}^{[R_{max}]})=max\_ans$. These results are consequences of Proposition 9 by considering $R_{min}=R_{2}$ and $R_{max}=R_{1}$ if $aggr=min$, and $R_{min}=R_{1}$ and $R_{max}=R_{2}$ if $aggr$ is $max$ or $count$ or if $M_{i}$-values are positive and $aggr=sum$. We also mention that if $aggr=sum$ with no restriction on the $M_{i}$-values, $min\_ans$, respectively $max\_ans$, is the least, respectively the largest, sum that can be obtained using all relevant sets $\sigma(M_{i})$. We thus have $min\_ans\leq glb$ and $max\_ans\geq lub$, which concludes this part of the proof.
The last item in the proposition regarding ${\sf C\_ans}^{*}(\mathcal{AQ})$ and $ans^{*}$ is an immediate consequence of Corollary 1. The proof of the proposition is thus complete. $\Box$

### v2

@@
 Abstract
-We present an approach to computing consistent answers to analytic queries in data warehouses operating under a star schema and possibly containing missing values and inconsistent data.
-Our approach is based on earlier work concerning consistent query answering for standard, non-analytic queries in multi-table databases.
-In that work we presented polynomial algorithms for computing either the exact consistent answer to a standard, non analytic query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
-We extend this approach to computing exact consistent answers of analytic queries over star schemas, provided that the selection condition in the query involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute).
-The main contributions of this paper are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema under the above restrictions on the selection condition, and (b) showing that, under the same restrictions the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse.
-Keywords: Star schema $\bullet$ analytic query $\bullet$ inconsistent data $\bullet$ consistent query answering
+We present an approach to computing consistent answers to queries possibly involving an aggregation operator in databases operating under a star schema and possibly containing missing values and inconsistent data.
+Our approach is based on earlier work concerning consistent query answering for standard queries (with no aggregate operator) in multi-table databases.
+In that work, we presented polynomial algorithms for computing either the exact consistent answer to a query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
+In the present work, we consider databases operating under a star schema.
+Calling data warehouses such databases, we extend our previous work to queries involving aggregate operators, called analytic queries.
+In this context, we propose specific algorithms for computing exact consistent answers to queries, whether analytic or not, provided that the selection condition in the query satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute).
+We show that the overall time complexity of these specific algorithms is in $O(W\cdot\log(W))$, where $W$ is the size of the data warehouse.
+Moreover, the case of analytic queries involving a having clause associated with a group-by clause is discussed in the context of our approach.
+Keywords: Star schema, analytic query, inconsistent data, consistent query answering
 1 Introduction
-The efficient processing of analytic queries is an important issue in databases and has motivated considerable research work since the last three decades.
+The efficient processing of queries involving an aggregation operator, hereafter referred to as analytic queries, is an important issue in databases and has motivated considerable research work since the last three decades.
 The main purpose of analytic queries is to extract relevant ‘statistics’ from huge volumes of data, resulting from the integration of heterogeneous databases and stored in what is called a data warehouse inmon96 ; Kimball .
@@
 A star schema consists of two types of relation schemas (also called tables): a fact table and a number of dimension tables.
-In this context, an analytic query can be seen as an SQL Group-by query involving some aggregate function such as $min$, $max$, $count$ or $sum$ operating over attributes in the fact table called measure attributes (or simply measures).
+In this context, an analytic query can be seen as an SQL Group-by query involving some aggregate operator such as minmin, maxmax, countcount or sumsum operating over attributes in the fact table called measure attributes (or simply measures).
 Let us see an example to illustrate the concepts of star schema and analytic query.
@@
   • Dimension tables. There are four dimension tables, one for each of the attributes $K_{B}$, $K_{P}$, $K_{C}$, $K_{D}$:
-    • Branch, defined over the attributes $K_{B}$, $B\_Loc$, $B\_Ctry$ standing respectively for the branch identifier, the town in which the branch is located and the country which this town belongs to.
-The attribute $K_{B}$ is the (primary) key of Branch, meaning that the table Branch must satisfy the functional dependencies $K_{B}\to B\_Loc$ and $K_{B}\to B\_Ctry$.
-    • Prod, defined over the attributes $K_{P}$, $P\_Type$, $Price$, where $K_{P}$ is the product identifier, and $P\_Type$ and $Price$ are respectively the type and the price of a product.
+    • Branch, defined over the attributes $K_{B}$, B_TownB\_Town, B_CtryB\_Ctry standing respectively for the branch identifier, the town in which the branch is located and the country which this town belongs to.
+The attribute $K_{B}$ is the (primary) key of Branch, meaning that the table Branch must satisfy the functional dependencies $K_{B}\to B\_Town$ and $K_{B}\to B\_Ctry$.
+    • Prod, defined over the attributes $K_{P}$, P_TypeP\_Type, PricePrice, where $K_{P}$ is the product identifier, and P_TypeP\_Type and $Price$ are respectively the type and the price of a product.
 The attribute $K_{P}$ is the (primary) key of Prod, meaning that the table Prod must satisfy the functional dependencies $K_{P}\to P\_Type$ and $K_{B}\to Price$.
-    • Cust, defined over the attributes $K_{C}$, $C\_Name$, $C\_Addr$ standing respectively for customer identifier, name and address of the customer.
+    • Cust, defined over the attributes $K_{C}$, C_NameC\_Name, C_AddrC\_Addr standing respectively for customer identifier, name and address of the customer.
 The attribute $K_{C}$ is the (primary) key of the table Cust, meaning that Cust must satisfy the functional dependencies $K_{C}\to C\_Name$ and $K_{C}\to C\_Addr$.
-    • Date, defined over the attributes $K_{D}$, $Month$, $Year$ standing respectively for the date identifier or key, the month and the year of the date.
+    • Date, defined over the attributes $K_{D}$, $Month$, YearYear standing respectively for the date identifier or key, the month and the year of the date.
 The attribute $K_{D}$ is the (primary) key of Date, meaning that Date must satisfy the functional dependencies $K_{D}\to Month$ and $K_{D}\to Year$.
@@
 In the case of our example these constraints are expressed through the following inclusions: $\pi_{K_{B}}(F)\subseteq\pi_{K_{B}}({\sf Branch})$, $\pi_{K_{P}}(F)\subseteq\pi_{K_{P}}({\sf Prod})$, $\pi_{K_{C}}(F)\subseteq\pi_{K_{C}}({\sf Cust})$ and $\pi_{K_{D}}(F)\subseteq\pi_{K_{D}}({\sf Date})$.
-In this setting a typical analytic query is to display the total quantity of all products sold in all branches during the year $2024$. This query can be expressed in SQL as follows:
+In this setting a typical analytic query is to display the total quantity of each product sold during the year $2024$. This query can be expressed in SQL as follows:
 select $K_{P}$, $sum(Qty)$ from $J$ where $Year=2024$ group by $K_{P}$
-Here $J$ denotes the (lossless) join of all dimension tables with the fact table $F$ (although the join can by simplified by involving only $F$, Prod and Date). $\Box$
+Here $J$ denotes the (lossless) join of all dimension tables with the fact table $F$ (although the join can be simplified by involving only $F$, Prod and Date). $\Box$
 How to efficiently evaluate analytic queries against huge volumes of data has been widely investigated and lies outside the scope of the present paper; the reader is referred to Ullman regarding standard SQL query optimization and to BellatrecheGL05 regarding more specific optimization techniques for analytic queries.
@@
 This is so because, when computing the answer to this query, the only needed attribute value among all attributes of the table Cust is the $K_{C}$-value of the tuple in $F$ being processed.
-A more problematic situation is if the selection condition in the query is $Year=2024$ and $C\_Addr=Paris$.
+A more problematic situation is if the selection condition in the query is $Year=2024$ and C_Addr=ParisC\_Addr=Paris.
 This is so because among all transactions regarding customers whose address may be Paris, some concern customers whose address may violate the dependency $K_{C}\to C\_Addr$ in the table Cust.
@@
 To cope with inconsistencies and missing values in data warehouses, our approach is based on our earlier work LauS25 dealing with consistent query answering for standard, non analytic queries in multi-table databases.
-In that work, we presented polynomial algorithms for computing either the exact consistent answer to a standard non analytic query or bounds of the exact answer, depending on whether the query involves or not a selection condition.
-In the present paper, we show that in the case of a star schema, under the restrictions that the selection condition involves no keys and satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute), the exact consistent answer can be effectively computed.
-In the following section, we summarize briefly the main results of the approach in LauS25 and then focus on analytic queries over star schemas.
-Considering queries, analytic or not, whose selection condition satisfies the two restrictions mentioned above, the main contribution of this paper is showing that: (a) computing the exact consistent answer to a usual projection-selection-join query over a star schema is in time polynomial in the size of the data warehouse (in contrast to LauS25 , where consistent answers to non analytic queries are approximated when involving a selection condition), and (b) the exact consistent answer to an analytic query over a star schema can be computed in time polynomial in the size of the data warehouse (with two exceptions where only approximations are given).
+In that work, we presented polynomial algorithms for computing either the exact consistent answer to a standard non analytic query or bounds of the exact answer, depending on whether the query involves a selection condition or not.
+In the present paper, we show how to compute effectively the exact consistent answer in the case of a star schema, under the restriction that the selection condition satisfies the property of independency (i.e., the condition can be expressed as a conjunction of conditions each involving a single attribute).
+The main contributions of this paper can be summarized as follows: We propose specific algorithms for computing exact consistent answers to queries, whether analytic or not, over a star schema, provided that the selection condition in the query satisfies the property of independency.
+We also show that the overall time complexity of our specific algorithms is in $O(W\cdot\log(W))$, where $W$ is the size of the data warehouse.
+We discuss the case of analytic queries involving a having clause associated with a group-by clause in the context of our approach.
 The paper is organized as follows: In Section 2 we recall the main features of our previous work in LauS25 , on which the present approach is based.
 In Section 3 we first recall the definition of a star schema and argue that the approach in LauS25 applies in this context.
-In Section 4 we investigate further the concept of repairs in the context of star schemas.
+In Section 4 we investigate the concept of repairs in the context of star schemas.
 Section 5 deals with consistent answers to queries in the case of standard projection-selection queries as well as in the case of analytic queries.
-In Section 6, we propose algorithms for efficiently computing the consistent answers, or in some few cases, an approximation of the consistent answers to analytic queries.
+In Section 6, we propose algorithms for efficiently computing the consistent answers to analytic queries, or in some few cases an approximation to the consistent answer.
 In Section 7 we compare our approach to other approaches from the literature and in Section 8 we summarize the contents of the paper and suggest research directions for future work.
 2 The Approach of LauS25
-2.1 The Approach at a Glance
+2.1 Basic Definitions and Notation
+As done in JIIS ; LauS25 , we rely on basic notions from relational databases as introduced in the literature Ullman .
+As usual in this context, we consider a finite set of attributes $U=\{A_{1},\ldots,A_{n}\}$, called the universe, and we assume that each attribute $A_{i}$ is associated with a domain of values denoted $dom(A_{i})$.
+A nonempty subset $R=\{A_{i_{1}},\ldots,A_{i_{p}}\}$ of $U$ is called a (relation) schema and simply denoted by $A_{i_{1}}\ldots A_{i_{p}}$ (i.e., by the concatenation of its attributes).
+Given such a schema $R=A_{i_{1}}\ldots A_{i_{p}}$, a tuple $t$ over $R$ is an element of the cross product $dom(A_{i_{1}})\times\ldots\times dom(A_{i_{p}})$, and $t$ is denoted by the concatenation of its values, that is $t=a_{i_{1}}\ldots a_{i_{p}}$ where $a_{i_{j}}$ is in $dom(A_{i_{p}})$, for every $j=1,\ldots,p$.
+In this case, $R$ is called the schema of $t$ and is also denoted by $sch(t)$.
+Moreover, if $S$ is a subset of $R$, the restriction of $t$ over $S$, denoted by $t.S$ is the tuple over $S$ whose values over attributes in $S$ are those of $t$.
+More formally, given tuple $t$ over $R$, $S\subseteq R$, and tuple $q$ over $S$, the fact that $q=t.S$ means that $sch(q)=S$ and that for every $A\in S$, $q.A=t.A$.
+In this setting a relation, also called a table, over schema $R$ is a finite set of tuples over $R$, and a multi-table database over universe $U$ is a set of such relations.
+Additionally, relations in a database are associated with constraints, whose goal is to restrict their content so as to model given properties.
+Functional dependencies, that we consider in our work, are recognized as the most popular constraints.
+Given a relation $r$ over schema $R$, a functional dependency over $R$ (or simply ‘dependency over $R$’) is an expression of the form $X\to Y$ where $X$ and $Y$ are subsets of $R$.
+A relation $r$ over $R$ is said to satisfy the dependency $X\to Y$ if for all tuples $t$ and $t^{\prime}$ in $r$ the following holds: if $t.X=t^{\prime}.X$ then $t.Y=t^{\prime}.Y$; and $r$ is said to satisfy a set $FD$ of dependencies over $R$ if $r$ satisfies every dependency of $FD$.
+A basic result from functional dependency theory Armstrong says that: $r$ satisfies a set $FD$ of dependencies over $R$ if and only if $r$ satisfies the set $FD^{\prime}=\{X\to A~|~X\to Y\in FD,A\in Y\}$.
+For example, if $r$ is a relation over $R=ABCDE$ then $r$ satisfies $FD=\{AB\to CD,D\to EC\}$ if and only if $r$ satisfies $FD^{\prime}=\{AB\to C,AB\to D,D\to E,D\to C\}$.
+This explains why in our work we restrict our attention to dependencies having a right hand-side consisting of just one attribute.
+Next, we address the issue of checking whether a given multi-table database satisfies a given set of dependencies.
+2.2 From $Chase$ to m_Chasem\_Chase
 Traditionally, to verify the consistency of a multi-table database with respect to a set $FD$ of functional dependencies one applies the well-known Chase algorithm Ullman .
@@
 else if $t.A=a$ and $t^{\prime}.A$ is null then assign $a$ to $t^{\prime}.A$
-In this context, a tuple $t$ in the current value of $Chase(T)$ is said to be conflicting if the following holds: there is a tuple $t^{\prime}$ in the current value of $Chase(T)$ and a dependency $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are both defined over $XA$, $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$.
+When running the Chase algorithm, a tuple $t$ in the current value of $Chase(T)$ is said to be conflicting if the following holds: there is a tuple $t^{\prime}$ in the current value of $Chase(T)$ and a dependency $X\to A$ in $FD$ such that $t$ and $t^{\prime}$ are both defined over $XA$, $t.X=t^{\prime}.X$ and $t.A\neq t^{\prime}.A$.
 A tuple $t$ is called non-conflicting if $t$ is not a conflicting tuple.
 Now, if Chase is successful (i.e., no conflicts are encountered and no more new tuples can be derived) then the database is declared consistent else conflicting.
-If the database is consistent then processing of queries (whether standard queries or analytic queries) proceeds as usual, else the following question arises: can we still extract useful (i.e.
-non conflicting) information from the conflicting database?
-The work in LauS25 gives a positive answer to this question based on an extended version of the Chase algorithm, called the $m\_Chase$ algorithm to be presented formally in the following subsection.
-The input of this algorithm is no more $T$ but a table $\mathcal{T}$ containing all tuples that can be built up from constants in the active domains of the database.
-The set $\mathcal{T}$ is called the ‘Universe of discourse’ of the $m\_Chase$ algorithm and the notion of ‘conflicting tuple’ remains the same but now concerns all tuples of $\mathcal{T}$ and not only those of $T$.
-It is shown in LauS25 that the tuples of $\mathcal{T}$ can be characterized in two orthogonal ways: a tuple of $\mathcal{T}$ can be either true or false, and it can be either conflicting or non-conflicting.
-This characterization can be intuitively described as follows.
-If Chase terminates successfully and $Chase(T)$ denotes the output table then a tuple $t$ of $\mathcal{T}$ is true if it appears in $Chase(T)$ and false otherwise (i.e.
-$t$ is false if it appears in $\mathcal{T}\setminus Chase(T)$).
+If the database is consistent then processing of queries (whether standard queries or analytic queries) proceeds as usual, else the following question arises: can we still extract useful (i.e., non conflicting) information from the conflicting database?
+The work in LauS25 gives a positive answer to this question based on an extended version of the Chase algorithm, called the m_Chasem\_Chase algorithm (to be presented formally in the following subsection).
+As will shall see, the m_Chasem\_Chase algorithm operates over the set of all tuples that can be built up from constants from the active domains of the database, rather than over the tuples in $T$.
+We recall that the active domain of an attribute $A$ in a multi-table database is the set of all values of the domain of $A$ that are present in the database.
+Note that every tuple appearing in the database is built up from constants in the active domains of database attributes but there are tuples that are built-up from constants in the active domains that are not present in the database.
+For example, if a database relation has only two tuples, $ab$ and $a^{\prime}b^{\prime}$, then the tuples $ab^{\prime}$ and $a^{\prime}b$ are built up from constants in the active domains (here $\{a,a^{\prime}\}$ and $\{b,b^{\prime}\}$) but they don’t belong to the relation.
+In order to deal with this important difference between Chase and m_Chasem\_Chase, we introduce the following notation: given a set of tuples $S$, we denote by $\mathcal{S}$ the set of all tuples that can be built up from constants occurring in $S$.
+In our previous example, where $S=\{ab^{\prime},a^{\prime}b\}$, we have $\mathcal{S}=\{ab,ab^{\prime},a^{\prime}b,a^{\prime}b^{\prime},a,a^{\prime},b,b^{\prime}\}$.
+It is important to note that when considering the table $T$, the set $\mathcal{T}$ is precisely the set of all tuples that can be built up from constants in the active domains of attributes in $U$.
+Hereafter, we shall refer to $\mathcal{T}$ as the universe of discourse of the m_Chasem\_Chase.
+To summarize, the $Chase$ algorithm operates over the set $T$ of all tuples in the database, whereas the m_Chasem\_Chase algorithm operates over the set $\mathcal{T}$.
+Moreover, in the m_Chasem\_Chase algorithm the notion of ‘conflicting tuple’ remains the same as for the Chase algorithm, but now this notion concerns all tuples of $\mathcal{T}$, not just those of $T$.
+More precisely, it is shown in LauS25 that the tuples of $\mathcal{T}$ can be characterized in two orthogonal ways: a tuple of $\mathcal{T}$ can be either true or false, and it can be either conflicting or non-conflicting.
+This characterization can be intuitively described as follows:
+  • If Chase algorithm terminates successfully and $Chase(T)$ denotes the output table, then a tuple $t$ of $\mathcal{T}$ is true if it appears in $Chase(T)$ and false otherwise (i.e., $t$ is false if it appears in $\mathcal{T}\setminus Chase(T)$).
+In this case no tuple in $\mathcal{T}$ is conflicting, entailing that all tuples in $\mathcal{T}$ are non-conflicting.
   • However, if the Chase algorithm fails then we don’t know which tuples are true and which are non-conflicting.
-The $m\_Chase$ algorithm remedies this deficiency by modifying the Chase algorithm as follows: instead of stopping the application of functional dependencies on table $T$ when a conflict is encountered, the application continues (and the true tuples are stored) until no more tuples are found.
+The m_Chasem\_Chase algorithm remedies this deficiency by modifying the Chase algorithm as follows: instead of stopping the application of functional dependencies on table $T$ when a conflict is encountered, the application continues (and the true tuples are stored) until no more tuples are found.
 In doing so, all true tuples and all conflicting tuples are computed - and therefore each tuple of $\mathcal{T}$ can be characterized as true/false and as conflicting/non-conflicting.
@@
 Note that every sub-tuple of a true tuple is true and that every sub-tuple of a consistent tuple is consistent.
-Finally, call ‘consistent subset’ of $\mathcal{T}$ any set $S$ of true tuples of $\mathcal{T}$ such that the set of all tuples inferred from $S$ using the functional dependencies contains no conflicting tuples in $\mathcal{S}$.
+Then, a set $S$ of true tuples of $\mathcal{T}$ is said to be a consistent subset of $\mathcal{T}$ if the set of all tuples inferred from $S$ using the functional dependencies contains no conflicting tuples in $\mathcal{S}$.
 Let us illustrate these concepts using the following example.
 Example 2
-Suppose $T=\{ab,bc,ac^{\prime}\}$ and $FD=\{A\to C,B\to C\}$.
-Then all tuples in $T$ are true and non conflicting (hence consistent), but the application of the functional dependencies on $T$ allows to infer the tuples $abc$ and $abc^{\prime}$, which are conflicting tuples of $\mathcal{T}$ inferred from $T$.
-In fact it can be seen that $(a)$ the true tuples of $\mathcal{T}$ are $abc$, $abc^{\prime}$ and all their sub-tuples, implying that any other tuple of $\mathcal{T}$ is false in $\mathcal{T}$, and $(b)$ the conflicting tuples in $\mathcal{T}$ are $abc$, $abc^{\prime}$, $ac$, $ac^{\prime}$, $bc$ and $bc^{\prime}$, implying that any other tuple in $\mathcal{T}$ is non conflicting in $\mathcal{T}$.
-In this example, the consistent tuples of $\mathcal{T}$ are $ab$, $a$, $b$, $c$ and $c^{\prime}$.
-In this context, the set $R=\{abc,c^{\prime}\}$ is a consistent subset of $\mathcal{T}$.
-Indeed, since $\mathcal{R}$ is the set of all tuples built up using constants $a$, $b$, $c$ and $c^{\prime}$, we have $\mathcal{R}=\mathcal{T}$.
-Moreover the tuples in $R$ are tuples true in $\mathcal{T}$ and no conflicting tuples in $\mathcal{R}$ can be generated from the tuples in $R$.
-It is important to notice that, although $abc$ is conflicting in $\mathcal{T}$, $abc$ is not conflicting in $\mathcal{R}$ because $ac^{\prime}$, $bc^{\prime}$ and $abc^{\prime}$ are not true in $\mathcal{R}$.
+Consider a universe of three attributes $A$, $B$ and $C$ over which the functional dependencies $FD=\{A\to C,B\to C\}$ are assumed to hold, along with a database consisting of three relations, namely $r_{1}=\{ab\}$, $r_{2}=\{bc\}$ and $r_{3}=\{ac^{\prime}\}$, respectively defined over schemas $AB$, BCBC and $AC$.
+Since no dependency can be enforced on $r_{1}$, and since $r_{2}$ and $r_{3}$ respectively satisfy $B\to C$ and $A\to C$, the database satisfies the dependencies in $FD$.
+However, when ‘chasing’ this database, its tuples are first collected in a single table $T$ defined over the full universe $ABC$ and containing the three tuples $\{ab,bc,ac^{\prime}\}$, in which missing values occur (for example, $ab$ is not defined over attribute $C$).
+At this stage, all tuples in $T$ are true (since they are in the database) and they are non conflicting (since the dependencies in $FD$ are satisfied).
+However, the application of these functional dependencies on $T$ allows to infer the tuples abcabc and $abc^{\prime}$, which are conflicting tuples of $\mathcal{T}$ inferred from $T$.
+In fact it can be seen that $(a)$ the true tuples of $\mathcal{T}$ are abcabc, $abc^{\prime}$ and all their sub-tuples, implying that any other tuple of $\mathcal{T}$ is false in $\mathcal{T}$, and $(b)$ the conflicting tuples in $\mathcal{T}$ are abcabc, $abc^{\prime}$, $ac$, $ac^{\prime}$, $bc$ and $bc^{\prime}$, implying that any other tuple in $\mathcal{T}$ is non conflicting in $\mathcal{T}$.
+In this example, the consistent tuples of $\mathcal{T}$ (i.e., the tuples that are true and non conflicting in $\mathcal{T}$) are $ab$, $a$, $b$, $c$ and $c^{\prime}$.
+In this context, for $R=\{abc,c^{\prime}\}$, we denote by $\mathcal{R}$ the set of all tuples that can be built up using the constants occurring in $R$, namely $a$, $b$, $c$ and $c^{\prime}$.
+Thus, we have $\mathcal{R}=\mathcal{T}$ and it can be seen that applying the $Chase$ algorithm to the tuples in $R$ does not change $R$, and does not fail.
+Therefore, no tuples of $\mathcal{R}$ are conflicting, and since $R$ is a set of tuples true in $\mathcal{R}$, $R$ is said to be a consistent subset of $\mathcal{T}$.
+We notice that although abcabc is conflicting in $\mathcal{T}$, abcabc is not conflicting in $\mathcal{R}$ because $ac^{\prime}$, $bc^{\prime}$ and $abc^{\prime}$ are not true in $\mathcal{R}$.
 $\Box$
-Finally, based on the concepts introduced so far, a repair of $T$ in LauS25 is defined to be a maximal and consistent subset of $\mathcal{T}$ containing a maximal and consistent set of tuples which are consistent in $\mathcal{T}$.
+Based on the concepts introduced so far, a repair of $T$ in LauS25 is defined to be a maximal and consistent subset of $\mathcal{T}$ containing a maximal and consistent set of tuples which are consistent in $\mathcal{T}$.
 In our previous example, the subset $R=\{abc,c^{\prime}\}$ of $\mathcal{T}$ is a repair of $T$ because $(a)$ as we have just seen, $R$ is a consistent subset of $\mathcal{T}$, $(b)$ $R$ is maximal because adding to $R$ a tuple true of $\mathcal{T}$ either does not bring any new true tuple in $\mathcal{R}$ (e.g., adding the tuple $ac$) or generates a conflicting tuple in $\mathcal{R}$ (e.g., adding the tuple $ac^{\prime}$), and $(c)$ all consistent tuples of $\mathcal{T}$ are true in $\mathcal{R}$.
 Note that similar arguments show that the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of $\mathcal{T}$, however, $S$ is not a repair of $T$, because $ab$ is a consistent tuple of $\mathcal{T}$ which is not true in $\mathcal{S}$.
-By the way, as we shall see in Section 4.1, our definition of repair is more restrictive than the usual definition ArenasBC99 ; Wijsen19 in which a repair is defined to be a maximal and consistent subset of $\mathcal{T}$.
-For example, the set $S=\{bc,ac^{\prime}\}$ is a repair of $T$ following ArenasBC99 ; Wijsen19 , but it is not a repair of $T$ following our approach.
+By the way, as we shall see in Section 4.1, our definition of repair is more restrictive than the usual definition ArenasBC99 ; Wijsen19 in which a repair is defined to be a maximal and consistent subset of tuples true in $\mathcal{T}$.
+In the context of the previous example, since the set $S=\{bc,ac^{\prime}\}$ is a maximal and consistent subset of tuples true in $\mathcal{T}$, $S$ is seen as a repair of $T$ following ArenasBC99 ; Wijsen19 , whereas $S$ is not a repair of $T$ following our approach, because $ab$ is consistent in $\mathcal{T}$ but not true in $\mathcal{S}$.
 2.3 The m-Chase Algorithm
-Clearly to apply the $m\_Chase$-based approach described above, one has to answer the following questions:
-  • Does the $m\_Chase$ algorithm terminate?
+Clearly to apply the m_Chasem\_Chase-based approach described above, one has to answer the following questions:
+  • Does the m_Chasem\_Chase algorithm terminate?
   • Is the result independent of the order in which the functional dependencies are applied?
-  • Does the result contain all true tuples and all conflicting tuples that the dependencies can derive? In other words: which underlying semantics ensure that $m\_Chase$ algorithm is correct?
+  • Does the result contain all true tuples and all conflicting tuples that the dependencies can derive? In other words: which underlying semantics ensure that m_Chasem\_Chase algorithm is correct?
 All these questions find positive answers in LauS25 , based on the set theoretic semantics introduced in CKS86 ; Spyratos87 , under the assumption that the set $FD$ is normalized.
-Following LauS25 , if $FD^{+}$ denotes the closure of $FD$ under the Armstrong’s axioms Armstrong , then $FD$ is said to be normalized if it contains all dependencies in $FD^{+}$ such that:
+Following LauS25 , given a set of attributes $U$ and a set of functional dependencies $FD$ over $U$, if $FD^{+}$ denotes the closure of $FD$ under the Armstrong’s axioms Armstrong , then $FD$ is said to be normalized if it contains all dependencies in $FD^{+}$ such that:
 FD1:
@@
 It is easy to see that the sets $FD$ considered in Example 1 and in Example 2 are both normalized and acyclic.
-In this section, we recall from LauS25 the basic formalism on which the algorithm $m\_Chase$ relies, namely that of multi-valued tuple.
+In this section, we recall from LauS25 the basic formalism on which the algorithm m_Chasem\_Chase relies, namely that of multi-valued tuple.
 A multi-valued tuple, or m-tuple, extends the notion of tuple in the sense that an m-tuple associates every attribute $A$ with a possibly empty subset of the active domain of $A$ as opposed to a single value from the active domain.
@@
 […]
+Similarly, the set theoretic union operation is extended to m-tuples in a ‘component-wise’ manner, that is: the union of m-tuples $\sigma$ and $\sigma^{\prime}$ is the m-tuple denoted by $\sigma\sqcup\sigma^{\prime}$ and for every $A$ in $U$, $(\sigma\sqcup\sigma^{\prime})(A)=\sigma(A)\cup\sigma^{\prime}(A)$.
+For example, if $A$, $B$ and $C$ are three attributes, $\sigma=(aa^{\prime})(b)$ and $\sigma^{\prime}=(a)$ are two m-tuples such that $sch(\sigma)=AB$ and $sch(\sigma^{\prime})=A$.
+In this case, we have $\sigma^{\prime}\sqsubseteq\sigma$ because $\sigma^{\prime}(A)\subseteq\sigma(A)$ and $\sigma^{\prime}(B)=\emptyset$ whereas $\sigma(B)=(b)$, and so, it holds that $\sigma\sqcup\sigma^{\prime}=\sigma$.
+Moreover, if $\sigma^{\prime\prime}=(a)(bb^{\prime})(c)$, then $sch(\sigma^{\prime\prime})=ABC$ and we have $\sigma\sqcup\sigma^{\prime\prime}=(aa^{\prime})(bb^{\prime})(c)$.
 We call m-table over $U$ any finite set of m-tuples over $U$.
-For all $\sigma$ and $\sigma^{\prime}$ in an m-table $\Sigma$, and $X\to A$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma^{\prime})$, the following rule called m-Chase rule generalizes the chase rule.
-$\bullet$ m-Chase rule: Let $\sigma_{1}=\sigma\cup\sigma^{\prime}(A)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}\cup\sigma(A)$
+For all $\sigma$ and $\sigma^{\prime}$ in an m-table $\Sigma$, and for every functional dependency $X\to A$ over $U$ such that $XA\subseteq sch(\sigma)$ and $XA\subseteq sch(\sigma^{\prime})$, the following rule called m_Chase rule generalizes the chase rule.
+$\bullet$ m_Chase rule: Let $\sigma_{1}=\sigma\sqcup\sigma^{\prime}(A)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}\sqcup\sigma(A)$
 Case of $\sigma_{1}\sqsubseteq\sigma^{\prime}_{1}$: replace $\sigma$ with $\sigma^{\prime}_{1}$, and remove $\sigma_{1}$
@@
 Otherwise: replace $\sigma$ and $\sigma^{\prime}$ with $\sigma_{1}$ and $\sigma^{\prime}_{1}$, respectively.
-As shown in Algorithm 1, our algorithm consists in applying the above m-Chase rule whenever ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ until no further transformation is possible.
-The output is an m-table denoted by $m\_Chase(T)$.
-It has been shown in LauS25 that this algorithm always terminates and that the partition semantics of tuples in $\mathcal{T}$ (as introduced in Spyratos87 and extended in JIIS ; LauS25 ), can be defined based on $m\_Chase(T)$ as follows:
+As shown in Algorithm 1, our approach consists in applying the above m_Chase rule whenever ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ until no further transformation is possible.
+According to this algorithm, at each step, if $\sigma$ and $\sigma^{\prime}$ are two m-tuples in the current state of $\Sigma$ whose schemas contain $XA$ where $X\to A$ is a functional dependency in $FD$, if their $X$-value share at least one tuple then their $A$-values are set to the same value $\sigma(A)\cup\sigma^{\prime}(A)$ (see the first and last line of the above m_Chase rule), and in case these changes imply that one m-tuple subsumes the other one, the redundancy is removed (as stated in the two intermediate lines of the above m_Chase rule).
+Similarly to the $Chase$ algorithm, this process is applied in Algorithm 1, until no further transformation of $\Sigma$ is possible.
+The output of Algorithm 1 is then an m-table denoted by m_Chase(T)m\_Chase(T).
+To illustrate this step in the context of Example 2, consider first $\sigma=(a)(b)$ and $\sigma^{\prime}=(b)(c)$ with the dependency $B\to C$.
+In this case $\sigma_{1}=(a)(b)(c)$ and $\sigma^{\prime}_{1}=\sigma^{\prime}$.
+As $\sigma^{\prime}_{1}\sqsubseteq\sigma_{1}$, only $\sigma_{1}$ is kept for the next computation step.
+Thus, for $\sigma=(a)(b)(c)$ and $\sigma^{\prime}=(a)(c^{\prime})$ with the dependency $A\to C$, we have $\sigma_{1}=(a)(b)(cc^{\prime})$ and $\sigma^{\prime}_{1}=(b)(cc^{\prime})$.
+In this case again, we have $\sigma^{\prime}_{1}\sqsubseteq\sigma_{1}$, and so, $\sigma^{\prime}_{1}$ is discarded, and we are left with the only m-tuple $(a)(b)(cc^{\prime})$.
+Since no further application of the m_Chase rule is possible, we obtain that $m\_Chase(T)=\{(a)(b)(cc^{\prime})\}$.
+[…]
+It has been shown in LauS25 that Algorithm 1 always terminates and that the partition semantics of tuples in $\mathcal{T}$ (as introduced in Spyratos87 and extended in JIIS ; LauS25 ), can be defined based on m_Chase(T)m\_Chase(T) according to the following proposition.
+To state these basic results, we use the following notation given a table $T$: the sets of all true tuples in $\mathcal{T}$, all conflicting tuples in $\mathcal{T}$, all consistent tuples in $\mathcal{T}$ are respectively denoted by ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$.
 Algorithm 1 The m-Chase Algorithm
 0:A table $T$ over $U$ and a normalized set $FD$ of functional dependencies over $U$.
-0:  An m-table denoted by $m\_Chase(T)$.
-1:  $\Sigma:=\{\sigma_{t}~{}|~{}t\in T\}$ // $\sigma_{t}$ is the m-tuple such that $\sigma_{t}(A)=\{t.A\}$ for $A\in sch(t)$
+0: An m-table denoted by m_Chase(T)m\_Chase(T).
+[…]
 2:$change:=true$
@@
 7:    if ${\sf tuples}(\sigma(X))\cap{\sf tuples}(\sigma^{\prime}(X))\neq\emptyset$ then
-8:              apply the m-Chase rule to $\sigma$ and $\sigma^{\prime}$
+8:     apply the m_Chase rule to $\sigma$ and $\sigma^{\prime}$
 9:     $change:=true$
-10:  $m\_Chase(T):=\Sigma$
-11:  return  $m\_Chase(T)$
+[…]
+11:return m_Chase(T)m\_Chase(T)
 Proposition 1
 Let $T$ be a table over universe $U$ with $FD$ as set of functional dependencies. The following holds:
-  1. 1.
-    A tuple $t$ in $\mathcal{T}$ is in ${\sf True}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
-    -- $sch(t)\subseteq sch(\sigma)$ (i.e., $\sigma$ has nonempty components over attributes on which $t$ is defined),
-    -- $t\sqsubseteq\sigma$.
-  2. 2.
-    A tuple $t$ in $\mathcal{T}$ is in ${\sf Confl}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
-    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
-    -- there exists $X\to A$ in $FD$ such that $XA\subseteq sch(t)$ and $|{\sf tuples}(\sigma(A)|>1$.
-  3. 3.
-    A tuple $t$ in $\mathcal{T}$ is in ${\sf Cons}(\mathcal{T})$ if and only if there exists $\sigma$ in $m\_Chase(T)$ such that:
-    -- $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
-    -- for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $|{\sf tuples}(\sigma(A))|=1$.
-  4. 4.
-    For every $\sigma$ in $m\_Chase(T)$ and every $S\subseteq sch(\sigma)$, either ${\sf tuples}(\sigma(S))\subseteq{\sf Cons}(\mathcal{T})$ or ${\sf tuples}(\sigma(S))\subseteq{\sf Confl}(\mathcal{T})$.
+  • A tuple $t$ in $\mathcal{T}$ is in ${\sf True}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
+    - $sch(t)\subseteq sch(\sigma)$ (i.e., $\sigma$ has nonempty components over attributes in $sch(t)$),
+    - $t\sqsubseteq\sigma$.
+  • A tuple $t$ in $\mathcal{T}$ is in ${\sf Confl}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
+    - $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
+    - there exists $X\to A$ in $FD$ such that $XA\subseteq sch(t)$ and $|{\sf tuples}(\sigma(A)|>1$.
+  • A tuple $t$ in $\mathcal{T}$ is in ${\sf Cons}(\mathcal{T})$ if and only if there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that:
+    - $sch(t)\subseteq sch(\sigma)$ and $t\sqsubseteq\sigma$,
+    - for every $X\to A$ in $FD$ such that $XA\subseteq sch(t)$, $|{\sf tuples}(\sigma(A))|=1$.
+  • For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $S\subseteq sch(\sigma)$, either ${\sf tuples}(\sigma(S))\subseteq{\sf Cons}(\mathcal{T})$ or ${\sf tuples}(\sigma(S))\subseteq{\sf Confl}(\mathcal{T})$.
 $\Box$
-As shown in LauS25 , the computation of $m\_Chase(T)$ is in $\mathcal{O}(|m\_Chase(T)|^{3}.\delta^{2})$, where $\delta$ is the maximal cardinality of the components of m-tuples in $m\_Chase(T)$, which is precisely the maximum number of $A$-values associated with $X$-values when $X\to A$ is a functional dependency in $FD$.
-As Algorithm 1 shows that $|m\_Chase(T)|\leq|T|$, we state that the computation of $m\_Chase(T)$ is in $\mathcal{O}(|T|^{3}.\delta^{2})$, i.e., polynomial in the size of $T$.
+We notice that when ${\sf Confl}({T})=\emptyset$, i.e., when no tuples in $\mathcal{T}$ are conflicting, then $T$ satisfies all functional dependencies in $FD$, which is denoted by $T\models FD$.
+In this case, as shown by Proposition 1, all components of all m-tuples in m_Chase(T)m\_Chase(T) contain one value, and so, m_Chase(T)m\_Chase(T) can be seen as a table, which is in fact the table $Chase(T)$.
+Regarding the time complexity of Algorithm 1, it has been shown in LauS25 that the computation of m_Chase(T)m\_Chase(T) is in $\mathcal{O}(|m\_Chase(T)|^{3}\cdot\delta^{2})$, where $\delta$ is the maximal cardinality of the components of m-tuples in m_Chase(T)m\_Chase(T), which is precisely the maximum number of $A$-values associated with $X$-values when $X\to A$ is a functional dependency in $FD$.
+As Algorithm 1 shows that $|m\_Chase(T)|\leq|T|$, we state that the computation of m_Chase(T)m\_Chase(T) is in $\mathcal{O}(|T|^{3}\cdot\delta^{2})$, i.e., polynomial in the size of $T$.
+We notice that in this complexity result, the size of the universe $U$ is a constant, contrary to the size of m-tuples.
+In fact the size of every m-tuple in m_Chase(T)m\_Chase(T) is bounded by $\delta\cdot|U|$.
+As will be seen shortly, these bounds can be refined when considering star-schemas.
+[…]
+Thus, using Proposition 1, the set ${\sf Confl}_{\min}(\mathcal{T})$ is characterized as follows: $t$ is in ${\sf Confl}_{\min}(\mathcal{T})$ if and only if one of the following two statements holds:
+  • there exist $\sigma$ in m_Chase(T)m\_Chase(T), $i_{0}$ in $\{1,\ldots,n\}$ and $A$ in $sch^{*}(D_{i_{0}})\cap sch(\sigma)\cap sch(t)$ such that $t=k_{i_{0}}a$ is in ${\sf tuples}(\sigma(K_{i_{0}}A))$ and $|{\sf tuples}(\sigma(A))|>1$,
+  • there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)\cap sch(t)$, there exists $M_{i}$ in $\mathbb{M}\cap sch(\sigma)\cap sch(t)$ such that $t=km_{i}$ is in ${\sf tuples}(\sigma(\mathbb{K}M_{i}))$ and $|{\sf tuples}(\sigma(M_{i}))|>1$.
+By complementation with respect to ${\sf True}(\mathcal{T})$, a tuple $t$ is in ${\sf Cons}(\mathcal{T})$ if and only if it has no sub-tuple satisfying one of the above statements.
 To illustrate Algorithm 1 and Proposition 1, consider again the context of Example 2 where $U=\{A,B,C\}$, $FD=\{A\to C,B\to C\}$ and $T=\{ab,bc,ac^{\prime}\}$.
-Running Algorithm 1 yields the following steps:
-  • The algorithm starts with the m-table $\Sigma=\{(a)(b),(b)(c),(a)(c^{\prime})\}$
-  • Applying $B\to C$ to the first two m-tuples, we obtain $\Sigma=\{(a)(b)(c),(a)(c^{\prime})\}$
+As seen earlier, running Algorithm 1 yields the following steps:
+- The algorithm starts with the m-table $\Sigma=\{(a)(b),(b)(c),(a)(c^{\prime})\}$.
+- Applying $B\to C$ to the first two m-tuples, we obtain $\Sigma=\{(a)(b)(c),(a)(c^{\prime})\}$.
 - Applying now $A\to C$ to these two m-tuples, we obtain $\Sigma=\{(a)(b)(cc^{\prime})\}$.
@@
 In other words, there are no false tuples in this example.
-  • ${\sf Confl}(\mathcal{T})=\{abc,abc^{\prime},ac,ac^{\prime},bc,bc^{\prime}\}$.
+  • ${\sf Confl}_{min}(\mathcal{T})=\{ac,ac^{\prime},bc,bc^{\prime}\}$ and so, we have ${\sf Confl}(\mathcal{T})=\{abc,abc^{\prime},ac,ac^{\prime},bc,bc^{\prime}\}$.
   • ${\sf Cons}(\mathcal{T})={\sf True}(\mathcal{T})\setminus{\sf Confl}(\mathcal{T})$, that is ${\sf Cons}(\mathcal{T})=\{ab,a,b,c,c^{\prime}\}$.
-In the following section, we first define the notion of star schema and then we show that the results from LauS25 that have just been recalled apply in this context as well.
+In the following section, we first recall the notion of star schema and then we show that the results from LauS25 that have just been recalled apply in this context as well.
 3 Star Schemas
 3.1 The Context of our Approach
-We first recall from the literature Ullman that a star schema, as considered in our approach, consists of the following tables and constraints:
+We first recall from the literature inmon96 ; Ullman that a star schema, as considered in our approach, consists of the following tables and constraints:
   • $n$ dimension tables $D_{1},\ldots,D_{n}$.
@@
   • a fact table $F$ defined over $K_{1},\ldots,K_{n},M_{1},\ldots,M_{p}$.
-The attributes $M_{1},\ldots,M_{p}$ are called measures, and we denote by $\mathbf{M}$ the set of all measures, that is $\mathbf{M}=\{M_{1},\ldots,M_{p}\}$.
+The attributes $M_{1},\ldots,M_{p}$ are called measures, and we denote by $\mathbb{M}$ the set of all measures, that is $\mathbb{M}=\{M_{1},\ldots,M_{p}\}$.
   • $FD=\bigcup_{i=1}^{i=n}\{K_{i}\to A_{i}^{j}~|~j=1,\ldots,d_{i}\}\cup\{K_{1}\ldots K_{n}\to M_{k}~|~k=1,\ldots,p\}$.
     In other words, for $i=1,\ldots,n$, $K_{i}$ is the key of $D_{i}$ and $K_{1}\ldots K_{n}$ is the key of $F$.
-We denote by $\mathbf{K}$ the set of all dimension keys that is $\mathbf{K}=\{K_{1},\ldots,K_{n}\}$.
-It is easy to see that if $FD$ is defined as above, then for every non trivial functional dependency $X\to A$ in $FD^{+}$ we have: $X\cap\mathbf{K}\neq\emptyset$.
-More precisely, if $A$ is in $\mathbf{K}$ then $A$ must occur in $X$, in which case $X\to A$ is trivial (because $FD$ contains no dependency whose right hand-side is in $\mathbf{K}$), if $A$ is in $sch^{*}(D_{i})$ then $X$ must contain $K_{i}$ and if $A$ is in $\mathbf{M}$ then $X$ must contain $\mathbf{K}$.
+We denote by $\mathbb{K}$ the set of all dimension keys that is $\mathbb{K}=\{K_{1},\ldots,K_{n}\}$.
+It is easy to see that if $FD$ is defined as above, then for every non trivial functional dependency $X\to A$ in $FD^{+}$, it holds that $X\cap\mathbb{K}\neq\emptyset$ and $A\not\in\mathbb{K}$.
+More precisely, given an attribute $A$, only one of the following three cases is possible: $(i)$ AA is in $\mathbb{K}$, or $(ii)$ AA is in $sch^{*}(D_{i})$ for some $i=1,\ldots,n$, or $(iii)$ AA is in $\mathbb{M}$.
+For each of these cases, we have the following: $(i)$ if $A$ is in $\mathbb{K}$ then $A$ must occur in $X$, in which case $X\to A$ is trivial (because $FD$ contains no dependency whose right hand-side is in $\mathbb{K}$); $(ii)$ if $A$ is in $sch^{*}(D_{i})$ then $X$ must contain $K_{i}$; $(iii)$ if $A$ is in $\mathbb{M}$ then $X$ must contain $\mathbb{K}$.
 Thus, for every non trivial functional dependency $X\to A$ in $FD^{+}$, there exists $X_{0}\to A$ in $FD$ such that $X_{0}\subseteq X$.
 Since the left hand-sides of the dependencies in $FD$ can not be reduced further, this means that $FD$ is normalized.
-On the other hand as the left hand-sides of functional dependencies in $FD$ are attributes in $\mathbf{K}$, not occurring in the right hand-sides of these dependencies, $FD$ is acyclic.
+On the other hand as the left hand-sides of functional dependencies in $FD$ are attributes in $\mathbb{K}$, not occurring in the right hand-sides of these dependencies, $FD$ is acyclic.
 As a consequence, all results in LauS25 apply in the context of star schemas.
@@
   2. 2.
-    For every $t$ in $F$, $t$ is defined over $\mathbf{K}$ and over at least one measure attribute in $\mathbf{M}$. We consider that storing a fact with no associated measure value makes no sense.
-Example 3
-We illustrate the above concepts using a toy example that will serve as a running example in the remainder of this paper.
-We consider two dimensions $D_{1}$ and $D_{2}$ such that $sch(D_{1})=K_{1}A_{1}^{1}A_{1}^{2}$ and $sch(D_{2})=K_{1}A_{2}^{1}A_{2}^{2}$.
-Moreover the fact table $F$ is such that $sch(F)=K_{1}K_{2}M_{1}$ meaning that we consider one measure attribute $M_{1}$.
-As specified above, we have $FD=\{K_{1}\to A_{1}^{1},K_{1}\to A_{1}^{2},K_{2}\to A_{2}^{1},K_{2}\to A_{2}^{2},K_{1}K_{2}\to M_{1}\}$.
-The content of the tables $D_{1}$, $D_{2}$ and $F$ is shown in the following Figure 1.
+    For every $t$ in $F$, $t$ is defined over $\mathbb{K}$ and over at least one measure attribute in $\mathbb{M}$.
+We consider that storing a fact in which some key values are missing makes no sense, since we consider that measure values depend on all dimensions.
+Similarly, considering a fact with no associated measure value makes no sense.
+We notice that the above restrictions follow common sense considerations (storing values without corresponding key values or storing key values alone with no other attribute values make no sense).
+However, we also notice that relaxing these restrictions can be supported by our approach, at the cost of technical complications.
 $D_{1}$ $K_{1}$ $A_{1}^{1}$ $A_{1}^{2}$
@@
 Figure 1: The tables of the data warehouse in our running example
+Example 3
+We illustrate the above concepts using a toy example that will serve as a running example in the remainder of this paper.
+We consider two dimensions $D_{1}$ and $D_{2}$ such that $sch(D_{1})=K_{1}A_{1}^{1}A_{1}^{2}$ and $sch(D_{2})=K_{2}A_{2}^{1}A_{2}^{2}$.
+Moreover the fact table $F$ is such that $sch(F)=K_{1}K_{2}M_{1}$ meaning that we consider one measure attribute $M_{1}$.
+As specified above, we have $FD=\{K_{1}\to A_{1}^{1},K_{1}\to A_{1}^{2},K_{2}\to A_{2}^{1},K_{2}\to A_{2}^{2},K_{1}K_{2}\to M_{1}\}$.
+The content of the tables $D_{1}$, $D_{2}$ and $F$ is shown in Figure 1.
 We observe that these tables are indeed those of a star schema and that they comply with the two restrictions above regarding missing values.
@@
 On the other hand $D_{2}$ satisfies its two associated functional dependencies $K_{2}\to A_{2}^{1}$ and $K_{2}\to A_{2}^{2}$.
-We also stress that the key value $k^{\prime\prime}_{1}$ occurs in $D_{1}$ but not in $F$, whereas the key value $k^{\prime\prime}_{2}$ over $K_{2}$ occurs in $F$ but not in $D_{2}$.
+We also stress that the above two restrictions on missing values are satisfied and that key value $k^{\prime\prime}_{1}$ occurs in $D_{1}$ but not in $F$, whereas the key value $k^{\prime\prime}_{2}$ over $K_{2}$ occurs in $F$ but not in $D_{2}$.
 These two cases respectively illustrate that key values in a dimension table may not occur in the fact table and that the foreign key constraint between a dimension table and the fact table may not be satisfied (contrary to what is generally assumed in the literature).
-$\Box$
-3.2 True, Consistent and Conflicting Tuples in a Star Schema
-The following proposition states important properties regarding m-tuples of the table $m\_Chase(T)$ where $T$ is the table collecting all tuples in the data warehouse.
-In the remainder of this paper, we refer to such table as a star-table.
-Proposition 2
-Let $T$ be a star-table over universe $U$. The following hold:
-  1. 1.
-    For every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$.
-Consequently, if $\mathbf{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbf{K}))|=1$.
-  2. 2.
-    For every tuple $k$ over $\mathbf{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$.
-  3. 3.
-    Moreover, $m\_Chase(T)$ contains the following two kinds of m-tuples:
-    1. (a)
-      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
-      -- $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=\{k_{i_{0}}\}$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$.
-      -- for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~{}|~{}(\exists q\in D_{i})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$
-    2. (b)
-      $\sigma$ such that $\mathbf{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbf{K}))=\{k\}$, and
-      -- for every $M_{j}\in\mathbf{M}$, $\sigma(M_{j})=\{m_{j}~{}|~{}(\exists t\in F)(t.\mathbf{K}=\sigma(\mathbf{K})\wedge t.M_{j}=m_{j})\}$
-      -- for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$, $\sigma(A)=\{a~{}|~{}(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
-Proof.
-1.
-This result comes from the fact that, in order to generate multi-valued components of an attribute $A$, the algorithm $m\_Chase$ has to consider a functional dependency whose right hand-side is $A$.
-As in the case of a star-table, no dependency in $FD$ has its right hand-side in $\mathbf{K}$, the proof of this item is complete.
-2.
-Let $\sigma_{1}$ and $\sigma_{2}$ be in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma_{i})$ for $i=1,2$ and $\sigma_{1}(\mathbf{K})=\sigma_{2}(\mathbf{K})$.
-Since $\sigma_{1}\neq\sigma_{2}$ (the algorithm eliminates duplicates) there exists $A$ not in $\mathbf{K}$ such that $\sigma_{1}(A)\neq\sigma_{2}(A)$.
-Since $A$ is not in $\mathbf{K}$, either there is $i_{0}$ in $\{1,\ldots,n\}$ such that $A\in sch^{*}(D_{i_{0}})$ or $A$ is in $\mathbf{M}$.
-In either case, $FD$ contains a dependency of the form $X\to A$ where $X=K_{i_{0}}$ in the former case, or $X=\mathbf{K}$ in the latter case.
-Since we have $\sigma_{1}(X)=\sigma_{2}(X)$, applying $m\_Chase$ to $m\_Chase(T)$ changes $\sigma_{1}(A)$ and $\sigma_{2}(A)$ into $\sigma_{1}(A)\cup\sigma_{2}(A)$.
-As by definition of $m\_Chase(T)$, we have $m\_Chase(m\_Chase(T))=m\_Chase(T)$, we obtain a contradiction which completes the proof.
-3.
-We first notice that, as stated above, for every $K_{i}$ in $\mathbf{K}$ and every $\sigma$ in $m\_Chase(T)$ such that $K_{i}$ is in $sch(\sigma)$, we have $|{\sf tuples}(\sigma(K_{i}))|=1$.
-This is so because no functional dependency in $FD$ has a right hand-side in $\mathbf{K}$, which makes it impossible to generate conflicts on these attributes.
-Moreover, the m-tuples of the first kind are the result of joining two rows $\rho_{1}$ and $\rho_{2}$ in the current table such that $sch(\rho_{1})=sch(\rho_{2})=sch(D_{i_{0}})$, $\rho_{1}(K_{i_{0}})=\rho_{2}(K_{i_{0}})$, and this key value does not occur in $F$ thus preventing these m-tuples to be combined with m-tuples over $sch(F)$.
-Similarly, the tuples of the second kind are obtained by joining every tuple $km$ in $F$ with tuples $t_{i}$ in $D_{i}$ such that $k.K_{i}=t_{i}.K_{i}$ ($i=1,\ldots,n$).
-We thus obtain that the set $\Sigma$ of all m-tuples as described above occur in $m\_Chase(T)$.
-To end the proof we now argue that applying the $m\_Chase$ procedure to $\Sigma$ has no effect.
-This is so because for every $X\to A$ in $FD$ and all rows $\rho_{1}$ and $\rho_{2}$ in $\Sigma$ such that $\rho_{1}(X)\cap\rho_{2}(X)\neq\emptyset$, the $X$-values of these rows are reduced to one tuple and thus we have $\rho_{1}(X)=\rho_{2}(X)$.
-By definition of $\Sigma$, it follows that $\rho_{1}(A)=\rho_{2}(A)$, which completes the proof.
-$\Box$
 $T$ $K_{1}$ $K_{2}$    $A_{1}^{1}$   $A_{1}^{2}$     $A_{2}^{1}$   $A_{2}^{2}$  $M_{1}$
@@
 $k^{\prime}_{1}$ $k^{\prime\prime}_{2}$ $m_{1}$
-$m\_Chase(T)$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
+m_Chase(T)m\_Chase(T) $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
 $(k_{1})$ $(k_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(b_{2})$ $(m_{1})$
@@
 $(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
-Figure 2: The star-table $T$ and the m-table $m\_Chase(T)$ of Example 3
-In the context of Example 3, the star-table $T$ and its associated m-table $m\_Chase(T)$ are shown in Figure 2.
-In particular, it can be seen that, in $m\_Chase(T)$, the last m-tuple satisfies item 3.a of Proposition 2 while the first three m-tuples satisfy item 3.b.
-As will be seen later, these three m-tuples are the ones to be relevant for analytic queries.
-[…]
-Using Proposition 1, the set ${\sf Confl}_{\min}(\mathcal{T})$ is characterized as follows: $t$ is in ${\sf Confl}_{\min}(\mathcal{T})$ if and only if one of the following two statements holds:
-  • there exist $\sigma$ in $m\_Chase(T)$, $i_{0}$ in $\{1,\ldots,n\}$ and $A$ in $sch^{*}(D_{i_{0}})\cap sch(\sigma)\cap sch(t)$ such that $t=k_{i_{0}}a$ is in ${\sf tuples}(\sigma(K_{i_{0}}A))$ and $|{\sf tuples}(\sigma(A))|>1$
-  • there exists $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\subseteq sch(\sigma)\cap sch(t)$, there exists $M_{i}$ in $\mathbf{M}\cap sch(\sigma)\cap sch(t)$ such that $t=km_{i}$ is in ${\sf tuples}(\sigma(\mathbf{K}M_{i}))$ and $|{\sf tuples}(\sigma(M_{i}))|>1$.
-By complementation with respect to ${\sf True}(\mathcal{T})$, a tuple $t$ is in ${\sf Cons}(\mathcal{T})$ if and only if it has no sub-tuple satisfying one of the above statements.
-Example 4
-Applying Proposition 1 in the context of Example 3, for which Figure 2 displays the star-table $T$ and the m-table $m\_Chase(T)$, the sets ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$ are as follows:
-  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of the tuples in ${\sf tuples}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. Thus, ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of:
-    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
-    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
-    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
-    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
+Figure 2: The star-table $T$ and the m-table m_Chase(T)m\_Chase(T) of Example 3
+In this context, the associated table $T$ over $U$ and its corresponding m-table m_Chase(T)m\_Chase(T) are shown in Figure 2.
+Applying Proposition 1, the sets ${\sf True}(\mathcal{T})$, ${\sf Confl}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})$ are as follows:
+  • ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of the tuples in ${\sf tuples}(\sigma)$ for every $\sigma$ in m_Chase(T)m\_Chase(T).
+Thus, ${\sf True}(\mathcal{T})$ is the set of all sub-tuples of:
+    - $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
+    - $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
+    - $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
+    - $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
   • ${\sf Confl}(\mathcal{T})$ is the set of all true super-tuples of tuples in ${\sf Confl}_{\min}(\mathcal{T})=\{k_{1}a_{1}$, $k_{1}a^{\prime}_{1}$, $k_{1}k^{\prime}_{2}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}m^{\prime\prime}_{1}\}$.
 The maximal tuples in ${\sf Confl}(\mathcal{T})$ are:
-    -- $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
-    -- $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$.
+    - $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
+    - $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$.
   • ${\sf Cons}(\mathcal{T})={\sf True}(\mathcal{T})\setminus{\sf Confl}(\mathcal{T})$. Thus ${\sf Cons}(\mathcal{T})$ is the set of all sub-tuples of:
-    -- $k_{1}k_{2}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
-    -- $k_{1}k^{\prime}_{2}a_{2}b_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
-    -- $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
-    -- $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$. $\Box$
+    - $k_{1}k_{2}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{2}a^{\prime}_{1}a_{2}b_{1}b_{2}m_{1}$,
+    - $k_{1}k^{\prime}_{2}a_{2}b_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k^{\prime}_{2}a^{\prime}_{1}a_{2}b_{1}m^{\prime\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime}_{1}$, $k_{1}a_{2}b_{1}m^{\prime\prime}_{1}$,
+    - $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$,
+    - $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$. $\Box$
+3.2 m_Chasem\_Chase Computation in a Star Schema
+In this section, we show that when considering a star schema, the time complexity of computing m_Chase(T)m\_Chase(T) for a star-table $T$ can be drastically lowered, as compared with the general case shown in LauS25 and recalled above.
+To do so, we first state important properties regarding m-tuples of m_Chase(T)m\_Chase(T) where $T$ is the table collecting all tuples in the data warehouse.
+In the remainder of this paper, we refer to such table as a star-table.
+The following proposition states important properties of the m-tuples in m_Chase(T)m\_Chase(T).
+Proposition 2
+Let $T$ be a star-table over universe $U$. The following hold:
+  1. 1.
+    For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$.
+Consequently, if $\mathbb{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbb{K}))|=1$.
+  2. 2.
+    For every tuple $k$ over $\mathbb{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$.
+  3. 3.
+    Moreover, m_Chase(T)m\_Chase(T) contains the following two kinds of m-tuples:
+    1. (a)
+      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
+      • $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$,
+      • for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
+    2. (b)
+      $\sigma$ such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$, and
+      • for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})=\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$,
+      • for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$,
+        $\sigma(A)=\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
+Proof. See Appendix A. $\Box$
+We point out that item 3 in Proposition 2 holds thanks to the restrictions on missing values earlier stated.
+Moreover, an important consequence of Proposition 2 is that m_Chase(T)m\_Chase(T) can be computed by an algorithm whose time complexity is significantly lower than that of Algorithm 1.
+This new algorithm, displayed in Algorithm 2, is shown to be correct in the following proposition.
+Algorithm 2 The m-Chase Algorithm for a star-table
+0: A data warehouse composed of $n$ dimensional tables $D_{1},\ldots,D_{n}$ and a fact table $F$.
+0: The m-table m_Chase(T)m\_Chase(T).
+1:for all $i=1,\ldots,n$ do
+2:  sort the m-tuples in $D_{i}$ according to their $K_{i}$-value
+3:  $D_{i}^{*}:=\emptyset$
+4:  for all $K_{i}$-value $k_{i}$ occurring in $D_{i}$ do
+5:   $\sigma:=(k_{i})s_{i}^{1}\ldots s_{i}^{d_{i}}$ where $s_{i}^{j}=\{a_{i}^{j}\in adom(A_{i}^{j})~|~(\exists t_{i}\in D_{i})(t_{i}.{K_{i}}=k_{i}\wedge A_{i}^{j}\in sch(t_{i})\wedge t_{i}.A_{i}^{j}=a_{i})\}$ for every $j=1,\ldots,d_{i}$
+6:   $D_{i}^{*}:=D_{i}^{*}\cup\{\sigma\}$
+7: sort the m-tuples in $F$ according to their $\mathbb{K}$-value
+[…]
+9:for all $\mathbb{K}$-value $k=k_{1}\ldots k_{n}$ occurring in $F$ do
+10:  $\sigma:=(k)\mu_{1}\ldots\mu_{p}$ where $\mu_{q}=\{m_{q}\in adom(M_{q})~|~(\exists t\in F)(t.{\mathbb{K}}=k\wedge M_{q}\in sch(t)\wedge t.M_{q}=m_{q})\}$ for every $q=1,\ldots,p$, and $\sigma(A)=\emptyset$ for every $A\in U\setminus(\mathbb{K}\cup\mathbb{M})$
+11:  for all $i=1,\ldots,n$ do
+12:   if $k_{i}$ occurs in $D_{i}^{*}$ then
+13:    Let $\sigma_{i}$ be the unique m-tuple in $D_{i}^{*}$ such that $\sigma(K_{i})=(k_{i})$
+14:    Mark $\sigma_{i}$ in $D_{i}^{*}$
+15:    for all $j=1,\ldots,d_{i}$ do
+16:     $\sigma(A_{i}^{j}):=\sigma_{i}(A_{i}^{j})$
+17:  $F^{*}:=F^{*}\cup\{\sigma\}$
+18:for all $i=1,\ldots,n$ do
+19:  for all $\sigma_{i}$ in $D_{i}^{*}$ do
+20:   if $\sigma_{i}$ is not marked then
+21:    $F^{*}:=F^{*}\cup\{\sigma_{i}\}$
+[…]
+23:return m_Chase(T)m\_Chase(T)
+Proposition 3
+Given a data warehouse composed of $n$ dimensional tables $D_{1},\ldots,D_{n}$ and a fact table $F$, and its associated star-table $T$, Algorithm 2 correctly computes m_Chase(T)m\_Chase(T).
+The time complexity of Algorithm 2 is in
+$\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))+|F|\cdot(\log(|F|))\right).$
+Proof.
+It can be seen that the loop lines 18-21 in Algorithm 2 computes all m-tuples of kind (a) in Proposition 2(3), whereas the loop lines 9-17 in Algorithm 2 computes all m-tuples of kind (b).
+The first part of the proof is therefore complete.
+Regarding the time complexity of Algorithm 2, we first notice that the statements line 2 and 7 are respectively in $\mathcal{O}(|D_{i}]\cdot\log(|D_{i}|))$ and $\mathcal{O}(|F]\cdot\log(|F|))$.
+Moreover, the loop lines 4-6 is linear in the size of $D_{i}$ as the table is sorted according to its key values.
+Thus the loop lines 1-6 is in $\mathcal{O}(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|)))$.
+Similarly, the loop lines 9-17 is in $\mathcal{O}(|F]\cdot\log(|F|))$ and the last loop lines 18-21 is in $\mathcal{O}(|D_{i}^{*}|)$.
+Since $\mathcal{O}(|D_{i}^{*}|)\approx\mathcal{O}(|D_{i}|)$, we obtain that Algorithm 2 is in $\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))\right)+\mathcal{O}(|F]\cdot\log(|F|))+\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|)\right)$, that is in $\mathcal{O}\left(\Sigma_{i=1}^{i=n}(|D_{i}|\cdot\log(|D_{i}|))\right)+\mathcal{O}(|F]\cdot\log(|F|))$.
+The proof is therefore complete.
+$\Box$
+We emphasize that, denoting by $W$ the overall size of the underlying data warehouse (i.e., the sum of the sizes of all its tables), the complexity result in Proposition 3 can be written as $\mathcal{O}(W\cdot\log(W))$, which is significantly lower than the generic time complexity stated in LauS25 , which can be approximated to $\mathcal{O}(W^{3}\cdot\delta^{2})$.
+Moreover, considering Algorithm 2 instead of Algorithm 1 to assess the size of the m-table m_Chase(T)m\_Chase(T) shows that the number of m-tuples in m_Chase(T)m\_Chase(T) is exactly the number of distinct $\mathbb{K}$-values present in the fact table $F$, plus for every $i=1,\ldots,n$, the number of all $K_{i}$-values occurring in $D_{i}$ but not in $F$.
+On the other hand it is easy to see that, if $T$ is a star-table, the size of any m-tuple in m_Chase(T)m\_Chase(T) is bounded by $|\mathbb{K}|+\delta\cdot|U\setminus\mathbb{K}|$, which is less than $\delta\cdot|U|$ as earlier stated.
+We now illustrate Proposition 3 in the context of Example 3.
+$D_{1}^{*}$ $K_{1}$ $A_{1}^{1}$ $A_{1}^{2}$
+$(k_{1})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$
+$(k^{\prime}_{1})$ $(a_{1})$
+$(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
+$D_{2}^{*}$ $K_{2}$ $A_{2}^{1}$ $A_{2}^{2}$
+$(k_{2})$ $(b_{1})$ $(b_{2})$
+$(k^{\prime}_{2})$ $(b_{1})$
+$F^{*}_{0}$ $K_{1}$ $K_{2}$ $M_{1}$
+$(k_{1})$ $(k_{2})$ $(m_{1})$
+$(k_{1})$ $(k^{\prime}_{2})$ $(m^{\prime}_{1}\,m^{\prime\prime}_{1})$
+$(k^{\prime}_{1})$ $(k^{\prime\prime}_{2})$ $(m_{1})$
+$F^{*}$ $K_{1}$ $K_{2}$ $A_{1}^{1}$ $A_{1}^{2}$ $A_{2}^{1}$ $A_{2}^{2}$ $M_{1}$
+$(k_{1})$ $(k_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(b_{2})$ $(m_{1})$
+$(k_{1})$ $(k^{\prime}_{2})$ $(a_{1}\,a^{\prime}_{1})$ $(a_{2})$ $(b_{1})$ $(m^{\prime}_{1}\,m^{\prime\prime}_{1})$
+$(k^{\prime}_{1})$ $(k^{\prime\prime}_{2})$ $(a_{1})$ $(m_{1})$
+$(k^{\prime\prime}_{1})$ $(a^{\prime}_{1})$ $(a^{\prime}_{2})$
+Figure 3: The m-tables when running Algorithm 2 in the context of Example 3
+Example 4
+Applying Algorithm 2 to the tables shown in Figure 1, yields the following m-tables shown in Figure 3:
+  • $D_{1}^{*}$ and $D_{2}^{*}$, when running the loop lines 1-6,
+  • $F_{0}^{*}$ when considering the m-tuples $\sigma$ on line 10,
+  • the first three rows in $F^{*}$ when completing the loop lines 2-5,
+  • the whole m-table $F^{*}$, after computing the loop lines 18-21.
+Indeed, as $k^{\prime\prime}_{1}$ does not occur in $F$, the last m-tuple in $D_{1}^{*}$ is not marked because the test line 12 fails.
+It is easy to see that the obtained m-table $F^{*}$ is equal to m_Chase(T)m\_Chase(T) as shown in Figure 2.
+Moreover, the number of m-tuples in m_Chase(T)m\_Chase(T) is $5$, corresponding respectively to the four $\mathbb{K}$-values occurring in $F$ and to the $K_{1}$-value $k_{1}^{\prime\prime}$ occurring in $D_{1}$ but not in $F$.
+$\Box$
 4 Repairs and their Basic Properties
@@
 As explained in LauS25 , contrary to most approaches to consistent query answering in the literature, it is not appropriate to define a repair of $T$ as a maximal subset of ${\sf True}(\mathcal{T})$ satisfying $FD$.
-This is so because it is intuitively justified to define a repair $R$ of $T$ so as ${\sf True}(\mathcal{R})$ contains ${\sf Cons}(\mathcal{T})$.
-Unfortunately as shown earlier in Example 2, it may happen that a maximal consistent subset S of ${\sf True}(\mathcal{T})$ does not contain all consistent tuples.
-On the other hand, it has also been shown in LauS25 , that in case of cyclic sets of functional dependencies (that is if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A\in Y$ and $B\in X$ hold), the set ${\sf Cons}(\mathcal{T})$ may not satisfy $FD$.
-As in the case of a star-table, the set $FD$ of functional dependencies is clearly acyclic, ${\sf Cons}(\mathcal{T})$ satisfies $FD$. We thus define repairs of star-tables as follows.
-Definition 2
-Let $T$ be a star-table over universe $U$. A repair $R$ of $T$ is a table over $U$ such that:
+This is so because, although it is intuitively justified to impose that all tuples of ${\sf Cons}(\mathcal{T})$ be true in every repair, it is shown in LauS25 that in case of cyclic sets of functional dependencies (i.e., if there exist $X\to A$ and $Y\to B$ in $FD$ such that $A\in Y$ and $B\in X$), the set ${\sf Cons}(\mathcal{T})$ may not satisfy $FD$.
+To cope with this issue, given universe $U$, a set $FD$ of functional dependencies over $U$ and a table $T$ over $U$, a repair $R$ of $T$ is defined in LauS25 as a table $R$ over $U$ such that:
   1. 1.
-    ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$
+    ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$.
   2. 2.
-    $R\models FD$
+    $R\models FD$.
   3. 3.
-    For every table $R^{\prime}$ satisfying 1, and 2 above, and such that ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}^{\prime})$.
-The set of all repairs of $T$ is denoted by ${\sf Rep}(T)$. $\Box$
-It has been shown in LauS25 that the following holds, based on Definition 2:
+    ${\sf True}(\mathcal{R})\cap{\sf Cons}(\mathcal{T})$ is a maximal consistent subset of ${\sf Cons}(\mathcal{T})$.
+  4. 4.
+    For every table $R^{\prime}$ satisfying 1, 2 and 3 above, and such that ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf True}(\mathcal{R})={\sf True}(\mathcal{T})$.
+Moreover, it is also shown in LauS25 that when $FD$ is acyclic, the set ${\sf Cons}(\mathcal{T})$ satisfies $FD$.
+Hence, it turns out from the above definition that when $FD$ is acyclic, every tuple in ${\sf Cons}(\mathcal{T})$ must be true in every repair.
+It is important to notice that, as shown in Example 2, even when $FD$ is acyclic, there exist maximal and consistent subsets of ${\sf True}(\mathcal{T})$ that do not contain all tuples in ${\sf Cons}(\mathcal{T})$.
+This shows that, even when considering acyclic sets of functional dependencies, our notion of repair does not exactly coincide with that of traditional approaches, in which repairs would be defined as maximal subsets of ${\sf True}(\mathcal{T})$ satisfying $FD$.
+However, we show in the next sub-section (see Proposition 4) that these two notions of repairs coincide, when considering star-tables as we do in the present approach.
+It has also been shown in LauS25 that when the set $FD$ is acyclic, then the following holds:
   • For every table $T$, ${\sf Rep}(T)\neq\emptyset$.
@@
 For example for $FD=\emptyset$ and $T=\{abc,ac\}$, $R=\{abc\}$ is a repair of $T$.
-Elaborating briefly on the remark in the second item above, we notice that two distinct tables defined over the same universe and the same set of functional dependencies can have the same sets of true tuples.
-Databases such as $T$ and $R$ above are said to be equivalent and when we refer to a table $T$ we in fact refer to any table $\widetilde{T}$ such that ${\sf True}(\mathcal{T})={\sf True}(\mathcal{\widetilde{T}})$.
-The following basic theorem has been shown in LauS25 to hold whenever the set $FD$ is acyclic.
+[…]
+Moreover, the following basic theorem has been shown in LauS25 to hold whenever the set $FD$ is acyclic.
 Theorem 4.1
@@
 $\Box$
+Coming back to star-tables, the set ${\sf Cons}(\mathcal{T})$ satisfies $FD$, because as seen earlier, the set $FD$ of functional dependencies is acyclic.
+We thus define repairs of star-tables as follows.
+Definition 2
+Let $T$ be a star-table over universe $U$. A repair $R$ of $T$ is a table over $U$ such that:
+  1. 1.
+    ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{T})$.
+  2. 2.
+    $R\models FD$.
+  3. 3.
+    For every table $R^{\prime}$ satisfying 1, and 2 above, and such that ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}^{\prime})$.
+The set of all repairs of $T$ is denoted by ${\sf Rep}(T)$. $\Box$
+It should be clear that all results recalled just above concerning acyclic sets of functional dependencies hold in the case of star-tables.
 4.2 Repairs of Star-Tables
 In this section, we show that, in the case of star-tables, repairs satisfy important properties that don’t hold in general.
-Another important specificity of star-tables is that our notion of repair coincides with that in the literature, applied to ${\sf True}(\mathcal{T})$.
+An important specificity of star-tables is that our notion of repair coincides with that in the literature, applied to ${\sf True}(\mathcal{T})$.
 More precisely, we show that if $T$ is a star-table and $S$ is a maximal consistent subset of ${\sf True}(\mathcal{T})$, then $S$ is a repair of $T$ in the sense of Definition 2, implying in particular that ${\sf True}(\mathcal{S})$ contains ${\sf Cons}(\mathcal{T})$.
@@
 Let $T$ be a star-table over universe $U$, $S$ a subset of ${\sf True}(\mathcal{T})$ and $t$ a tuple in ${\sf Cons}(\mathcal{T})$. If $S\models FD$ then $S\cup\{t\}\models FD$.
-Proof.
-Since $S\models FD$, every $\sigma$ in the m-table $m\_Chase(S)$ is such that, for every $A$ in $sch(\sigma)$, $\sigma(A)=(a)$.
-Thus $m\_Chase(S)$ can be seen as a table that we denote by $S^{*}$.
-Moreover, denoting by $S_{t}$ the table $S\cup\{t\}$, we have $m\_Chase(S_{t})=m\_Chase(S^{*}\cup\{t\})$.
-Let $S^{*}_{t}=S^{*}\cup\{t\}$, and let us consider the computation of $m\_Chase(S^{*}_{t})$.
-To this end, given $q_{1}$ and $q_{2}$ in $S^{*}_{t}$ and $X\to A$ in $FD$ such that $q_{1}.X=q_{2}.X$, the only possible cases are as follows:
--- If $q_{1}$ and $q_{2}$ are in $S^{*}_{t}$, then either $q_{1}$ and $q_{2}$ are not defined over $A$ or, they are both defined such that $q_{1}.A=q_{2}.A$.
-In this case $m\_Chase$ does not change $S^{*}_{t}$.
--- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are not defined over $A$, then $m\_Chase$ does not change $S^{*}_{t}$.
--- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are defined over $A$.
-Since $t\in{\sf Cons}(\mathcal{T})$, it is not possible that $q_{1}.A\neq t.A$.
-Thus in this case again, $m\_Chase$ does not change $S^{*}_{t}$.
--- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is defined over $A$ and $q_{2}$ is not defined over $A$. Then $m\_Chase$ changes $t$ into $ta$ where $a=q_{1}.A$.
--- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is not defined over $A$ and $q_{2}$ is defined over $A$. Then $m\_Chase$ changes $q_{1}$ into $qa$ where $a=t.A$.
-Based on the previous cases, we denote by $\Sigma$ the table obtained from $S^{*}_{t}$ by the following transformations:
-$(1)$ For every $q$ in $S^{*}$ and every $X\to A$ in $FD$ such that $q.X=t.X$, $q$ is not defined over $A$ and $t.A=a$, in $\Sigma$, $q.A$ is set to $a$.
-$(2)$ For every $X\to A$ in $FD$ such that $q.X=t.X$, $q.A=a$ and $t$ is not defined over $A$, in $\Sigma$, $t.A$ is set to $a$.
-Since in a star schema, for every attribute $A$ in $U$ there is at most one functional dependency whose right hand-side is $A$, the construction of $\Sigma$ cannot generate conflicts, thus entailing that $\Sigma$ contains no conflicts.
-We now show that $m\_Chase(\Sigma)=\Sigma$.
-Indeed, let $q_{1}$ and $q_{2}$ be in $\Sigma$ and $X\to A$ be in $FD$.
-The only possibility for $m\_Chase$ to change $\Sigma$ is that $q_{1}.X=q_{2}.X$, $q_{1}.A=a$ and $q_{2}$ is not defined over $A$.
-As this does not happen in $S^{*}$ (because $m\_Chase(S^{*})=S^{*}$), and as this does not happen in $\Sigma$ (by definition of $\Sigma$), we obtain that $m\_Chase(\Sigma)=\Sigma$.
-It follows that $m\_Chase(S^{*}_{t})=\Sigma$, and since we have seen that no conflicts occur in $\Sigma$, $S^{*}_{t}\models FD$ holds.
-The proof is therefore complete.
-$\Box$
-The following lemma is a consequence of Lemma 1.
+Proof. See Appendix B. $\Box$
 Lemma 2
@@
 $\Box$
-As a consequence of Proposition 3, we show that in the case of a star-table, given a true tuple $t$, there exists a repair in which this tuple is true, and that moreover, if this tuple is conflicting, there exists a repair in which this tuple is not true.
+We emphasize that, Proposition 4 shows that, if ${\sf True}(\mathcal{T})$ is seen as a table over $U$, then a repair $R$ in ${\sf Rep}(T)$ can be seen as a repair of ${\sf True}(\mathcal{T})$, in the sense of traditional approaches ArenasBC99 .
+On the other hand, we show below that Proposition 4 implies that, in the case of a star-table, given a true tuple $t$, there exists a repair in which $t$ is true, and that moreover, if $t$ is conflicting, there exists a repair in which $t$ is not true.
 It should be noticed here that, the second result is shown in LauS25 , whereas the first one is not.
@@
 Thus the proof of the first statement of the proposition is complete.
-The second statement being the subject of Proposition 4 of LauS25 , the proof of the proposition is complete.
+The second statement of this proposition has been proved in LauS25 as Proposition 4.
+Therefore the proof of this proposition is complete.
 $\Box$
@@
 Let $T$ be a star-table over universe $U$.
-For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{k}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$.
-Then, for every $R$ in ${\sf Rep}(T)$, there exists $i_{0}$ in $\{1,\ldots,k\}$ such that $xa_{i_{0}}$ is in ${\sf True}(\mathcal{R})$.
-Proof.
-It is easy to show that the proposition holds for $k=1$, since in this case $xa_{1}$ is in ${\sf Cons}(\mathcal{T})$, and thus in ${\sf True}(\mathcal{R})$ for every $R$ in ${\sf Rep}(T)$.
-We now assume that $k>1$, which implies that for every $\sigma$ in $m\_Chase(T)$ such that $XA\subseteq sch(\sigma)$ and $x$ is in ${\sf tuples}(\sigma(X))$, ${\sf tuples}(\sigma(XA))=(xa_{1},\ldots,xa_{k})$ with $k>1$.
-Moreover, given the form of functional dependencies in $FD$, either $(a)$ $X=\mathbf{K}$ and $A$ is in $\mathbf{M}$, or $(b)$ there exists $i_{0}$ in $\{1,\ldots,n\}$ such that $X=K_{i_{0}}$ and $A$ is in $sch^{*}(D_{i_{0}})$.
-Let $R$ be in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no $xa_{l}$ from $\{xa_{1},\ldots,xa_{k}\}$, and let $l_{0}$ in $\{1,\ldots,k\}$.
-Denoting by $R^{\prime}$ the table $R\cup\{xa_{l_{0}}\}$ we show that in either case $(a)$ or $(b)$ above, $R^{\prime}\models FD$.
-To this end, we first notice that, since for every $k$ over $\mathbf{K}$ and for any $k_{i}$ over $K_{i}$, $k$ and $k_{i}$ are in ${\sf Cons}(\mathcal{T})$, the tuple $x$ is in ${\sf True}(\mathcal{R})$.
-Hence, ${\sf True}(\mathcal{R})$ contains tuples whose $X$-value is $x$ and for every such tuple, the $A$-value is missing.
-The cases $(a)$ and $(b)$ are then handled as follows when computing $m\_Chase(R^{\prime})$: every $t$ in ${\sf True}(\mathcal{R}^{\prime})$ such that $t.X=x$ is considered and its $A$-value is set to $a_{l_{0}}$.
-Denoting by $\Sigma$ the resulting m-table and let $X^{\prime}\to A^{\prime}$ be a dependency in $FD$, and let $\sigma_{1}$ and $\sigma_{2}$ be two m-tuples in $\Sigma$ such that $\sigma_{1}(X^{\prime})=\sigma_{2}(X^{\prime})$ where $X^{\prime}$ is either $\mathbf{K}$ or $K_{j}$.
--- If $X^{\prime}=X$ then by definition of $\Sigma$, $\sigma_{1}(A)=\sigma_{2}(A)=(a_{l_{0}})$. Thus $\Sigma$ is not changed.
--- If $X^{\prime}=K_{j}$, $X=K_{i}$ and $i\neq j$, $\sigma_{1}$ and $\sigma_{2}$ are both in $R^{\prime}$ and their $A^{\prime}$-components have not been changed when building $\Sigma$ (because $K_{j}\neq K_{i}$ implies $A^{\prime}\neq A$).
-Since $m\_Chase({\sf True}(\mathcal{R}^{\prime}))={\sf True}(\mathcal{R}^{\prime})$ and $R^{\prime}\models FD$, either $A^{\prime}$ is in $sch(\sigma_{1})\cap sch(\sigma_{2})$ and $\sigma_{1}(A^{\prime})=\sigma_{2}(A^{\prime})$, or $A^{\prime}$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$.
-In either case, $\Sigma$ is not changed.
--- If $X^{\prime}=K_{j}$ and $X=\mathbf{K}$, since as above we have $A^{\prime}\neq A$, $m\_Chase$ does not change $\Sigma$.
--- If $X^{\prime}=\mathbf{K}$ and $X=K_{i}$, then in ${\sf True}(\mathcal{R}^{\prime})$, we have $\sigma_{1}(X)=\sigma_{2}(X)=x$, and $A$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$.
-By definition of $\Sigma$, $\sigma_{1}(A)$ and $\sigma_{2}(A)$ are set to $(a_{l_{0}})$, and then here again, $m\_Chase$ does not change $\Sigma$.
-We thus obtain that $\Sigma=m\_Chase(\mathcal{R}^{\prime})$ and that $R^{\prime}\models FD$, because $R\models FD$ and no conflicts occur in $\Sigma$.
-Moreover, as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}^{\prime})$.
-It follows from Definition 2 that $R$ can not be in ${\sf Rep}(T)$ since the inclusion ${\sf True}(\mathcal{R})\subset{\sf True}(\mathcal{R}^{\prime})$ is strict (because $xa_{l_{0}}$ is in ${\sf True}(\mathcal{R}^{\prime})$ but not in ${\sf True}(\mathcal{R})$).
-As this is a contradiction with our hypothesis that $R$ is in ${\sf Rep}(T)$, the proof is complete.
-$\Box$
-We now show how to characterize all repairs of a star-table $T$, based on $m\_Chase(T)$. To this end, we define the following process (P):
+For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{q}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$.
+Then, for every $R$ in ${\sf Rep}(T)$, there exists $q_{0}$ in $\{1,\ldots,q\}$ such that $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R})$.
+Proof. See Appendix C. $\Box$
+We now show how to characterize all repairs of a star-table $T$, based on m_Chase(T)m\_Chase(T). To this end, we define the following three-step process (P):
 Step P1.
-For every $i$ in $\{1,\ldots,n\}$ and every $A_{i}^{j}$ in $sch^{*}(D_{i})$, if $k_{i}$ is a $K_{i}$-value occurring in $m\_Chase(T)$, we choose one $A_{i}^{j}$-value in the set $\sigma(A_{i}^{j})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$.
-Denoting by $\varphi_{i}^{j}(k_{i})$ this value, we notice that $\varphi_{i}^{j}(k_{i})$ is well defined because, thanks to $K_{i}\to A_{i}^{j}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$ have the same value over $A_{i}^{j}$.
+For every $i$ in $\{1,\ldots,n\}$ and every $A_{i}^{j}$ in $sch^{*}(D_{i})$, if $k_{i}$ is a $K_{i}$-value occurring in m_Chase(T)m\_Chase(T), we choose one $A_{i}^{j}$-value in the set $\sigma(A_{i}^{j})$ where $\sigma$ is any m-tuple in m_Chase(T)m\_Chase(T) such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$.
+Denoting by $\varphi_{i}^{j}(k_{i})$ this value, we notice that $\varphi_{i}^{j}(k_{i})$ is well defined because, thanks to $K_{i}\to A_{i}^{j}$ in $FD$, all $\sigma$ in m_Chase(T)m\_Chase(T) such that $K_{i}A_{i}^{j}\subseteq sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$ have the same value over $A_{i}^{j}$.
 Step P2.
-For every $l$ in $\{1,\ldots,p\}$, if $k$ is a $\mathbf{K}$-value occurring in $m\_Chase(T)$, we choose one $M_{l}$-value in the set $\sigma(M_{l})$ where $\sigma$ is any m-tuple in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$.
-Denoting by $\varphi_{l}(k)$ this value, we notice that, as above, $\varphi_{l}(k)$ is well defined because, thanks to $\mathbf{K}\to M_{l}$ in $FD$, all $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$ have the same value over $M_{l}$.
+For every $l$ in $\{1,\ldots,p\}$, if $k$ is a $\mathbb{K}$-value occurring in m_Chase(T)m\_Chase(T), we choose one $M_{l}$-value in the set $\sigma(M_{l})$ where $\sigma$ is any m-tuple in m_Chase(T)m\_Chase(T) such that $\mathbb{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$.
+Denoting by $\varphi_{l}(k)$ this value, we notice that, as above, $\varphi_{l}(k)$ is well defined because, thanks to $\mathbb{K}\to M_{l}$ in $FD$, all $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}M_{l}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$ have the same value over $M_{l}$.
 Step P3.
-Denoting by $\varphi$ the set of all $\varphi_{i}^{j}(k_{i})$ and all $\varphi_{l}(k)$ defined above, for every $\sigma$ in $m\_Chase(T)$, let $t_{\varphi}(\sigma)$ be the tuple such that $sch(t_{\varphi}(\sigma))=sch(\sigma)$, defined by:
+Denoting by $\varphi$ the set of all $\varphi_{i}^{j}(k_{i})$ and all $\varphi_{l}(k)$ defined above, for every $\sigma$ in m_Chase(T)m\_Chase(T), let $t_{\varphi}(\sigma)$ be the tuple such that $sch(t_{\varphi}(\sigma))=sch(\sigma)$, defined by:
   • For every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ and $\sigma(K_{i})=(k_{i})$, then $t_{\varphi}(\sigma).K_{i}=k_{i}$ and for every $j$ such that $A_{i}^{j}\in sch(\sigma)$, $t_{\varphi}(\sigma).A_{i}^{j}=\varphi_{i}^{j}(k_{i})$.
-  • If $\mathbf{K}\subseteq sch(\sigma)$ and $\sigma(\mathbf{K})=(k)$, for every $l=1,\ldots,p$ such that $M_{l}\in sch(\sigma)$, let $t_{\varphi}(\sigma).M_{l}=\varphi_{l}(k)$.
+  • If $\mathbb{K}\subseteq sch(\sigma)$ and $\sigma(\mathbb{K})=(k)$, for every $l=1,\ldots,p$ such that $M_{l}\in sch(\sigma)$, let $t_{\varphi}(\sigma).M_{l}=\varphi_{l}(k)$.
 Denoting by $R_{\varphi}$ the table $R_{\varphi}=\{t_{\varphi}(\sigma)~|~\sigma\in m\_Chase(T)\}\cup{\sf Cons}(\mathcal{T})$, the following proposition states that $R_{\varphi}$ is a repair and that all repairs are obtained through such process.
@@
 Let $T$ be a star-table over $U$. $R$ is a repair of $T$ if and only if there is $\varphi$ as defined above such that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$.
-Proof. See Appendix A. $\Box$
+Proof. See Appendix D. $\Box$
 We illustrate repairs of star-tables in the context of our running Example 3 as follows.
@@
 Example 5
-In the context of Example 3, $T$ has four repairs as shown in Figure 3. It can be seen from the m-table $m\_Chase(T)$ in Figure 2 that, in each of these four tables:
-  • the first two rows are generated using the first m-tuple in $m\_Chase(T)$, namely
+In the context of Example 3, $T$ has four repairs as shown in Figure 4. It can be seen from the m-table m_Chase(T)m\_Chase(T) in Figure 2 that, in each of these four tables:
+  • the first two rows are generated using the first m-tuple in m_Chase(T)m\_Chase(T), namely
     $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$,
-  • the next five rows are generated using the second m-tuple in $m\_Chase(T)$, namely
+  • the next five rows are generated using the second m-tuple in m_Chase(T)m\_Chase(T), namely
     $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$,
-  • the last two rows are generated using respectively the last two m-tuples in $m\_Chase(T)$, namely $(k^{\prime}_{1})(k^{\prime\prime}_{2})(a_{1})(m_{1})$ and $(k^{\prime\prime}_{1})(a^{\prime}_{1})(a^{\prime}_{2})$.
+  • the last two rows are generated using respectively the last two m-tuples in m_Chase(T)m\_Chase(T), namely $(k^{\prime}_{1})(k^{\prime\prime}_{2})(a_{1})(m_{1})$ and $(k^{\prime\prime}_{1})(a^{\prime}_{1})(a^{\prime}_{2})$.
 To illustrate how process (P) works, we apply it to generate the repair $R_{1}$ shown in Figure 4.
@@
   • Step P2.
-The only measure attribute is here $M_{1}$ and the $\mathbf{K}$-values to be considered are $k_{1}k_{2}$, $k_{1}k^{\prime}_{2}$ and $k^{\prime}_{1}k^{\prime\prime}_{2}$ which are respectively associated with $M_{1}$-values $m_{1}$, $m^{\prime\prime}_{1}$ and $m_{1}$.
+The only measure attribute is here $M_{1}$ and the $\mathbb{K}$-values to be considered are $k_{1}k_{2}$, $k_{1}k^{\prime}_{2}$ and $k^{\prime}_{1}k^{\prime\prime}_{2}$ which are respectively associated with $M_{1}$-values $m_{1}$, $m^{\prime\prime}_{1}$ and $m_{1}$.
   • Step P3.
-Considering every m-tuple $\sigma$ of the table $m\_Chase(T)$ shown in Figure 2, we obtain the following four tuples: $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$ and $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
-We obtain that $R_{\varphi}$ contains the four tuples above along with the tuples in ${\sf Cons}(\mathcal{T})$ that have been characterized in Example 3.
+Considering every m-tuple $\sigma$ of the table m_Chase(T)m\_Chase(T) shown in Figure 2, we obtain the following four tuples: $k_{1}k_{2}a_{1}a_{2}b_{1}b_{2}m_{1}$, $k_{1}k^{\prime}_{2}a_{1}a_{2}b_{1}m^{\prime}_{1}$, $k^{\prime}_{1}k^{\prime\prime}_{2}a_{1}m_{1}$ and $k^{\prime\prime}_{1}a^{\prime}_{1}a^{\prime}_{2}$.
+Hence, $R_{\varphi}$ contains the four tuples above along with the tuples in ${\sf Cons}(\mathcal{T})$ that have been characterized in Example 3.
 Moreover, inspecting thoroughly the table $R_{1}$ and the set $R_{\varphi}$ shows that the two sets yield the same set of true tuples, corresponding to the same repair of $T$.
@@
   1. 1.
-    Standard queries or simply queries, to be defined in the next section.
+    Standard queries (or simply queries), to be defined in the next section.
 These queries are simple projection-selection queries of the form select $X$ from $T$ where $\Gamma$, where $X$ is a list of attributes and $\Gamma$ an optional selection condition.
   2. 2.
-    Analytic queries with no group-by statement, defined in Section 5.3.
-These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$, where $X$ and $\Gamma$ are as above, and where $aggr$ is an aggregation operator and $M_{i}$ is a measure attribute.
+    Analytic queries with no group-by clause, defined in Section 5.3.
+These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$, where $X$ and $\Gamma$ are as above, and where aggraggr is an aggregation operator and $M_{i}$ is a measure attribute.
   3. 3.
-    Analytic queries with a group-by statement, also defined in Section 5.3.
-These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above.
+    Analytic queries with a group-by clause, also defined in Section 5.3.
+These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$, where $X$, $\Gamma$, aggraggr and $M_{i}$ are as above.
   4. 4.
-    Analytic queries with a group-by-having statement, dealt with in Section 6.4.
-These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$, where $X$, $\Gamma$, $aggr$ and $M_{i}$ are as above, and where $\Upsilon$ is a boolean expression involving aggregates.
+    Analytic queries with a group-by-having clause, dealt with in Section 6.4.
+These queries are of the form select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$, where $X$, $\Gamma$, aggraggr and $M_{i}$ are as above, and where $\Upsilon$ is a boolean expression involving aggregates.
 5.1 Standard Queries
@@
 The set of all attributes occurring in $\Gamma$ is called the schema of $\Gamma$, denoted by $sch(\Gamma)$; and the attribute set $X\cup sch(\Gamma)$ is called the schema of $Q$, denoted by $sch(Q)$.
+Moreover, in what follows, given an attribute set $X$, $\mathcal{T}(X)$ denotes the set of all tuples in $\mathcal{T}$ whose schema is $X$.
 A selection condition $\Gamma$ is a well-formed formula involving the usual connectors $\neg$, $\vee$ and $\wedge$ and built up from atomic Boolean comparisons of one of the following forms: $A\,\theta\,a$ or $A\,\theta\,A^{\prime}$, where $\theta$ is a comparison predicate, $A$ and $A^{\prime}$ are attributes in $U$ whose domain elements are comparable through $\theta$, and $a$ is in $dom(A)$.
-Given a selection condition $\Gamma$, we denote by $Sat(\Gamma)$ the set of all tuples in $\mathcal{T}(sch(\Gamma))$ satisfying $\Gamma$, as defined below:
+Given a selection condition $\Gamma$, the set of tuples satisfying $\Gamma$, denoted by $Sat(\Gamma)$, is inductively defined as follows:
 $\bullet$ if $\Gamma$ is of the the form $A\,\theta\,a$, $Sat(\Gamma)=\{t\in\mathcal{T}(sch(\Gamma))~|~t.A\,\theta\,a\}$,
@@
 $\Box$
-Moreover, based on Proposition 1, it is shown in LauS25 that the two bounds ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ can be easily computed by means of one scan of the m-table $m\_Chase(T)$.
+Moreover, based on Proposition 1, it is shown in LauS25 that the two bounds ${\sf C\_ans}^{-}(Q)$ and ${\sf C\_ans}^{+}(Q)$ can be easily computed by means of one scan of the m-table m_Chase(T)m\_Chase(T).
 It turns out that approximating the consistent answer to any query in our approach is polynomial.
@@
 5.2 Consistent Query Answering and Star Schemas
-As mentioned in our introductory section, we can compute efficiently exact consistent answers (instead of bounds of exact consistent anwers as in LauS25 ) if the selection condition meets two requirements.
-The first requirement is that satisfaction of the selection condition can be tested attribute by attribute, as stated in the following definition.
+As mentioned in our introductory section, we can compute efficiently exact consistent answers (instead of bounds of exact consistent answers as in LauS25 ) if the selection condition is independent, as stated in the following definition.
 Definition 4
@@
 By Definition 4, if $\Gamma=\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ is independent, then $sch(\Gamma)=A_{1}\ldots A_{k}$ and a tuple $\gamma$ over $sch(\Gamma)$ is in $Sat(\Gamma)$ if and only if for every $i=1,\ldots,k$, $\gamma.A_{i}$ is in $Sat(\Gamma(A_{i}))$.
-For example, $\Gamma=(A_{1}\leq 0)\vee((A_{1}\geq 4)\wedge((A_{2}=a))$ can be written as $((A_{1}\leq 0)\vee(A_{1}\geq 4))\wedge((A_{1}\leq 0)\vee(A_{2}=a))$.
-Thus, $\Gamma$ is not independent.
+For example, let $\Gamma=(A_{1}\leq 0)\vee((A_{1}\geq 4)\wedge(A_{2}=a))$.
+Clearly, the conjunctive normal form of $\Gamma$ is $((A_{1}\leq 0)\vee(A_{1}\geq 4))\wedge((A_{1}\leq 0)\vee(A_{2}=a))$.
+Due to the second conjunct, it turns out that $\Gamma$ cannot be written as the conjunction of disjunctions involving only one attribute.
+Therefore, $\Gamma$ is not independent.
 On the other hand, $\Gamma^{\prime}=((A_{1}\leq 4)\vee(A_{1}\geq 0))\wedge(A_{2}=a)$ is obviously independent with $\Gamma^{\prime}(A_{1})=(A_{1}\geq 4)\vee(A_{1}\leq 0)$ and $\Gamma^{\prime}(A_{2})=(A_{2}=a)$, and we have $Sat(\Gamma^{\prime})=\{(a_{1},a_{2})\in adom(A_{1})\times adom(A_{2})~|~(a_{1}\leq 0\vee a_{1}\geq 4)\wedge(a_{2}=a)\}$.
@@
 Therefore, and contrary to most existing approaches to consistent query answering ArenasBC99 ; Wijsen09 , our approach does not restrict selection conditions to be conjunctive.
-Another restriction that we consider on queries is that key attributes in $\mathbf{K}$ are not allowed in selection conditions.
-We point out that this syntactical restriction on queries does not rule out queries of practical interest, because in most relevant queries, selections are expressed using meaningful attributes rather than numerical abstract key-values.
-Notice in this respect that key attributes are mainly used in selection conditions to express join conditions, which are not explicit in our chase-based approach.
-  • From now on, it is assumed, even when not explicitly mentioned, that selection conditions are independent and involve no key attributes from $\mathbf{K}$.
-Actually, what the above assumption implies is that, in a star-table, given a query $Q:{\tt select}$ $X$ from $T$ where $\Gamma$, for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$ holds, we have $Y\subseteq X\setminus sch(\Gamma)$.
+  • From now on, it is assumed, even when not explicitly mentioned, that selection conditions are independent.
 The following proposition states one of the main contributions of the paper, namely that, under the restriction above, the consistent answer to any query having an independent selection condition can be easily computed.
 Proposition 9
-Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition.
-Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
+Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition.
+Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that
   1. 1.
@@
   2. 2.
-    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
-  3. 3.
-    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
-Proof. See Appendix B. $\Box$
-An important consequence of Proposition 8 is that the data complexity of consistent query answering to projection-selection-join queries on star schemes is polynomial, under the restriction that the selection condition is independent and involves no keys.
-The following corollary shows that the third item in Proposition 8 can be more simply stated when $\mathbf{K}$ is a subset of $X$.
+    for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$:
+    1. (a)
+      if $B\in X$, then $|{\sf tuples}(\sigma(B))|=1$,
+    2. (b)
+      if $B\in sch(\Gamma)$, then ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
+Proof. See Appendix E. $\Box$
+An important consequence of Proposition 9 is that the time complexity of consistent query answering to projection-selection-join queries on star schemes is quasi-linear (i.e., in $\mathcal{O}(W\cdot\log(W))$, where $W$ is the size of the data warehouse), under the restriction that the selection condition is independent.
+The following corollary shows that the items in Proposition 9 can be more simply stated when $\mathbb{K}$ is a subset of $X$.
 Corollary 1
-Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition.
-If $\mathbf{K}\subseteq X$, ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
+Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition.
+If $\mathbb{K}\subseteq X$, ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that
   1. 1.
-    $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$,
+    $sch(Q)\subseteq sch(\sigma)$,
   2. 2.
-    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
+    $|{\sf tuples}(\sigma(X))|=1$, i.e., ${\sf tuples}(\sigma(X))=\{x\}$,
   3. 3.
@@
 Proof.
-As for every $A$ in $U$, there exists $Y\to A$ in $FD$ with $Y\subseteq\mathbf{K}$, if $\mathbf{K}\subseteq X$, for every $B$ in $sch(\Gamma)$ there exists $Y\to B$ in $FD$ with $Y\subseteq X$.
-Therefore, in this case, item 3 in Proposition 8 can be stated as: for every $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B)$.
+The inclusion of $\mathbb{K}$ in $X$ implies that for every $B$ in $sch(Q)\setminus\mathbb{K}$, there exists $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$.
+Considering moreover that for every $Y\subseteq\mathbb{K}$, $|{\sf tuples}(\sigma(Y))|=1$ always holds, items 1 and 2(a) in Proposition 9 are equivalent to item 2 in the corollary.
+Similarly, for every attribute $B$ in $sch(\Gamma)$, either $B$ is in $\mathbb{K}$ or there exists $Y\to B$ in $FD$ such that $Y\subseteq\mathbb{K}$.
+Thus, considering the latter case, item 2(b) in Proposition 9 is equivalent to: for every $B\in sch(\Gamma)\setminus\mathbb{K}$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
+On the other hand if $B$ is in $\mathbb{K}$, we have ${\sf tuples}(\sigma(B))=(b)$ where $b$ is in $Sat(\Gamma(B))$, by item 1 in Proposition 9.
+Hence item 1 and item 2(b) imply that for every $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B)\subseteq Sat(\Gamma(B))$.
 Since $\Gamma$ is independent, as mentioned right after Definition 4, this is equivalent ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
@@
 $\Box$
-In the following example, we illustrate Proposition 8 and Corollary 1, and we show the impact of our two restrictions on the form of the selection conditions.
+In the following example, we illustrate Proposition 9 and Corollary 1, along with the importance of our restriction on the selection conditions.
 Example 6
-In the context of Example 3, we consider the star-table $T$, and its associated m-table $m\_Chase(T)$ as shown in Figure 2, let $Q_{2}$ be the query defined by:
+In the context of Example 3, we consider the star-table $T$, and its associated m-table m_Chase(T)m\_Chase(T) as shown in Figure 2, let $Q_{2}$ be the query defined by:
 $Q_{2}:$ select $K_{1}$, $K_{2}$, $A_{2}^{1}$, $M_{1}$, from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$).
 As already noticed after stating Proposition 8, we have ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$.
-On the other hand, the selection condition in this query obviously satisfies our restrictions, since it is independent and involves no key attributes.
-Applying Proposition 8 or Corollary 1 (because $K_{1}K_{2}\subseteq sch(Q_{2})$) yields the following:
-  • The m-tuples $\sigma$ in $m\_Chase(T)$ such that $sch(Q_{i})\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$, are $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ and $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$.
-  • Among these m-tuples, only $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ is such that for every $Y\to B$ in $FD$ such that $YB\subseteq X_{i}$, $|{\sf tuples}(\sigma(B))|=1$.
-  • This m-tuple satisfies ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
+As the selection condition in this query is obviously independent, applying Proposition 9 or Corollary 1 (because $K_{1}K_{2}\subseteq sch(Q_{2})$) yields the following:
+  • The m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) such that $sch(Q_{i})\subseteq sch(\sigma)$, $x\in\sigma(X)$ and $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$, are $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ and $(k_{1})(k^{\prime}_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(m^{\prime}_{1}\,m^{\prime\prime}_{1})$.
+  • Among these m-tuples, only $(k_{1})(k_{2})(a_{1}\,a^{\prime}_{1})(a_{2})(b_{1})(b_{2})(m_{1})$ is such that: for every $Y\to B$ in $FD$ such that $YB\subseteq X_{i}$, $|{\sf tuples}(\sigma(B))|=1$.
+  • The unique m-tuple $\sigma$ of the previous item is such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
 We therefore obtain the expected consistent answer, namely ${\sf C\_ans}(Q_{2})=\{k_{1}k_{2}b_{1}m_{1}\}$. Now, let $Q_{3}$ be defined by:
-$Q_{3}:$ select $K_{1},K_{2},A^{2}_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ and $M_{1}=m_{1}$) or ($A_{1}^{1}=a^{\prime}_{1}$ and $M_{1}=m^{\prime\prime}_{1}$).
+$Q_{3}:$ select $K_{1},K_{2},A^{2}_{1}$ from $T$ where
+($A_{1}^{1}=a_{1}$ and $M_{1}=m_{1}$) or ($A_{1}^{1}=a^{\prime}_{1}$ and $M_{1}=m^{\prime\prime}_{1}$).
 In this case, the selection condition $\Gamma_{3}$ is clearly not independent, and it can be seen from Figure 4 that ${\sf C\_ans}(Q_{3})$ is in fact empty, because:
@@
 On the other hand, considering the third item in Proposition 9, we emphasize the following: $k_{1}k^{\prime}_{2}a_{2}$ is discarded from the consistent answer because $a_{1}m^{\prime}_{1}$ is in ${\sf tuples}(\sigma_{2}(A_{1}^{1}M_{1}))$ but not in $Sat(\Gamma_{3})$.
-However, $\sigma_{1}$ can be seen as satisfying the third item in the proposition for the following reasons: the dependencies to consider are $K_{1}\to A_{1}^{1}$ and $K_{1}K_{2}\to M_{1}$, and ${\sf tuples}(\sigma_{1}(A^{1}_{1}))$, respectively ${\sf tuples}(\sigma_{1}(M_{1}))$, is a sub-set of the set of all $A_{1}^{1}$-values, respectively all $M_{1}$-values, occurring in $Sat(\Gamma_{3})$.
+However, $\sigma_{1}$ can be seen as satisfying the third item in the proposition for the following reasons: the dependencies to consider are $K_{1}\to A_{1}^{1}$ and $K_{1}K_{2}\to M_{1}$, and ${\sf tuples}(\sigma_{1}(A^{1}_{1}))$, respectively ${\sf tuples}(\sigma_{1}(M_{1}))$, is a subset of the set of all $A_{1}^{1}$-values, respectively all $M_{1}$-values, occurring in $Sat(\Gamma_{3})$.
 A characterization of the consistent answer to a query involving a non independent selection condition is currently unknown to the authors.
-Now, to illustrate our restriction on the key-attributes with respect to selection conditions, let $Q_{4}:$ select $A_{1}$ from $T$ where ($K_{1}=k_{1}$).
-It should be clear that the selection condition of $Q_{4}$ is independent, and that ${\sf C\_ans}(Q_{4})=\emptyset$ (because, by Figure 3, neither $a_{1}$ nor $a^{\prime}_{1}$ is associated with $k_{1}$ is every repair).
-On the other hand, it is easy to see that the conditions in Proposition 8 or Corollary 1 are satisfied by the first two m-tuples of $m\_Chase(T)$ (see Figure 2).
-A characterization of the consistent answer to a query with a selection condition involving a key attribute is currently unknown to the authors.
 $\Box$
@@
 5.3 Analytic Queries and their Consistent Answers
-In the literature Ullman , an analytic query is a query involving an aggregate function among $count$, $min$, $max$, $sum$, and possibly a group-by clause.
-Formally, given a data warehouse over a star schema, an analytic query $\mathcal{AQ}$ has one of the following two forms:
+In the literature Ullman , an analytic query is a query involving an aggregate operator among countcount, minmin, maxmax, sumsum, and possibly a group-by clause.
+Formally, given a data warehouse, an analytic query $\mathcal{AQ}$ has one of the following two forms:
   • $\mathcal{AQ}:{\tt select}$ $aggr(M_{i})$ from $\varphi$ where $\Gamma$, or
   • $\mathcal{AQ}:{\tt select}$ $X$, $aggr(M_{i})$ from $\varphi$ where $\Gamma$ group by $X$
-where $\varphi$ is the join of the fact table $F$ with all dimension tables $D_{1},\ldots,D_{n}$, $X$ is a relation schema, and for $j=1,\ldots,p,$ $aggr$ is an aggregate function such as $count$, $min$, $max$, $sum$ and $M_{i}$ is a measure attribute.
+where $\varphi$ is the join of the fact table $F$ with all dimension tables $D_{1},\ldots,D_{n}$, $X$ is a relation schema, and for $j=1,\ldots,p,$ aggraggr is an aggregate operator such as countcount, minmin, maxmax, sumsum and $M_{i}$ is a measure attribute.
 In all traditional approaches to data warehouses, it is generally assumed that the tables in the data warehouse have no missing values and that the key-foreign-key constraints between $F$ and all dimension tables $D_{i}$ are satisfied.
@@
   • If $\mathcal{AQ}$ involves no group-by clause, ${\sf Ans}(\mathcal{AQ})$ is the value of the aggregation evaluated over the set of all tuples in $\varphi$ satisfying the condition $\Gamma$.
-We notice that the aggregate may involve no attribute, when expressed as $count(*)$.
+We notice that the aggregate may involve no attribute, when expressed as count()count(*).
   • If $\mathcal{AQ}$ involves a statement group-by $X$, ${\sf Ans}(\mathcal{AQ})$ is the set of all pairs $\langle x,v_{x}\rangle$ where $x$ is such that there exists a tuple in $\varphi$ satisfying $\Gamma$ and whose $X$-value is $x$, and where $v_{x}$ is the value of the aggregation evaluated over all tuples in $\varphi$ satisfying $\Gamma$ and whose $X$-value is $x$.
 Now, if the underlying data warehouse does not satisfy all functional dependencies, this traditional semantics of answer to analytic queries has to be revisited.
-When no dependency of the form $\mathbf{K}\to M_{i}$ is considered, this has been done in ArenasBCHRS03 , in the case of queries with no group-by clause and in FuxmanFM05 when a group-by clause occurs in the query.
-As shown in DixitK19 , in either case, the ‘repair semantics’ of the consistent answer to an analytic query consists intuitively in producing an interval in which aggregate values fall, when answering the query in any repair.
+When no dependency of the form $\mathbb{K}\to M_{i}$ is considered, this has been done in ArenasBCHRS03 , in the case of queries with no group-by clause and in FuxmanFM05 when a group-by clause occurs in the query.
+As shown in DixitK22 , in either case, the ‘repair semantics’ of the consistent answer to an analytic query consists intuitively in producing an interval in which aggregate values fall, when answering the query in any repair.
 The reader is referred to the forthcoming Section 7 for a more precise relationship between these approaches and our work.
-In our approach, we follow the same line regarding analytic queries and their consistent answers, and we emphasize that, as in the previous section, all selection conditions are assumed to satisfy the restrictions that are independent and that they involve no attribute from $\mathbf{K}$.
+In our approach, we follow the same line regarding analytic queries and their consistent answers, and we emphasize that, as in the previous section, all selection conditions are assumed to be independent.
 Definition 5
@@
 (A) If $\mathcal{AQ}$ involves no group-by clause, then:
-  • If there exists $R$ in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})={\tt NULL}$ if $aggr\neq count$, and ${\sf C\_ans}(\mathcal{AQ})=0$ if $aggr$ is $count$.
+  • If there exists $R$ in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})={\tt NULL}$ if $aggr\neq count$, and ${\sf C\_ans}(\mathcal{AQ})=0$ if aggraggr is countcount.
   • If for every $R$ in ${\sf Rep}(T)$, ${\sf True}(\mathcal{R})$ contains at least one tuple over $sch(Q)$ satisfying $\Gamma$, then ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$ such that:
@@
 (B) If $\mathcal{AQ}$ involves a group-by clause, ${\sf C\_ans}(\mathcal{AQ})$ is a set of pairs of the form $\langle x,[glb,lub]\rangle$ such that:
-  • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})$ contains $(x\,d)$,
-  • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})$ respectively contain $(x\,glb)$ and $(x\,lub)$. $\Box$
+  • for every $R$ in ${\sf Rep}(T)$ there exists $d$ in $[glb,lub]$ such that ${\sf Ans}(\mathcal{AQ}^{[R]})$ contains (xd)(x\,d),
+  • there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})$ and ${\sf Ans}(\mathcal{AQ}^{[R_{2}]})$ respectively contain $(x,glb)$ and $(x,lub)$. $\Box$
 The following remarks are in order regarding Definition 5:
-  • In case (A), a NULL answer means that the consistent answer to the query $Q:~{}{\tt select}$ $\mathbf{K}$ from $T~{}{\tt where}~{}\Gamma$ is empty.
-In this case the aggregate cannot be computed if different than $count$, and if the aggregate is $count$, then the expected answer is 00.
+  • In case (A), a NULL answer means that the consistent answer to the query $Q:~{\tt select}$ $\mathbb{K}$ from $T~{\tt where}~\Gamma$ is empty.
+In this case the aggregate value cannot be computed if the aggregate operator is different than countcount, and if the aggregate operator is countcount, then the expected answer is 0.
 This explains why we return NULL in the former case.
@@
 Consequently, if no $X$-value fits the conditions in the definition, ${\sf C\_ans}(\mathcal{AQ})$ is simply set to $\emptyset$.
-Additionally to consistent answers as defined above, we argue that our approach allows for another kind of consistent answers to analytic queries, relying on the tuples in ${\sf Cons}(\mathcal{T})$, or by Theorem 4.1 on tuples true in every repair.
-To define this kind of consistent answer, a given analytic query $\mathcal{AQ}:~{}$select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] is associated with a query involving no aggregate, defined by
-$Q_{\mathcal{AQ}}:$ select $\mathbf{K},$[$X$]$,M_{i}$ from $T$ where $\Gamma$.
-Then, relying on the consistent answer to $Q_{\mathcal{AQ}}$, we define the strongly consistent answer to $\mathcal{AQ}$, denoted by ${\sf C\_ans}^{*}(\mathcal{AQ})$, as the answer to the following analytic query:
-select [$X$]$,aggr(M_{i})$ from ${\sf C\_ans}(Q_{\mathcal{AQ}})$ [group-by $X$].
-We emphasize that, based on Corollary 1, since the select clause of $Q_{\mathcal{AQ}}$ contains $\mathbf{K}$, ${\sf C\_ans}(Q_{\mathcal{AQ}})$ can be effectively computed, assuming that $m\_Chase(T)$ is available.
-The important issue of efficiently computing answers to analytic queries is addressed in the next section.
+An alternative way to deal with consistent answers to analytic queries consists of computing first the consistent answer to an appropriate non analytic query and then, of applying the aggregation operators to this answer.
+More formally, a given analytic query $\mathcal{AQ}:$~select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] is associated with the non analytic query defined by
+[…]
+and this alternative consistent answer to $\mathcal{AQ}$ is defined as follows.
+Definition 6
+Let $T$ be a star-table over universe $U$ and let $\mathcal{AQ}$ be an analytic query defined as follows: $\mathcal{AQ}:$~select [$X$], $aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$].
+The strongly consistent answer to $\mathcal{AQ}$, denoted by ${\sf C\_ans}^{*}(\mathcal{AQ})$, is the answer to the analytic query:
+[…]
+where ${\sf C\_ans}(Q_{\mathcal{AQ}})$ is the consistent answer to the (non-analytic) query $Q_{\mathcal{AQ}}$ as defined above. $\Box$
+We emphasize that, since the select clause of $Q_{\mathcal{AQ}}$ contains $\mathbb{K}$, ${\sf C\_ans}(Q_{\mathcal{AQ}})$ can be effectively computed based on Corollary 1, assuming that m_Chase(T)m\_Chase(T) is available.
+Moreover, the semantics of such answers are radically different than those defined in Definition 5, simply because strongly consistent answers do not involve intervals, whereas consistent answers do.
+Moreover, as shown in the following example, aggregate values in the strongly consistent answer may not belong to the corresponding interval of the consistent answer.
 Example 7
-In the context of Example 3, $\mathcal{AQ}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$) is an example of analytic query with no group-by clause and involving the aggregate $\max$.
+In the context of Example 3, $\mathcal{AQ}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$) is an example of analytic query with no group-by clause and involving the aggregate sumsum.
 The expected result is the interval containing all possible sums of $M_{1}$-values among all tuples defined over $K_{1}K_{2}$ and associated with $A_{1}^{1}$-value $a_{1}$ in all repairs.
@@
 $\bullet$ ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{3}]})=m_{1}+m^{\prime\prime}_{1}+m_{1}$, and ${\sf Ans}(\mathcal{AQ}_{1}^{[R_{4}]})=m_{1}$.
-Assuming that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive numbers and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{1})=[glb,lub]$ and where $glb=m_{1}$ and $lub=2.m_{1}+m^{\prime\prime}_{1}$.
-On the other hand, we have $Q_{\mathcal{AQ}_{1}}:~{}$select $\mathbf{K},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), and it can be seen from Figure 3, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{1}})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$.
+Assuming that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive numbers and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{1})=[glb,lub]$ where $glb=m_{1}$ and $lub=2\cdot m_{1}+m^{\prime\prime}_{1}$.
+On the other hand, we have $Q_{\mathcal{AQ}_{1}}:$~select $\mathbb{K},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$), and it can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{1}})=\{k^{\prime}_{1}k^{\prime\prime}_{2}m_{1}\}$.
 Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})=m_{1}$.
-To illustrate that if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=m$, then it does not always hold that $glb\leq m\leq lub$, consider the aggregate $count(*)$ instead of $sum(M_{1})$.
-Indeed, in this case, ${\sf C\_ans}(\mathcal{AQ})=[3,3]$ and ${\sf C\_ans}^{*}(\mathcal{AQ})=1$.
+Notice that in this case, the value in ${\sf C\_ans}^{*}(\mathcal{AQ}_{1})$ belongs to the interval in ${\sf C\_ans}(\mathcal{AQ}_{1})$.
 As another example of analytic query with no group-by clause and slightly different from $\mathcal{AQ}_{1}$, consider $\mathcal{AQ}^{\prime}_{1}:{\tt select}$ $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a^{\prime}_{1}$).
@@
 In this case, by Definition 5(A), the expected consistent answer is NULL.
+On the other hand, we have $Q_{\mathcal{AQ}^{\prime}_{1}}:$~select $\mathbb{K},M_{1}$ from $T$ where ($A_{1}^{1}=a^{\prime}_{1}$), and it can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}^{\prime}_{1}})=\emptyset$.
+Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}^{\prime}_{1})={\tt NULL}$.
 The query $\mathcal{AQ}_{2}:{\tt select}$ $A_{2}^{1}$, $sum(M_{1})$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$) group-by $A_{2}^{1}$ is an analytic query involving a group-by clause.
@@
 Assuming as above that $m_{1}$, $m^{\prime}_{1}$ and $m^{\prime\prime}_{1}$ are positive and that $m^{\prime}_{1}\leq m^{\prime\prime}_{1}$, we obtain that ${\sf C\_ans}(\mathcal{AQ}_{2})=\{\langle b_{1},[glb,lub]\rangle\}$ where $glb=m_{1}+m^{\prime}_{1}$ and $lub=m_{1}+m^{\prime\prime}_{1}$.
-On the other hand, we have $Q_{\mathcal{AQ}_{2}}:~{}$select $K,A_{1}^{2},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$).
+On the other hand, we have $Q_{\mathcal{AQ}_{2}}:$~select $\mathbb{K},A_{1}^{2},M_{1}$ from $T$ where ($A_{1}^{1}=a_{1}$ or $A_{1}^{1}=a^{\prime}_{1}$).
 It can be seen from Figure 4, that ${\sf C\_ans}(Q_{\mathcal{AQ}_{2}})=\{k_{1}k_{2}b_{1}m_{1}\}$.
 Therefore, ${\sf C\_ans}^{*}(\mathcal{AQ}_{2})=\{(b_{1},m_{1})\}$.
-This example illustrates that if $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$ and $x\,m$ is in ${\sf C\_ans}^{*}(\mathcal{AQ})$, then it does not always hold that $glb\leq m\leq lub$.
-Indeed, for $m_{1}=1$, $m^{\prime}_{1}=2$ and $m^{\prime\prime}_{1}=3$, we have $glb=3$, $lub=4$ showing that for $x=b_{1}$, $m\not\in[3,4]$.
+It is important to notice that this case illustrates that if $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$ and xmx\,m is in ${\sf C\_ans}^{*}(\mathcal{AQ})$, then it does not always hold that $glb\leq m\leq lub$.
+Indeed, for $m_{1}=1$, $m^{\prime}_{1}=2$ and $m^{\prime\prime}_{1}=3$, we have $glb=3$, $lub=4$ showing that for $x=b_{1}$, $m\not\in[glb,lub]$.
 $\Box$
@@
 We also consider in this section the case of analytic queries involving a group-by-having clause, whereas the introduction of the clause distinct is the subject of the last sub-section.
-We emphasize again that the selection conditions in queries are independent and involve no key attributes.
+We emphasize again that the selection conditions in queries are assumed to be independent.
 6.1 Basic Result
-Before presenting our algorithms, we state important properties of repairs and analytic queries.
-To this end, we introduce the following additional notation.
-Given a star-table $T$ and an analytic query $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$], let:
-$\bullet$ $\Sigma(\mathcal{AQ})=\{\sigma\in m\_Chase(T)~{}|~{}(\mathbf{K}\cup M_{i}\cup sch(\Gamma)\subseteq sch(\sigma))\wedge(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset)\}$
+[…]
+$\bullet$ $\Sigma(\mathcal{AQ})=\{\sigma\in m\_Chase(T)~|~(\mathbb{K}\cup M_{i}\cup sch(\Gamma)\subseteq sch(\sigma))\wedge(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset)\}$
 $\bullet$ $\Sigma^{+}(\mathcal{AQ})=\{\sigma\in\Sigma(\mathcal{AQ})~|~\sigma(sch(\Gamma))\subseteq Sat(\Gamma)\}$.
 Proposition 10
-Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}:~{}$select [$X$],$\,aggr(M_{i})$ from $T$ where $\Gamma$ [group by $X$] an analytic query where $\Gamma$ is independent, i.e., $\Gamma=\Gamma(A_{1})\wedge\ldots\wedge\Gamma(A_{k})$ where $sch(\Gamma)=A_{1}\ldots A_{k}$.
-If $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, there exist $R_{1}$ and $R_{2}$ in ${\sf Rep}(T)$ such that
+[…]
   1. 1.
-    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1})$ such that $t.sch(\Gamma)\not\in Sat(\Gamma)$,
+    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1}))(t.sch(\Gamma)\not\in Sat(\Gamma))$,
   2. 2.
-    for every $\sigma$ in $\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$, there exists $t$ in ${\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2})$ such that $t.sch(\Gamma)\in Sat(\Gamma)$.
-Proof.
-The proof relies on the fact that for every repair $R$ of $T$ and every $\sigma$ in $\Sigma(\mathcal{AQ})$, ${\sf True}(\mathcal{R})$ contains exactly one tuple $t$ such that $t\in{\sf tuples}(\sigma)$ and $t.K\in\sigma(\mathbf{K})$ (the existence is a consequence of Proposition 4, unicity follows from Proposition 2(1) because $|\sigma(\mathbf{K})|=1$, $\mathbf{K}$ is a key of $R$ and $R\models FD$).
-As a consequence of Proposition 6, the existence of repairs $R_{1}$ and $R_{2}$ is explicitly shown using the process (P).
-1. Regarding $R_{1}$, we build $\varphi$ as follows: For every $\sigma$ in $\Sigma(\mathcal{AQ})$:
-Step P1. For every $A_{p}^{k}$ in $sch^{*}(D_{p})\cap sch(\sigma)$ such that $R_{1}$ contains no tuple $k_{p}a_{p}^{k}$ from ${\sf tuples}(\sigma(K_{p}A_{p}^{k}))$
-$(a)$ if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then choose an $A_{p}^{k}$-value in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
-$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\setminus Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
-If ${\sf tuples}(\sigma(A_{p}^{k}))\subseteq Sat(\Gamma(A_{k}^{p}))$, choose any $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
-It is important to notice that in this case, there always exists at least one attribute $A_{p}^{k}$ such that ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, because otherwise $\sigma$ would be in $\Sigma^{+}(\mathcal{AQ})$.
-Step P2.
-If $M_{i}\not\in sch(\Gamma)$ or if $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma^{+}(\mathcal{AQ})$, choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k_{p})=m$.
-Otherwise (i.e., $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$), choose an $M_{i}$-value $m$ in $\sigma(M_{i})\setminus Sat(\Gamma(M_{i}))$ for $\varphi_{i}(K_{p})$.
-Once all m-tuples in $\Sigma(\mathcal{AQ})$ have been considered, the two steps above are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values as done in the generic description of process (P).
-Consequently, the corresponding set $\varphi$ in process (P) is properly defined.
-Step P3. Build up the tuples $t_{\varphi}(\sigma)$ for every $\sigma$ in $m\_Chase(T)$. We notice that
-$(a)$ if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\in Sat(\Gamma)$,
-$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\not\in Sat(\Gamma)$.
-By Proposition 6, we therefore obtain a repair $R_{1}$ satisfying item 1 in the proposition.
-2. Regarding the item 2, the repair $R_{2}$ is obtained as above, by changing $(b)$ in Step P1 and Step P2 as follows:
-Step P1.
-$(b)$ if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\cap Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
-It is important to notice that in this case, every attribute $A_{p}^{k}$ in $sch(\Gamma)$ is such that ${\sf tuples}(\sigma(A_{p}^{k}))\cap Sat(\Gamma(A_{k}^{p}))\neq\emptyset$, because otherwise ${\sf tuples}(\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ would not be possible.
-Step P2. Choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k)=m$ where $k=\sigma(\mathbf{K})$.
-Then, as above, these two steps are completed by considering all non-processed $K_{i}$- or $\mathbf{K}$-values, and Step P3 is considered.
-By Proposition 6, we obtain a repair $R_{2}$ that satisfies item 2 in the proposition.
-The proof is therefore complete.
-$\Box$
+    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2}))(t.sch(\Gamma)\in Sat(\Gamma))$.
+Proof. See Appendix F. $\Box$
 The following example shows that Proposition 10 does not always hold for non independent selection conditions.
@@
 Let $U=\{K_{1},K_{2},A_{1},A_{2},M_{1}\}$ and $FD=\{K_{1}\to A_{1},K_{2}\to A_{2},K_{1}K_{2}\to M_{1}\}$, and the following tables
-$D_{1}$ $K_{1}$ $A_{1}$ $k_{1}$ $10$ $k^{\prime}_{1}$ $-5$ $D_{2}$ $K_{2}$ $A_{2}$ $k_{2}$ $20$ $k^{\prime}_{2}$ $-1$ $k^{\prime}_{2}$ $30$ $F$ $K_{1}$ $K_{2}$ $M_{1}$ $k_{1}$ $k_{2}$ $-10$ $k_{1}$ $k^{\prime}_{2}$ $2$ $k^{\prime}1$ $k^{\prime}_{2}$ $-100$
-Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
-[…]
+$D_{1}$ $K_{1}$ $A_{1}$ $k_{1}$ $10$ $k^{\prime}_{1}$ 5-5 $D_{2}$ $K_{2}$ $A_{2}$ $k_{2}$ 2020 $k^{\prime}_{2}$ 1-1 $k^{\prime}_{2}$ $30$ $F$ $K_{1}$ $K_{2}$ $M_{1}$ $k_{1}$ $k_{2}$ 10-10 $k_{1}$ $k^{\prime}_{2}$ 22 $k^{\prime}1$ $k^{\prime}_{2}$ 100-100
+Thus m_Chase(T)m\_Chase(T) contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
+$\sigma_{1}=(k_{1})(k_{2})(10)(20)(-10)$, $\sigma_{2}=(k_{1})(k^{\prime}_{2})(10)(-1\,30)(2)$, $\sigma_{3}=(k^{\prime}_{1})(k^{\prime}_{2})(-5)(-1\,30)(-100)$
 and $T$ has two repairs $R_{1}$ and $R_{2}$.
-Moreover, the tuples in ${\sf True}(\mathcal{R}_{i})$ ($i=1,2$) whose schemas contain $\mathbf{K}=K_{1}K_{2}$, respectively denoted $R^{\mathbf{K}}_{1}$ and $R^{\mathbf{K}}_{2}$, are:
-$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,-1,2),(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)\}$
-$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,30,2),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$.
+Moreover, the tuples in ${\sf True}(\mathcal{R}_{i})$ ($i=1,2$) whose schemas contain $\mathbb{K}=K_{1}K_{2}$, respectively denoted $R^{\mathbb{K}}_{1}$ and $R^{\mathbb{K}}_{2}$, are:
+$R^{\mathbb{K}}_{1}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,-1,2),(k^{\prime}_{1},k^{\prime}_{2},-5,-1,-100)\}$
+$R^{\mathbb{K}}_{2}=\{(k_{1},k_{2},10,20,-10),(k_{1},k^{\prime}_{2},10,30,2),(k^{\prime}_{1},k^{\prime}_{2},-5,30,-100)\}$.
 For $\Gamma=(M_{1}\leq A_{1})\wedge(M_{1}\geq 0\Rightarrow A_{2}<0)\wedge(M_{1}<-50\Rightarrow A_{2}>20)$ and $\mathcal{AQ}$ an analytic query involving $\Gamma$, we have $\Sigma(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ and $\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1}\}$.
@@
 $\Box$
-In the following two sections 6.2 and 6.3, we successively introduce our algorithms for computing consistent answers to analytic queries, depending on whether they involve or not a group-by clause.
-In each case, the main algorithm relies on a procedure, called Compute_aggregate and shown in Algorithm 2, whose role is to scan (either entirely or partially) the m-table $m\_Chase(T)$ and return values that will appear in the answers.
-6.2 Analytic Queries with no Group-by Clause
-If the query involves no group-by clause, as shown in Algorithm 3, the procedure Compute_aggregate of Algorithm 2 is called to scan the whole m-table $m\_Chase(T)$.
-When running this call, line 1 in Algorithm 3, the main loop lines 6-32 of Algorithm 2 scans $m\_Chase(T)$ and computes values appearing in the answers.
-The following proposition shows that Algorithm 3 is correct, except when $aggr=sum$.
-In other words, if $aggr\neq sum$ Algorithm 3 returns ${\sf C\_ans}(\mathcal{AQ})$.
-Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 3 in any case.
 Algorithm 3 Procedure Compute_Aggregate
@@
 0:
-$proc\_change$ // a boolean variable
-$proc\_min$ // meant to be equal to the lower bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
-$proc\_max$ // meant to be equal to the upper bound of ${\sf C\_ans}(\mathcal{AQ})$ if $proc\_change$ has value true
-$proc\_ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$ if $proc\_change$ has value true
+proc_changeproc\_change // a boolean variable
+proc_minproc\_min // meant to be equal to the lower bound of ${\sf C\_ans}(\mathcal{AQ})$ if proc_changeproc\_change has value true
+proc_maxproc\_max // meant to be equal to the upper bound of ${\sf C\_ans}(\mathcal{AQ})$ if proc_changeproc\_change has value true
+$proc\_ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$ if proc_changeproc\_change has value true
 1:  if $aggr\neq count$ then
-[…]
+2:    proc_min:=proc\_min:=-dummy ; proc_max:=proc\_max:=+dummy ; $proc\_ans^{*}:=$-dummy// -dummy and +dummy are values not in $adom(M_{i})$ such that for every $m$ in $adom(M_{i})$,// $aggr($-dummy,m)=aggr(,m)=aggr(+dummy,m)=m,m)=m and $\min($-$m)=\max(+dummy$,m)=m,m)=m
 3:  if $aggr=count$ then
-4:      $proc\_min:=0$ ; $proc\_max:=0$ ; $proc\_ans^{*}:=0$
+4:    proc_min:=0proc\_min:=0 ; proc_max:=0proc\_max:=0 ; $proc\_ans^{*}:=0$
 5:  $proc\_change:={\tt false}$
 6:  for all $\sigma$ in $S$ do
-7:      if $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ then
+7:    if $\mathbb{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ and $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ then
 8:     if $aggr\neq count$ then
 9:       if $M_{i}\in sch(\Gamma)$ then
-10:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
-11:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
-12:               else
-13:                    $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i})))$
-14:                    $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i})))$
-15:           if ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$ then
-16:               $proc\_change:={\tt true}$
-17:               $proc\_min:=aggr(min_{\sigma},proc\_min)$
-18:               $proc\_max:=aggr(max_{\sigma},proc\_max)$
+10:        $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$ ; $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i}))\cap Sat(\Gamma(M_{i})))$
+11:       else
+12:        $min_{\sigma}:=\min({\sf tuples}(\sigma(M_{i})))$ ; $max_{\sigma}:=\max({\sf tuples}(\sigma(M_{i})))$
+13:     if ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$ then
+14:       $proc\_change:={\tt true}$
+15:       if $aggr=count$ then
+16:        proc_min:=proc_min+1proc\_min:=proc\_min+1
+17:       else
+18:        $proc\_min:=aggr(min_{\sigma},proc\_min)$ ; $proc\_max:=aggr(max_{\sigma},proc\_max)$
 19:       if ${\sf tuples}(\sigma(M_{i}))=\{m_{0}\}$ then
@@
 26:       if $aggr=count$ then
-27:                    $proc\_max:=proc\_max+1$
+27:        proc_max:=proc_max+1proc\_max:=proc\_max+1
 28:       if $aggr=sum$ then
@@
 32:          $proc\_max:=proc\_max+max_{\sigma}$
-33:  return  $proc\_change$, $proc\_min$, $proc\_max$, $proc\_ans^{*}$
-Proposition 10
-Let $T$ be a star-table over universe $U$ and let $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$ be an analytic query with no group-by clause. Then we have:
-  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
-  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
-    Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$.
-  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
-Proof. See Appendix C. $\Box$
+33:  return proc_changeproc\_change, proc_minproc\_min, proc_maxproc\_max, $proc\_ans^{*}$
+In the following two sections 6.2 and 6.3, we successively introduce our algorithms for computing consistent answers to analytic queries, depending on whether they involve or not a group-by clause.
+In each case, the main algorithm relies on a procedure, called Compute_aggregate and shown in Algorithm 3, whose role is to scan (either entirely or partially) the m-table m_Chase(T)m\_Chase(T) and return values that will appear in the answers.
+6.2 Analytic Queries with no Group-by Clause
+If the query involves no group-by clause, as shown in Algorithm 4, the procedure Compute_aggregate of Algorithm 3 is called to scan the whole m-table m_Chase(T)m\_Chase(T).
+When running this call, line 1 in Algorithm 4, the main loop lines 6-32 of Algorithm 3 scans m_Chase(T)m\_Chase(T) and computes values appearing in the answers.
+The following proposition shows that Algorithm 4 is correct, except when $aggr=sum$.
+In other words, if $aggr\neq sum$ Algorithm 4 returns ${\sf C\_ans}(\mathcal{AQ})$.
+Moreover, it is also shown that ${\sf C\_ans}^{*}(\mathcal{AQ})$ is correctly computed by Algorithm 4 in any case.
 Algorithm 4 Consistent answer to analytic queries with no group-by clause
-0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $aggr(M_{i})$ from $T$ where $\Gamma$
+0:  The m-table m_Chase(T)m\_Chase(T) and $\mathcal{AQ}:$~select $aggr(M_{i})$ from $T$ where $\Gamma$
 0:
-$[min\_ans,max\_ans]$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
+[min_ans,max_ans][min\_ans,max\_ans] // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
 $ans^{*}$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$
-1:  Call procedure Compute_aggregate with input parameters $m\_Chase(T)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
+1:  Call procedure Compute_aggregate with input parameters m_Chase(T)m\_Chase(T), $\mathcal{AQ}$ andwith output parameters change_min_maxchange\_min\_max, min_ansmin\_ans, max_ansmax\_ans, $ans^{*}$
 2:  if $change\_min\_max={\tt true}$ then
-3:      return  $([min\_ans,max\_ans]$, $ans^{*}$)
+3:    return ([min_ans,max_ans]([min\_ans,max\_ans], $ans^{*}$)
 4:  else
@@
 7:    if $aggr=count$ then
-8:           return  ($[0,0]$, 00)
-In the following example we show that Algorithm 3 may fail to return exact values for $glb$ and $lub$ when the aggregate operator is $sum$, operating on positive values and negative values.
+8:     return ($[0,0]$, 0)
+Proposition 11
+Let $T$ be a star-table over universe $U$ and let $\mathcal{AQ}:$~select $aggr(M_{i})$ from $T$ where $\Gamma$ be an analytic query with no group-by clause. Then we have:
+  • If aggraggr is minmin, maxmax or countcount, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where min_ansmin\_ans and max_ansmax\_ans are returned by Algorithm 4.
+  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then min_ansmin\_ans and max_ansmax\_ans as returned by Algorithm 4 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
+    Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$.
+  • For every aggregate operator and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
+Proof. See Appendix G. $\Box$
+In the following example we show that Algorithm 4 may fail to return exact values for glbglb and lublub when the aggregate operator is sumsum, operating on positive values and negative values.
 Example 9
@@
 $k_{1}$ $10$
-$k^{\prime}_{1}$ $-15$
-$k^{\prime}_{1}$ $20$
+$k^{\prime}_{1}$ 15-15
+$k^{\prime}_{1}$ 2020
 $D_{2}$ $K_{2}$ $A_{2}$
-$k_{2}$ $2$
-$k^{\prime}_{2}$ 00
-$k^{\prime}_{2}$ $3$
+$k_{2}$ 22
+$k^{\prime}_{2}$ 0
+$k^{\prime}_{2}$ 33
 $F$ $K_{1}$ $K_{2}$ $M_{1}$
 $k_{1}$ $k_{2}$ $30$
-$k^{\prime}_{1}$ $k^{\prime}_{2}$ $-10$
+$k^{\prime}_{1}$ $k^{\prime}_{2}$ 10-10
 $k_{1}$ $k^{\prime}_{2}$ $100$
-Thus $m\_Chase(T)$ contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
-[…]
-and $T$ has four repairs denoted by $R_{i}$ for $i=1,\ldots,4$, whose sets $R^{\mathbf{K}}_{i}$ of tuples over $U$ are defined by:
-$R^{\mathbf{K}}_{1}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
-$R^{\mathbf{K}}_{2}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$
-$R^{\mathbf{K}}_{3}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
-$R^{\mathbf{K}}_{4}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$.
+Thus m_Chase(T)m\_Chase(T) contains three m-tuples $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ defined over $U$ as follows:
+$\sigma_{1}=(k_{1})(k_{2})(10)(2)(30)$, $\sigma_{2}=(k^{\prime}_{1})(k^{\prime}_{2})(-15\,20)(0\,3)(-10)$, $\sigma_{3}=(k_{1})(k^{\prime}_{2})(10)(0\,3)(100)$
+and $T$ has four repairs denoted by $R_{i}$ for $i=1,\ldots,4$, whose sets $R^{\mathbb{K}}_{i}$ of tuples over $U$ are defined by:
+$R^{\mathbb{K}}_{1}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
+$R^{\mathbb{K}}_{2}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},-15,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$
+$R^{\mathbb{K}}_{3}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,0,-10),(k_{1},k^{\prime}_{2},10,0,100)\}$
+$R^{\mathbb{K}}_{4}=\{(k_{1},k_{2},10,2,30),(k^{\prime}_{1},k^{\prime}_{2},20,3,-10),(k_{1},k^{\prime}_{2},10,3,100)\}$.
 For $\Gamma=(A_{1}>0)\wedge(A_{2}>0)$ and $\mathcal{AQ}:$ select $aggr(M_{1})$ from $T$ where $\Gamma$, we have $\Sigma(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ and $\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1}\}$, because ${\sf tuples}(\sigma_{1}(A_{1}A_{2}))\subseteq Sat(\Gamma)$ holds, ${\sf tuples}(\sigma_{2}(A_{1}A_{2}))\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{2}(A_{1}A_{2}))\subseteq Sat(\Gamma)$, and ${\sf tuples}(\sigma_{3}(A_{1}A_{2}))\cap Sat(\Gamma)\neq\emptyset$ holds but not ${\sf tuples}(\sigma_{3}(A_{1}A_{2}))\subseteq Sat(\Gamma)$.
@@
   • If $aggr=min$ then, for every $i=1,\ldots,3$, ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=30$ and ${\sf Ans}(\mathcal{AQ}^{[R_{4}]})=-10$. Thus, ${\sf C\_ans}(\mathcal{AQ})=[-10,30]$.
-    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\min(\{30,\min\{-10,30\}\})=-10$ when processing $\sigma_{2}$, and then to $\min(\{-10,\min\{100,-10\}\})=-10$ when processing $\sigma_{3}$.
-Similarly, $max\_ans$ is first set to $30$ when processing $\sigma_{1}$, then to $\max(\{30,\min\{-10,30\}\})=30$ when processing $\sigma_{2}$, and then to $\max(\{30,\min\{100,30\}\})=30$ when processing $\sigma_{3}$.
-Hence, Algorithm 3 returns $[-10,30]$ as expected.
+    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is first set to $30$ when processing $\sigma_{1}$, then to $\min(\{30,\min\{-10,30\}\})=-10$ when processing $\sigma_{2}$, and then to $\min(\{-10,\min\{100,-10\}\})=-10$ when processing $\sigma_{3}$.
+Similarly, max_ansmax\_ans is first set to $30$ when processing $\sigma_{1}$, then to $\max(\{30,\min\{-10,30\}\})=30$ when processing $\sigma_{2}$, and then to $\max(\{30,\min\{100,30\}\})=30$ when processing $\sigma_{3}$.
+Hence, Algorithm 4 returns [10,30][-10,30] as expected.
   • Similarly, if $aggr=max$, we have ${\sf C\_ans}(\mathcal{AQ})=[30,100]$, which is also returned by Algorithm 4.
@@
 Thus, ${\sf C\_ans}(\mathcal{AQ})=[1,3]$.
-    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is first set to $1$ when processing $\sigma_{1}$ and then unchanged when processing $\sigma_{2}$ and $\sigma_{3}$.
-Moreover, $max\_ans$ is increased by $1$ for each m-tuple $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$.
+    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is first set to $1$ when processing $\sigma_{1}$ and then unchanged when processing $\sigma_{2}$ and $\sigma_{3}$.
+Moreover, max_ansmax\_ans is increased by $1$ for each m-tuple $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$.
 Thus Algorithm 4 returns $[1,3]$ as expected.
@@
 Thus, ${\sf C\_ans}(\mathcal{AQ})=[30,130]$.
-    When running the call of the procedure Compute_aggregate in Algorithm 3, $min\_ans$ is successively set to $30$, $(30-10)=20$ when processing $\sigma_{1}$ and then $\sigma_{2}$, and unchanged when processing $\sigma_{3}$.
-On the other hand, $max\_ans$ is successively set to $30$ when processing $\sigma_{1}$, unchanged when processing $\sigma_{2}$ and set to $(30+100)=130$ when processing $\sigma_{3}$.
+    When running the call of the procedure Compute_aggregate in Algorithm 4, min_ansmin\_ans is successively set to $30$, (3010)=20(30-10)=20 when processing $\sigma_{1}$ and then $\sigma_{2}$, and unchanged when processing $\sigma_{3}$.
+On the other hand, max_ansmax\_ans is successively set to $30$ when processing $\sigma_{1}$, unchanged when processing $\sigma_{2}$ and set to $(30+100)=130$ when processing $\sigma_{3}$.
 Thus the call of the procedure Compute_aggregate in Algorithm 4 returns $[20,130]$.
   • On the other hand, assuming that the second tuple in $F$ is $(k^{\prime}_{1}k^{\prime}_{2},10)$ (instead of $(k^{\prime}_{1}k^{\prime}_{2},-10)$), then ${\sf C\_ans}(\mathcal{AQ})=[30,140]$, because ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})$ is $30$ if $i=1$ or $i=3$, $130$ if $i=2$ and $140$ if $i=4$.
-It can be seen that when running the call of the procedure Compute_aggregate in Algorithm 3, we obtain $min\_ans=30$ and $max\_ans=140$.
+It can be seen that when running the call of the procedure Compute_aggregate in Algorithm 4, we obtain min_ans=30min\_ans=30 and max_ans=140max\_ans=140.
 $\Box$
 Algorithm 5 Consistent answer to analytic queries with a group-by clause
-0:  The m-table $m\_Chase(T)$ and $\mathcal{AQ}:~{}$select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$
+0:  The m-table m_Chase(T)m\_Chase(T) and $\mathcal{AQ}:$~select $X,aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$
 0:
-$Cons\_Ans$: a set of pairs $\langle x,[min\_ans,max\_ans]\rangle$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
+Cons_AnsCons\_Ans: a set of pairs $\langle x,[min\_ans,max\_ans]\rangle$ // meant to be equal to ${\sf C\_ans}(\mathcal{AQ})$
 $Ans^{*}$: a set of pairs $(x,ans^{*})$ // meant to be equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$
 1:  $Cons\_Ans:=\emptyset$ ; $Ans^{*}:=\emptyset$ ; $Temp:=\emptyset$
-2:  for all $\sigma$ in $m\_Chase(T)$ do
-3:      if $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ then
-4:           if $(Y\to B$ in $FD$ and $YB\subseteq X\Rightarrow|{\sf tuples}(\sigma(B))|=1)$ and $({\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$then
+2:  for all $\sigma$ in m_Chase(T)m\_Chase(T) do
+3:    if $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$ then
+4:     if $(Y\to B$ in $FD$ and $YB\subseteq X\Rightarrow|{\sf tuples}(\sigma(B))|=1)$ and $({\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$) then
 5:       $Temp:=Temp\cup\{\sigma\}$
-6:  for all $X$-value $x$ occurring in $Temp$ do
+6:  for all $X$-value $x$ occurring in TempTemp do
 7:    $Temp(x):=\{\sigma\in Temp~|~x\in{\sf tuples}(\sigma(X))\}$
-8:      Call procedure Compute_aggregate with input parameters $Temp(x)$, $\mathcal{AQ}$ andwith output parameters $change\_min\_max$, $min\_ans$, $max\_ans$, $ans^{*}$
+8:    Call procedure Compute_aggregate with input parameters $Temp(x)$, $\mathcal{AQ}$ andwith output parameters change_min_maxchange\_min\_max, min_ansmin\_ans, max_ansmax\_ans, $ans^{*}$
 9:    if $change\_min\_max={\tt true}$ then
@@
 As in the case of analytic queries with no group-by clause, when the query involves a group-by clause, the computation of the consistent answers also involves a call of the procedure Compute_aggregate, as shown line 8 of Algorithm 5.
-This algorithm works as follows: a first scan of $m\_Chase(T)$, lines 2-5 retrieves in the set $Temp$ all relevant m-tuples $\sigma$, namely, such that $sch(\sigma)$ contains all attributes occurring in $\mathcal{AQ}$ and such that the $X$-values are consistent and may be associated with a tuple over $sch(\Gamma)$ satisfying the condition $\Gamma$.
+This algorithm works as follows: a first scan of m_Chase(T)m\_Chase(T), lines 2-5 retrieves in the set $Temp$ all relevant m-tuples $\sigma$, namely, such that $sch(\sigma)$ contains all attributes occurring in $\mathcal{AQ}$ and such that the $X$-values are consistent and may be associated with a tuple over $sch(\Gamma)$ satisfying the condition $\Gamma$.
 Then, a subsequent loop lines 6-11 operates a call of the procedure Compute_aggregate for each $X$-value $x$ occurring in the m-tuples of the set $Temp$.
@@
 Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}$ an analytic query with a group-by clause.
-  • If $aggr=min$, $aggr=max$ or $aggr=count$, then $Cons\_Ans$ as returned by Algorithm 4 is equal to ${\sf C\_ans}(\mathcal{AQ})$.
-  • If $aggr=sum$, then $\langle x,[min\_ans,max\_ans]\rangle$ is in $Cons\_Ans$ as returned by Algorithm 4 if and only if ${\sf C\_ans}(\mathcal{AQ})$ contains $\langle x,[lub,glb]\rangle$ such that $min\_ans\leq glb$ and $max\_ans\geq lub$.
+  • If $aggr=min$, $aggr=max$ or $aggr=count$, then Cons_AnsCons\_Ans as returned by Algorithm 5 is equal to ${\sf C\_ans}(\mathcal{AQ})$.
+  • If $aggr=sum$, then $\langle x,[min\_ans,max\_ans]\rangle$ is in Cons_AnsCons\_Ans as returned by Algorithm 5 if and only if ${\sf C\_ans}(\mathcal{AQ})$ contains $\langle x,[lub,glb]\rangle$ such that $min\_ans\leq glb$ and $max\_ans\geq lub$.
     Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then $Cons\_Ans={\sf C\_ans}(\mathcal{AQ})$.
-  • For every aggregate function and every selection condition $\Gamma$, $Ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
+  • For every aggregate operator and every selection condition $\Gamma$, $Ans^{*}$ as returned by Algorithm 5 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
 Proof.
-First, the loop lines 2-5 scans $m\_Chase(T)$ and collects in the set $Temp$ the only m-tuples necessary to compute the consistent answer.
-Indeed, the collected m-tuples are all m-tuples $\sigma$ such that: they are defined over a super-set of $\mathbf{K}\cup sch(\sigma)$, their $X$-component contains consistent tuples, and their m-tuples over $sch(\Gamma)$ contain at least one tuple of $Sat(\Gamma)$.
-This is so because any m-tuple in $m\_Chase(T)$ not satisfying the above conditions can not contribute to the consistent answer.
+First, the loop lines 2-5 scans m_Chase(T)m\_Chase(T) and collects in the set $Temp$ the only m-tuples necessary to compute the consistent answer.
+Indeed, the collected m-tuples are all m-tuples $\sigma$ such that: they are defined over a super-set of $\mathbb{K}\cup sch(\sigma)$, their $X$-component contains consistent tuples, and their m-tuples over $sch(\Gamma)$ contain at least one tuple of $Sat(\Gamma)$.
+This is so because any m-tuple in m_Chase(T)m\_Chase(T) not satisfying the above conditions can not contribute to the consistent answer.
 Then, in the next steps for every collected $X$-value, its associated aggregate value is evaluated as done in Algorithm 4.
@@
 $\Box$
-We emphasize that Algorithm 3 and Algorithm 4 show that the consistent answers to a given analytic query $\mathcal{AQ}$, with or without a group-by clause, are computed in linear time in the size of $m\_Chase(T)$.
-Recalling that the computation of $m\_Chase(T)$ has been shown in LauS25 to be polynomial in time with respect to the size of $T$, it turns out that ${\sf C\_ans}(\mathcal{AQ})$ and ${\sf C\_ans}^{*}(\mathcal{AQ})$ can be computed in polynomial time with respect to the size of $T$.
+We emphasize that Algorithm 4 and Algorithm 5 show that the consistent answers to a given analytic query $\mathcal{AQ}$, with or without a group-by clause, are computed in linear time in the size of m_Chase(T)m\_Chase(T).
+Recalling that the computation of m_Chase(T)m\_Chase(T) has been shown earlier to be quasi-linear with respect to the size of $T$ (or in $\mathcal{O}(W\cdot\log(W))$ where $W$ is the size of the data warehouse), it turns out that ${\sf C\_ans}(\mathcal{AQ})$ and ${\sf C\_ans}^{*}(\mathcal{AQ})$ can be computed in quasi-linear time with respect to the size of the data warehouse.
 6.4 Analytic Queries with a group-by-having Clause
@@
   • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=15$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,20],[5,15]\rangle\}$ and so, ${\sf C\_ans}(\mathcal{AQ}_{3})$ is empty.
-Notice incidentally that it is easily seen from Figure 3 that there exist repairs $R$ in ${\sf Rep}(T)$ for which ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})=\{\langle a_{2},20,15\rangle\}$, implying that ${\sf Ans}(\mathcal{AQ}_{3}^{[R]})=\emptyset$.
+Notice incidentally that it is easily seen from Figure 4 that there exist repairs $R$ in ${\sf Rep}(T)$ for which ${\sf Ans}(\mathcal{AQ}_{3}^{\prime[R]})=\{\langle a_{2},20,15\rangle\}$, implying that ${\sf Ans}(\mathcal{AQ}_{3}^{[R]})=\emptyset$.
   • If we assume that $m_{1}=5$, $m^{\prime}_{1}=5$ and $m^{\prime\prime}_{1}=8$, then ${\sf C\_ans}(\mathcal{AQ}^{\prime}_{3})=\{\langle a_{2},[10,13],[5,8]\rangle\}$ and so, $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
-In this case, it is easily seen from Figure 3 that for every repair $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}_{3}^{{}^{\prime}[R]})$ contains a pair $\langle a_{2},\mu,\nu\rangle$ where $10\leq\mu\leq 13$ and $\nu<10$, implying that $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
+In this case, it is easily seen from Figure 4 that for every repair $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}_{3}^{\prime[R]})$ contains a pair $\langle a_{2},\mu,\nu\rangle$ where $10\leq\mu\leq 13$ and $\nu<10$, implying that $\langle a_{2},[10,13]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ}_{3})$.
 In the light of this example, we argue that our approach can deal with analytic queries involving a having clause with condition $\Upsilon$, under the following restrictions:
   1. 1.
-[…]
+    $\Upsilon$ is a conjunctive boolean expression built up from atoms of the form $aggr(M_{i})\,\theta\,\alpha$ where $M_{i}$ is a measure attribute in $\mathbb{M}$, $\theta$ is a comparison predicate in ${<,,>,}\{$$\leq$$,\geq\}$ and $\alpha$ is a number.
   2. 2.
@@
 $\mathcal{AQ}:{\tt select}~X$, $aggr(M_{i})$ from $T$ where $\Gamma$ group by $X$ having $\Upsilon$
-such that the above restrictions are met. If $\Upsilon=\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$, let $\mathcal{AQ}^{\prime}$ be the following analytic query with no having clause:
+such that the above two restrictions regarding $\Upsilon$ are met.
+If $\Upsilon=\Upsilon_{1}\wedge\ldots\wedge\Upsilon_{h}$, let $\mathcal{AQ}^{\prime}$ be the following analytic query with no having clause:
 $\mathcal{AQ}^{\prime}:{\tt select}~X$, $aggr(M_{i})$, $aggr_{1}(M_{i_{1}})$, $\ldots$ , $aggr_{h}(M_{i_{h}})$ from $T$ where $\Gamma$ group by $X$
-Then, $\langle x,[glb,lub]\rangle$ belongs to ${\sf C\_ans}(\mathcal{AQ})$ if and only if there exists $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$.
+The tuple $\langle x,[glb,lub]\rangle$ belongs to ${\sf C\_ans}(\mathcal{AQ})$ if and only if there exists a tuple of the form $\langle x,[glb,lub],$ $glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ holds for every $p=1,\ldots,h$.
 Proof.
-Let $\langle x,[glb,lub],$ $[glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$.
-Given $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{p}\rangle$ such that $\mu\in[glb,lub]$ and $\mu_{p}\in[glb_{p},lub_{p}]$ for $p=1,\ldots,h$ hold.
+Let $\langle x,[glb,lub],$ $glb_{1},lub_{1}],\ldots,[glb_{h},lub_{h}]\rangle$ in ${\sf C\_ans}(\mathcal{AQ}^{\prime})$ such that for every $p=1,\ldots,h$, $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$.
+Given $R$ in ${\sf Rep}(T)$, ${\sf Ans}(\mathcal{AQ}^{\prime[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{p}\rangle$ such that $\mu\in[glb,lub]$ and $\mu_{p}\in[glb_{p},lub_{p}]$ for $p=1,\ldots,h$ hold.
 Hence, for $p=1,\ldots,h$, $\mu_{p}$ is in $Sat(\Upsilon_{p})$, implying that $\langle x,[glb,lub]\rangle$ is in ${\sf C\_ans}(\mathcal{AQ})$.
-Conversely, let $\langle x,[glb,lub]\rangle$ be in ${\sf C\_ans}(\mathcal{AQ})$ and let $p$ such that $[glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ does not hold.
+Conversely, let $\langle x,[glb,lub]\rangle$ be in ${\sf C\_ans}(\mathcal{AQ})$ and let $p$ such that $glb_{p},lub_{p}]\subseteq Sat(\Upsilon_{p})$ does not hold.
 Since $Sat(\Upsilon_{p})$ is an interval, this implies that $glb_{p}\not\in Sat(\Upsilon_{p})$ or that $lub_{p}\not\in Sat(\Upsilon_{p})$.
-Since there exists a repair $R_{min}$, respectively a repair $R_{max}$, such that $aggr_{p}(M_{i_{p}})=glb_{p}$, respectively $aggr_{p}(M_{i_{p}})=lub_{p}$, it turns out that there exists at least one repair $R$ such that $ans(\mathcal{AQ}^{{}^{\prime}[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{h}\rangle$ such that $\mu_{p}\not\in Sat(\Upsilon_{p})$.
+Since there exists a repair $R_{min}$, respectively a repair $R_{max}$, such that $aggr_{p}(M_{i_{p}})=glb_{p}$, respectively $aggr_{p}(M_{i_{p}})=lub_{p}$, it turns out that there exists at least one repair $R$ such that $ans(\mathcal{AQ}^{\prime[R]})$ contains a tuple $\langle x,\mu,\mu_{1},\ldots,\mu_{h}\rangle$ such that $\mu_{p}\not\in Sat(\Upsilon_{p})$.
 Thus $\langle x,\mu\rangle$ is not in $ans(\mathcal{AQ}^{[R]})$, which implies that $\langle x,[glb,lub]\rangle$ is not in ${\sf C\_ans}(\mathcal{AQ})$; a contradiction which completes the proof.
 $\Box$
-It should be clear from Proposition 11 and Proposition 12 that, under the restrictions earlier stated, the consistent answer of an analytic query involving a group-by-having clause can be computed in polynomial time with respect to the size of $T$.
+It should be clear from Proposition 12 and Proposition 13 that, under the restrictions regarding $\Upsilon$ earlier stated, the consistent answer of an analytic query involving a group-by-having clause can be computed in quasi-linear time with respect to the size of the data warehouse.
 To end this section, we illustrate why in our approach, the sets $Sat(\Upsilon_{p})$ ($p=1,\ldots,h$) must be intervals.
@@
 6.5 Handling distinct Clauses
-In this section, we show how to handle the distinct clause associated with the operator $count$.
-Recalling that $count({\tt distinct}\,M_{i})$ counts each $M_{i}$-value only once, independently from the number of its occurrences, we consider analytic queries $\mathcal{AQ}$ with no group-by clause, knowing that the presence of a group-by clause does not raise particular difficulties.
-Given such a query $\mathcal{AQ}:$ select $count({\tt distinct}~{}M_{i})$ from $T$ where $\Gamma$, we let ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$.
-Similarly to the general case of the aggregate $sum$, an effective way of computing ${\sf C\_ans}(\mathcal{AQ})$ is currently unknown to the authors.
-This is so because, according to Algorithm 2, in order to determine $lub$ (respectively $glb$) without counting the same value more than once, every choice of one $M_{i}$-value in each set ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$ (respectively in $\Sigma^{+}(\mathcal{AQ})$) must be analyzed separately.
-As such a processing clearly requires to compute all possible choices, it is not polynomial, and so not tractable.
-Consequently, instead of computing the exact values of $lub$ and $glb$, we propose a tractable approximation of ${\sf C\_ans}(\mathcal{AQ})$.
-To this end, we first notice that the values to be counted for determining $lub$ are obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma(\mathcal{AQ})$.
-Therefore, the set resulting from such a choice can not contain $(a)$ more distinct values than there are m-tuples in $\Sigma(\mathcal{AQ})$ and $(b)$ more distinct values than there are distinct values in the union of all sets ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$.
+The clause distinct is used in SQL queries to avoid the processing of duplicates.
+The most frequent cases in which this key word is used are (1) associated with a select clause, in which case repeated values are not displayed to the user, and (2) associated with aggregation operator countcount, in which case every value is counted only once.
+In our approach, it turns out that the first case does not raise specific difficulties, since as usual, this case simply necessitates to check whether a given value has been displayed earlier or not.
+However, case (2) is much more involved because in order to compute the interval in the consistent answer to a clause of the form select countcount(distinct $M_{i}$)..., it must be known for each repair which $M_{i}$-values have been counted, and this is clearly not tractable.
+In order to show how to handle the distinct clause associated with the operator countcount, we focus on analytic queries of the form $\mathcal{AQ}:$ select $count({\tt distinct}~M_{i})$ from $T$ where $\Gamma$, because the presence of a group-by clause does not raise any particular difficulties.
+Similarly to the general case of the aggregate sumsum, an effective way of computing the interval $[glb,lub]$ of ${\sf C\_ans}(\mathcal{AQ})$ is currently unknown to the authors.
+This is so because, as mentioned above, repairs must be considered separately in order to compute lublub and glbglb, which cannot be done when running Algorithm 3.
+Consequently, instead of computing the exact values of lublub and glbglb, we propose a tractable approximation of ${\sf C\_ans}(\mathcal{AQ})$, easily implemented based on Algorithm 3 and Algorithm 4.
+To this end, we first notice that the values to be counted for determining lublub are obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma(\mathcal{AQ})$.
+Therefore, the set resulting from such a choice can not contain $(a)$ more distinct values than the number of m-tuples in $\Sigma(\mathcal{AQ})$ and $(b)$ more distinct values than the number of distinct values in the union of all sets ${\sf tuples}(\sigma(M_{i}))$ where $\sigma$ is in $\Sigma(\mathcal{AQ})$.
 This can be formally stated as $(a)$ $lub\leq|\Sigma(\mathcal{AQ})|$ and $(b)$ $lub\leq|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$, which we write as
 $lub\leq\inf\left(|\Sigma(\mathcal{AQ})|,|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|\right)$.
-If $count\_max$ denotes the infimum defined above, we emphasize that Algorithm 2 can be modified so as to compute $count\_max$ in polynomial time.
-Indeed, it is easy to see that $|\Sigma(\mathcal{AQ})|$ is equal to the value of $max\_ans$ returned by Algorithm 3 in the case of $count$.
-However, computing the second cardinality requires extra storage and extra computation to ensure that every value in the union is counted only once.
-Regarding the approximation of $glb$, we recall from Algorithm 2 that the values to be counted for determining $glb$ are distinct values obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$.
-Thus, if $\Sigma^{+}(\mathcal{AQ})=\emptyset$, $glb=0$, and otherwise, a trivial lower bound for $glb$ is $1$, because all sets ${\sf tuples}(\sigma(M_{i}))$ for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, are nonempty.
-To find a more accurate value of this lower bound, we notice that any choice of one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, concerns at least as many distinct $M_{i}$-values as there are in the union of all ${\sf tuples}(\sigma(M_{i}))$ of cardinality $1$, for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$.
+If count_maxcount\_max denotes the infimum defined above, we emphasize that Algorithm 3 can be modified so as to compute count_maxcount\_max.
+Indeed, it is easy to see that $|\Sigma(\mathcal{AQ})|$ is equal to the value of proc_maxproc\_max returned by Algorithm 3 in the case of countcount.
+Moreover, the computation of $|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$ can be implemented in Algorithm 3 in a similar way, at the cost of an extra storage to ensure that every value in the union is counted only once.
+Then, once the two cardinalities are known, count_maxcount\_max can easily be computed and returned by modifying Algorithm 4 accordingly.
+Regarding the approximation of glbglb, we recall from Algorithm 3 that the values to be counted for determining glbglb are distinct values obtained by picking exactly one $M_{i}$-value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$.
+Thus, if $\Sigma^{+}(\mathcal{AQ})=\emptyset$, $glb=0$, and otherwise, a trivial lower bound for glbglb is $1$, because all sets ${\sf tuples}(\sigma(M_{i}))$ for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, are nonempty.
+To find a more accurate lower bound, we notice that any choice of one value in ${\sf tuples}(\sigma(M_{i}))$ for every $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$, concerns at least as many distinct $M_{i}$-values as there are in the union of all ${\sf tuples}(\sigma(M_{i}))$ of cardinality 11, for $\sigma$ in $\Sigma^{+}(\mathcal{AQ})$.
 Denoting by $\Sigma^{+}_{1}(\mathcal{AQ})$ the set of these m-tuples, that is $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma\in\Sigma^{+}(\mathcal{AQ})~|~|{\sf tuples}(\sigma(M_{i}))|=1\}$, whenever $\Sigma^{+}(\mathcal{AQ})\neq\emptyset$, we have:
 $glb\geq|\bigcup_{\sigma\in\Sigma^{+}_{1}(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$.
-Denoting by $count\_min$ the cardinality shown above, we notice that, as for $count\_max$, Algorithm 3 can be modified so as to compute $count\_min$ in polynomial time.
-Another important remark regarding distinct is that if for every $\sigma$ in $m\_Chase(T)$, we have $|{\sf tuples}(\sigma(M_{i}))|=1$ (i.e., the dependency $\mathbf{K}\to M_{i}$ is satisfied), then $\Sigma^{+}_{1}(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$, and:
+Denoting by count_mincount\_min the cardinality shown above, we notice that, as for count_maxcount\_max, Algorithm 3 and Algorithm 4 can be modified so as to compute count_mincount\_min.
+Another important remark regarding distinct is that if for every $\sigma$ in m_Chase(T)m\_Chase(T), we have $|{\sf tuples}(\sigma(M_{i}))|=1$ (i.e., the dependency $\mathbb{K}\to M_{i}$ is satisfied), then $\Sigma^{+}_{1}(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$, and:
 $\bullet$ $count\_max=lub=|\bigcup_{\sigma\in\Sigma(\mathcal{AQ})}{\sf tuples}(\sigma(M_{i}))|$     $(1)$
@@
 be an analytic query involving a distinct clause, but no selection condition.
-In this case, we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$ and it can be seen from Figure 2 that the only m-tuples $\sigma$ in $m\_Chase(T)$ such that $\mathbf{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ are the first three m-tuples in $m\_Chase(T)$ that we denote by $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$.
-Thus, $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$ are equal to $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ that are such that ${\sf tuples}(\sigma_{1}(M_{1}))={\sf tuples}(\sigma_{3}(M_{1}))=\{m_{1}\}$ and ${\sf tuples}(\sigma_{2}(M_{1}))=\{m^{\prime}_{1},m^{\prime\prime}_{1}\}$.
-Referring to the four repairs of $m\_Chase(T)$ shown in Figure 3, we have ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=2$ for $i=1,\ldots,4$, implying that ${\sf C\_ans}(\mathcal{AQ})=2$.
-On the other hand, applying the approximations shown above, we have $|\Sigma(\mathcal{AQ})|=|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=3$ and thus, $count\_max=3$.
+In this case, we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})$ and it can be seen from Figure 2 that the only m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\cup sch(\mathcal{AQ})\subseteq sch(\sigma)$ are the first three m-tuples in m_Chase(T)m\_Chase(T) that we denote by $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$.
+Thus we have $\Sigma(\mathcal{AQ})=\Sigma^{+}(\mathcal{AQ})=\{\sigma_{1},\sigma_{2},\sigma_{3}\}$, along with ${\sf tuples}(\sigma_{1}(M_{1}))={\sf tuples}(\sigma_{3}(M_{1}))=\{m_{1}\}$ and ${\sf tuples}(\sigma_{2}(M_{1}))=\{m^{\prime}_{1},m^{\prime\prime}_{1}\}$.
+Referring to the four repairs of m_Chase(T)m\_Chase(T) shown in Figure 4, we have ${\sf Ans}(\mathcal{AQ}^{[R_{i}]})=2$ for $i=1,\ldots,4$, implying that ${\sf C\_ans}(\mathcal{AQ})=2$.
+On the other hand, applying the approximations shown above, we have $|\Sigma(\mathcal{AQ})|=|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=3$ and thus, count_max=3count\_max=3.
 Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$.
-Hence, we obtain $count\_min=1$, producing the interval $[1,3]$, which approximates, but is not equal to ${\sf C\_ans}(\mathcal{AQ})$.
+Hence, we obtain count_min=1count\_min=1, producing the interval $[1,3]$, which approximates, but is not equal to ${\sf C\_ans}(\mathcal{AQ})$.
 Changing $m^{\prime}_{1}$ to $m_{1}$ in $\sigma_{2}$ yields that ${\sf Ans}(\mathcal{AQ}^{[R_{1}]})={\sf Ans}(\mathcal{AQ}^{[R_{2}]})=1$ and ${\sf Ans}(\mathcal{AQ}^{[R_{3}]})={\sf Ans}(\mathcal{AQ}^{[R_{4}]})=2$.
 Thus ${\sf C\_ans}(\mathcal{AQ})=[1,2]$.
-On the other hand, applying the approximations as above, we have $|\Sigma(\mathcal{AQ})|=3$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=2$ and so, $count\_max=2$.
+On the other hand, applying the approximations as above, we have $|\Sigma(\mathcal{AQ})|=3$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{2}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=2$ and so, count_max=2count\_max=2.
 Moreover, $\Sigma^{+}_{1}(\mathcal{AQ})=\{\sigma_{1},\sigma_{3}\}$ and $|{\sf tuples}(\sigma_{1}(M_{1}))\cup{\sf tuples}(\sigma_{3}(M_{1}))|=1$.
-Hence, we obtain again $count\_min=1$, producing the interval $[1,2]$, which is now equal to ${\sf C\_ans}(\mathcal{AQ})$.
+Hence, we obtain again count_min=1count\_min=1, producing the interval $[1,2]$, which is now equal to ${\sf C\_ans}(\mathcal{AQ})$.
 $\Box$
@@
 These efforts were first focused on conjunctive and self-join free queries involving no aggregate operators.
-The case of queries with aggregate operators and possibly with group-by clauses has been introduced in ArenasBCHRS03 and then further investigated in FuxmanFM05 and in KhalfiouiW23 ; KhalfiouiW24 .
+The case of queries with aggregate operators and possibly with group-by clauses has been introduced in ArenasBCHRS03 and then further investigated in FuxmanFM05 .
 Focusing on relations whose constraints are unique key-constraints, these approaches rely on earlier work (see Wijsen19 ) characterizing classes of queries that can be rewritten in polynomial time, into SQL queries whose evaluation yields the consistent answer as defined in ArenasBCHRS03 .
-In parallel, the issue of consistent conjunctive query answering to queries involving aggregate operators has been addressed in DixitK19 , based on a radically different technique, mainly that of reducing the problem to the well-known SAT problem.
-Here again, the constraints of interest are key-constraints associated with relations, as in FuxmanFM05 .
-Comparing our approach with the work in DixitK19 or in FuxmanFM05 , the main differences are the following:
-  1. 1.
-    The functional dependencies in DixitK19 or in FuxmanFM05 do not exactly match those usually considered in star schemas.
-Indeed, the dependencies $\mathbf{K}\to M_{i}$ are not in the scope of these approaches, contrary to the present approach.
-  2. 2.
-    In our approach, we allow missing values in the input tables, which is not the case in DixitK19 or in FuxmanFM05 .
-  3. 3.
-    Whereas in these approaches conjunctive queries are considered, in our approach, we allow specific disjunctions that we call independent selection conditions.
-We recall that independent selection conditions generalize conjunctive selection conditions.
-  4. 4.
-    Dealing with aggregate operators $sum$ and $count({\tt distinct}\,.)$ in our approach may lead to approximate query answers, which is not the case in DixitK19 and in FuxmanFM05 .
-Therefore, it turns out that the two approaches are hardly comparable.
-However, we argue in this section that all results stated in the previous sections also hold when the functional dependencies $\mathbf{K}\to M_{i}$ are not considered.
-Moreover, in this case, when the aggregate is $count({\tt distinct}\,.)$ on a measure attribute, the consistent answer can be effectively evaluated, instead of being approximated as earlier explained.
-More precisely, we show that our approach works in the context of DixitK19 or of FuxmanFM05 , assuming that the key of the fact table $F$ consists of all its attributes (namely all key attributes along with all measure attributes).
-Indeed, if $FD$ contains no functional dependency of the form $\mathbf{K}\to M_{i}$ then for every $\sigma$ in $m\_Chase(T)$ and every $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is a singleton.
+In parallel, the issue of consistent conjunctive query answering to queries involving or not aggregate operators has been addressed in DixitK19 ; DixitK22 , based on a radically different technique, mainly that of reducing the problem to the well-known SAT problem.
+By relying on $Chase$ semantics, our approach does not strictly follow the same semantics as those cited above.
+Indeed, as argued in JIIS ; LauS25 and illustrated in Example 2, in our approach, database consistency is not checked against the database itself but rather, against its ‘chased’ version.
+It follows that, as explained in Section 4.1, the notion of repair in our approach does not coincide with the usual one in the literature, where a repair is a maximal and consistent subset of the tuples in the database.
+It follows that, although our definitions of consistent answer to a query, whether standard or analytic, follow the usual definitions (i.e., equal to the intersection of the answers to the query in every repair for standard queries, and equal to the interval containing all aggregate values found in the repairs for analytic queries), our semantics to consistent query answering does not coincide with those in traditional approaches.
+Nevertheless, since in the present work we concentrate on databases operating over a star schema, our approach is a particular case of databases whose relations have a unique primary key as in DixitK19 ; DixitK22 ; FuxmanFM05 ; KhalfiouiW23 ; KhalfiouiW24 .
+Moreover, it has been argued (see Proposition 4 in Section 4.2) that, in this case, our notion of repair can be seen as the usual notion of repair applied to the set of all true tuples (which is super-set of the tuples actually stored in the database).
+We also notice that in DixitK19 ; DixitK22 , and contrary to our approach, constraints other than key dependencies are considered since these approaches also consider denial constraints, which are known to generalize functional dependencies.
+[…]
+Another important specificity of queries in our approach, whether standard queries or analytic, is that their expression involves a single table (in fact the m-table m_Chase(T)m\_Chase(T)), which can be seen as an extension of the lossless join of all tables in the data warehouse.
+As a consequence, using the terminology in FuxmanFM05 , standard queries are tree-queries, whose join graph is the tree having the fact table $F$ as its root and all dimension tables $D_{i}$ as its leaves.
+Hence, as stated by Theorem 1 in FuxmanFM05 , all standard queries in our approach, are SQL-rewritable, meaning that each query $Q$ can be rewritten into an SQL query $Q^{\prime}$ whose answer is the consistent answer to $Q$.
+The case of analytic queries is however less clear.
+Indeed, although a similar result has been stated in FuxmanFM05 , it is argued in KhalfiouiW24 that the SQL rewriting for analytic queries involving the aggregate sumsum is flawed, because this aggregate operator is not monotonic.
+In fact, a consequence of the separation Theorem 1.1 in KhalfiouiW24 is that for any analytic query in our approach whose aggregate operator is monotonic and associative, it is possible to generate two SQL queries, each of them computing a bound of the consistent answer.
+The proof of this flaw, that can be found in DixitK22 or in KhalfiouiW24 , relies on a database schema expressed in our approach as follows: the database has three relations $D_{1}$, $D_{2}$ and $F$, respectively, defined over attributes sets $K_{1}A_{1}$, $K_{2}A_{2}$ and $K_{1}K_{2}M_{1}$, over which the two functional dependencies $K_{1}\to A_{1}$ and $K_{2}\to A_{2}$ are assumed.
+That is, $K_{1}$, respectively $K_{2}$, is the key of $D_{1}$, respectively $D_{2}$, and the key of $F$ is $K_{1}K_{2}M_{1}$ (i.e., $F$ has no non-key attributes).
+Hence, this schema can be seen as a star schema, whose fact table has no embedded functional dependency.
+In this context, it is shown in DixitK22 and in KhalfiouiW24 that computing the consistent answer to the analytic query $\mathcal{AQ}:$ select $sum(M_{1})$ from $T$ where $A_{1}=a_{1}$ and $A_{2}=a_{2}$ (where $a_{1}$ and $a_{2}$ are constants respectively in $dom(A_{1})$ and $dom(A_{2})$) is NP-hard.
+Although the above case does not follow the context of our work, since the functional dependencies $\mathbb{K}\to M_{i}$ are missing, we argue that all results stated earlier also hold in this case.
+Indeed, if $FD$ contains no functional dependency of the form $\mathbb{K}\to M_{i}$ then for every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is a singleton.
 Immediate consequences are that, in this context:
   1. 1.
-    $FD$ is normalized and acyclic, and so, the results stated in LauS25 hold.
+    FDFD is normalized and acyclic, and so, the results stated in LauS25 hold.
   2. 2.
-    Proposition 2(2) is not true: m-tuples in $m\_Chase(T)$ may be distinct and have the same $\mathbf{K}$-value.
+    Proposition 2(2) is not true: m-tuples in m_Chase(T)m\_Chase(T) may be distinct and have the same $\mathbb{K}$-value.
   3. 3.
-    Step P2 in process (P).
-is irrelevant and so, should be ignored.
-Moreover, the second item of Step P3 should be removed because for every $\sigma$ such that $\mathbf{K}M_{i}\subseteq sch(\sigma)$ for $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is reduced to one tuple, and this tuple is consistent.
-However, it should be emphasized that Proposition 6 still holds in this setting because adapting the proof amounts to discard all arguments related to inconsistencies of tuples involving $\mathbf{K}$-values associated with measure values.
+    Step P2 in process (P) is irrelevant and so, should be ignored.
+Moreover, the second item of Step P3 should be removed because for every $\sigma$ such that $\mathbb{K}M_{i}\subseteq sch(\sigma)$ for $i=1,\ldots,p$, ${\sf tuples}(\sigma(M_{i}))$ is reduced to one tuple, and this tuple is consistent.
+However, it should be emphasized that Proposition 7 still holds in this setting because adapting the proof amounts to discard all arguments related to inconsistencies of tuples involving $\mathbb{K}$-values associated with measure values.
   4. 4.
@@
   5. 5.
-    Despite the first paragraph in the proof of Proposition 9 (mentioning that this proof relies on the fact that $m\_Chase(T)$ cannot contain distinct m-tuples with the same $\mathbf{K}$-value), the proposition still holds.
+    Despite the first paragraph in the proof of Proposition 10 (mentioning that this proof relies on the fact that m_Chase(T)m\_Chase(T) cannot contain distinct m-tuples with the same $\mathbb{K}$-value), the proposition still holds.
 Indeed, as Proposition 7 still applies, for any of the two items in Proposition 10, the modified process (P) generates a repair which satisfies the conditions in the proposition.
@@
     Dealing with aggregate operator $count({\tt distinct}\,.)$ in this context yields the exact consistent answer, instead of an approximation.
-This is so because in this case all $\sigma(M_{i})$ contain one value, in which case the equalities (1) and (2) in the previous sub-section hold.
-An important consequence is that Algorithm 2, and thus Algorithm 3 and Algorithm 4, work in this context, showing that consistent answers to analytic queries can be computed in polynomial time, even when no functional dependency of the form $\mathbf{K}\to M_{i}$ is considered.
-It can thus be stated that our approach extends approaches from the literature because we allow missing values in the input tables and we can handle some disjunctive selection conditions (namely independent selection conditions involving no key attributes), at the cost of approximate consistent answers when the aggregate operator is $sum$, operating on positive or negative numbers.
+This is so because in this case all $\sigma(M_{i})$ contain one value, implying that the equalities (1) and (2) in the previous sub-section hold.
+An important consequence is that Algorithm 3, and thus Algorithm 4 and Algorithm 5, work in this context, showing that consistent answers to analytic queries can be computed in quasi-linear time, even when no functional dependency of the form $\mathbb{K}\to M_{i}$ is considered.
+As a consequence, based on the results in DixitK22 or in KhalfiouiW24 , determining whether analytic queries in a star schema according to our approach is SQL-rewritable, is an open issue to be investigated.
+On the other hand, this also shows that our approach extends approaches from the literature, in the sense that we allow missing values in the input tables and we can handle some disjunctive selection conditions (namely independent selection conditions involving no key attributes), at the cost of approximate consistent answers when the aggregate operator is sumsum, operating on positive or negative numbers, or when the aggregate operator is of the form countcount(distinct .).
+To sum up the above remarks, we recall that our approach works according to different semantics than in existing works dealing with consistent answering.
+Although this prevents from accurate comparisons, we emphasize the following two main important features of our work:
+  1. 1.
+    Contrary to any other approach, ours supports the presence of missing values in the data warehouse, at the cost of approximate consistent answers when the aggregate operator is sumsum, operating on positive or negative numbers or when the aggregate operator is of the form countcount(distinct .).
+  2. 2.
+    Our complexity results outperform those from the literature, since we have shown that consistent answers are computed in quasi-linear time, i.e., in $\mathcal{O}(W\cdot\log(W))$, where $W$ is the size of the data warehouse.
 8 Concluding Remarks
-Data analytics requires the collection and integration of data coming from different sources and stored in what is called a data warehouse that is a database operating under a specific type of schema called a star schema.
+Data analytics requires the collection and integration of data coming from different sources and stored in a database operating under a specific type of schema called a ‘star schema’.
+Such a database is referred to in the literature as a data warehouse inmon96 .
 As a consequence, more often than not, such integrated data contain inconsistencies and missing values.
 In this paper, we have seen an approach to efficiently compute consistent answers to analytic queries in data warehouses.
-Extending earlier work concerning the consistent query answering of standard, non-analytic queries in multi-table databases, we presented in this paper an approach to computing exact consistent answers of analytic queries over star schemas under two assumptions: (a) the selection condition in the query involves no keys and (b) the selection condition is independent.
-Our main contributions are: (a) a polynomial algorithm for computing the exact consistent answer to a usual projection-selection-join query over a star schema, and (b) showing that the exact consistent answer to an analytic query over a star schema can be computed also in time polynomial in the size of the data warehouse.
+Extending earlier work concerning the consistent query answering of standard, non-analytic queries in multi-table databases, we presented in this paper an approach to computing exact consistent answers of analytic queries over star schemas under the assumption that the selection condition is independent.
+Our main contribution is to propose a framework for star schemas in which missing values are allowed, and in which computing the consistent answer to a usual projection-selection-join query, or to an analytic query, is quasi-linear in the size of the data warehouse, i.e.
+in $\mathcal{O}(W\cdot\log(W))$ where $W$ is the size of the data warehouse.
 Our current work follows two main lines of research.
 First, the implementation of our approach is an important issue under consideration.
-We emphasize in this respect that since the functional dependencies in a star schema have a simple form, the generic algorithm $m\_Chase$ as presented here can be optimized.
-Moreover, as $m\_Chase(T)$ has to be ‘re-computed’ after each update, the question of incrementally implementing the changes in $m\_Chase(T)$ is crucial regarding efficiency.
-We are currently investigating these issues.
+We notice in this respect that, as m_Chase(T)m\_Chase(T) has to be ‘re-computed’ after each update, the question of incrementally implementing the changes in m_Chase(T)m\_Chase(T) is crucial regarding efficiency.
+This issue is currently investigated based on Algorithm 2 and on the statements in Proposition 2.
 A second line of research is investigating the introduction of hierarchies over dimensions.
@@
 No funds, grants, or other support was received for conducting this study.
+Acknowledments
+The authors thank the reviewers for their thorough reading of our manuscript. Their comments and suggestions helped us improve the contents and the presentation of the modified version.
 Appendix A Proof of Proposition 2
+Proposition 2  Let $T$ be a star-table over universe $U$. The following hold:
+  1. 1.
+    For every $\sigma$ in m_Chase(T)m\_Chase(T) and every $i=1,\ldots,n$, if $K_{i}\in sch(\sigma)$ then $|{\sf tuples}(\sigma(K_{i}))|=1$.
+Consequently, if $\mathbb{K}\subseteq sch(\sigma)$ then $|{\sf tuples}(\sigma(\mathbb{K}))|=1$.
+  2. 2.
+    For every tuple $k$ over $\mathbb{K}$ in $\mathcal{T}$, there exists at most one $\sigma$ in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$.
+  3. 3.
+    Moreover, m_Chase(T)m\_Chase(T) contains the following two kinds of m-tuples:
+    1. (a)
+      $\sigma$ for which there exists $i_{0}\in\{1,\ldots,n\}$ such that:
+      • $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and for every $t\in F$, $t.K_{i_{0}}\neq k_{i_{0}}$,
+      • for every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)=\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
+    2. (b)
+      $\sigma$ such that $\mathbb{K}\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(\mathbb{K}))=(k)$, and
+      • for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})=\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$,
+      • for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$,
+        $\sigma(A)=\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
+Proof.
+1.
+The result comes from the fact that, to generate multi-valued components over an attribute $A$, the algorithm m_Chasem\_Chase considers a functional dependency whose right hand-side is $A$.
+As in the case of a star-table, no dependency in $FD$ has its right hand-side in $\mathbb{K}$, it impossible to generate conflicts on key attributes.
+The proof of this item is therefore complete.
+2.
+Let $\sigma_{1}$ and $\sigma_{2}$ be in m_Chase(T)m\_Chase(T) such that $\mathbb{K}\subseteq sch(\sigma_{i})$ for $i=1,2$ and $\sigma_{1}(\mathbb{K})=\sigma_{2}(\mathbb{K})$.
+Since $\sigma_{1}\neq\sigma_{2}$ (the algorithm eliminates duplicates) there exists $A$ not in $\mathbb{K}$ such that $\sigma_{1}(A)\neq\sigma_{2}(A)$.
+Since $A$ is not in $\mathbb{K}$, either there is $i_{0}$ in $\{1,\ldots,n\}$ such that $A\in sch^{*}(D_{i_{0}})$ or $A$ is in $\mathbb{M}$.
+Hence, $FD$ contains a dependency $X\to A$ where $X=K_{i_{0}}$ in the former case, or $X=\mathbb{K}$ in the latter case.
+Since, as stated above, $|\sigma_{1}(X)|=|\sigma_{2}(X)|=1$, we have $\sigma_{1}(X)=\sigma_{2}(X)$.
+Thus, applying the m_Chase rule to m_Chase(T)m\_Chase(T) changes $\sigma_{1}(A)$ and $\sigma_{2}(A)$ into $\sigma_{1}(A)\cup\sigma_{2}(A)$.
+As by definition of m_Chase(T)m\_Chase(T), m_Chase(m_Chase(T))=m_Chase(T)m\_Chase(m\_Chase(T))=m\_Chase(T), we obtain a contradiction which completes the proof of this item.
+3.
+To prove the results, we let $\left(\Sigma_{l}\right)_{l\geq 0}$ be the sequence of m-tables successively generated when running Algorithm 1, and we let $\lambda$ be the least integer such that $\Sigma_{\lambda}=\Sigma_{\lambda+1}=m\_Chase(T)$.
+Since $\Sigma_{0}$ is obtained from $T$ by ‘converting’ tuples in the data warehouse into m-tuples, according to the restrictions on missing values earlier stated, every m-tuple $\sigma$ in $\Sigma_{0}$ satisfies the following:
+  • (A)
+    If $\mathbb{K}\not\subseteq sch(\sigma)$ then $\sigma$ has been obtained from a tuple in a dimension table.
+Thus, there exists $i_{0}\in\{1,\ldots,n\}$ such that $sch(\sigma)\subseteq sch(D_{i_{0}})$, ${\sf tuples}(\sigma(K_{i_{0}}))=(k_{i_{0}})$ and or every $A\in sch^{*}(D_{i_{0}})$, $\sigma(A)\subseteq\{a~|~(\exists q\in D_{i_{0}})(q.K_{i_{0}}=k_{i_{0}}\wedge q.A=a)\}$.
+  • (B)
+    Otherwise, $\sigma$ has been obtained from a tuple in the fact table, and thus, for every $M_{j}\in\mathbb{M}$, $\sigma(M_{j})\subseteq\{m_{j}~|~(\exists t\in F)(t.\mathbb{K}=\sigma(\mathbb{K})\wedge t.M_{j}=m_{j})\}$.
+Moreover, for every $i=1,\ldots,n$, for every $A\in sch^{*}(D_{i})$, we have $\sigma(A)=\emptyset$ (because $A\not\in sch(F)$), implying that $\sigma(A)\subseteq\{a~|~(\exists t\in D_{i})(t.K_{i}=k.K_{i}\wedge t.A=a)\}$.
+We show by induction that, for $l\geq 0$, the m-tuples in $\Sigma_{l}$ satisfy one of the two items (A) or (B) above.
+Indeed, assuming that all m-tuples in $\Sigma_{l}$ satisfy (A) and (B), let $\sigma$ and $\sigma^{\prime}$ be in $\Sigma_{l}$ and $Y\to B$ the dependency in $FD$ used by the m_Chase rule when generating $\Sigma_{l+1}$.
+Thus, $YB\subseteq sch(\sigma)$ and $YB\subseteq sch(\sigma^{\prime})$, and applying the m_Chase rule to $\sigma$ and $\sigma^{\prime}$ modifies $\sigma(B)$ and $\sigma^{\prime}(B)$ to produce $\sigma_{1}$ and $\sigma^{\prime}_{1}$ such that $\sigma_{1}(B)=\sigma^{\prime}_{1}(B)=\sigma(B)\cup\sigma^{\prime}(B)$.
+Since $B\not\in\mathbb{K}$, it follows that $sch(\sigma)\cap\mathbb{K}=sch(\sigma_{1})\cap\mathbb{K}$ and $sch(\sigma^{\prime})\cap\mathbb{K}=sch(\sigma^{\prime}_{1})\cap\mathbb{K}$, thus that $\sigma_{1}$ falls in case (A), respectively in case (B), if and only if $\sigma$ falls in case (A), respectively in case (B), and the same happens for $\sigma^{\prime}$ and $\sigma^{\prime}_{1}$.
+Moreover, since by our induction hypothesis, $\sigma$ and $\sigma^{\prime}$ satisfy the inclusions stated in items (A) and (B), it is easy to see that such is also the case for $\sigma_{1}$ and $\sigma^{\prime}_{1}$.
+Therefore, it turns out that the m-tuples in $\Sigma_{l+1}$ satisfy the inclusions stated in items (A) and (B) above; in particular, such is the case for m_Chase(T)m\_Chase(T).
+We now prove that, when considering m_Chase(T)m\_Chase(T), the inclusions involving $\sigma(A)$ in item (A) or (B) are in fact equalities.
+In case of item (A), let $\sigma\in m\_Chase(T)$ and $i_{0}\in\{1,\ldots,n\}$ such that $K_{i_{0}}\in sch(\sigma)$, $\mathbb{K}\not\subseteq sch(\sigma)$.
+If there exists $A\in sch^{*}(D_{i_{0}})$ and $q\in D_{i_{0}}$ such that $A\in sch(q)$, $\sigma(K_{i_{0}})=q.K_{i_{0}}$ and $q.A\not\in\sigma(A)$, then by construction of m_Chase(T)m\_Chase(T) there must exist $\sigma_{q}\in m\_Chase(T)$ such that $q\in{\sf tuples}(\sigma_{q}(sch(q)))$.
+It follows that $\sigma(A)\neq\sigma_{q}(A)$ whereas $\sigma(K_{i_{0}})=\sigma_{q}(K_{i_{0}})$, thus that applying the m_Chase rule to $\sigma$, $\sigma_{q}$ and $K_{i_{0}}\to A$ in $FD$, changes m_Chase(T)m\_Chase(T).
+In other words this implies that $m\_Chase(m\_Chase(T))\neq m\_Chase(T)$, which is a contradiction.
+Since similar arguments hold in the case where $\sigma$ satisfies item (B), this part of the proof is complete.
+The last step in this proof consists in showing that for every $\sigma$ in m_Chase(T)m\_Chase(T) of kind (a) in the proposition and such that $\sigma(K_{i_{0}})=k_{i_{0}}$, $k_{i_{0}}$ does not occur in $F$.
+Indeed, if such is the case, m_Chase(T)m\_Chase(T) contains $\sigma^{\prime}$ such that $\mathbb{K}\subseteq sch(\sigma^{\prime})$ and $\sigma^{\prime}(K_{i_{0}})=\sigma(K_{i_{0}})=k_{i_{0}}$.
+Since m_Chase(T)m\_Chase(T) is closed under the m_Chase rule, every attribute $A$ in $sch(\sigma)\setminus K_{i_{0}}$ belongs to $D_{i_{0}}^{*}$ and so, as $K_{i_{0}}\to A$ is in $FD$, AA is in $sch(\sigma^{\prime})$ and $\sigma(A)=\sigma^{\prime}(A)$.
+Hence, $\sigma\sqsubseteq\sigma^{\prime}$, and so, applying the m_Chase rule to $\sigma$ and $\sigma^{\prime}$ changes m_Chase(T)m\_Chase(T) by removing $\sigma$, a contradiction to the fact that m_Chase(T)m\_Chase(T) is closed under the m_Chase rule.
+The proof is therefore complete.
+$\Box$
+Appendix B Proof of Lemma 1
+Lemma 1  Let $T$ be a star-table over universe $U$, $S$ a subset of ${\sf True}(\mathcal{T})$ and $t$ a tuple in ${\sf Cons}(\mathcal{T})$. If $S\models FD$ then $S\cup\{t\}\models FD$.
+Proof.
+Since $S\models FD$, every $\sigma$ in the m-table m_Chase(S)m\_Chase(S) is such that, for every $A$ in $sch(\sigma)$, $\sigma(A)=(a)$.
+Thus m_Chase(S)m\_Chase(S) can be seen as a table that we denote by $S^{*}$.
+Moreover, denoting by $S_{t}$ the table $S\cup\{t\}$, we have $m\_Chase(S_{t})=m\_Chase(S^{*}\cup\{t\})$.
+Let $S^{*}_{t}=S^{*}\cup\{t\}$, and let us consider the computation of $m\_Chase(S^{*}_{t})$.
+To this end, given $q_{1}$ and $q_{2}$ in $S^{*}_{t}$ and $X\to A$ in $FD$ such that $q_{1}.X=q_{2}.X$, the only possible cases are as follows:
+- If $q_{1}$ and $q_{2}$ are in $S^{*}_{t}$, then either $q_{1}$ and $q_{2}$ are not defined over $A$ or, they are both defined such that $q_{1}.A=q_{2}.A$.
+In this case m_Chasem\_Chase does not change $S^{*}_{t}$.
+- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are not defined over $A$, then m_Chasem\_Chase does not change $S^{*}_{t}$.
+- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, and $q_{1}$ and $q_{2}$ are defined over $A$.
+Since $t\in{\sf Cons}(\mathcal{T})$, it is not possible that $q_{1}.A\neq t.A$.
+Thus in this case again, m_Chasem\_Chase does not change $S^{*}_{t}$.
+- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is defined over $A$ and $q_{2}$ is not defined over $A$. Then m_Chasem\_Chase changes $t$ into $ta$ where $a=q_{1}.A$.
+- If $q_{1}\in S^{*}_{t}$, $q_{2}=t$, $q_{1}$ is not defined over $A$ and $q_{2}$ is defined over $A$. Then m_Chasem\_Chase changes $q_{1}$ into $qa$ where $a=t.A$.
+Based on the previous cases, we denote by $\Sigma$ the table obtained from $S^{*}_{t}$ by the following transformations:
+$(1)$ For every $q$ in $S^{*}$ and every $X\to A$ in $FD$ such that $q.X=t.X$, $q$ is not defined over $A$ and $t.A=a$, in $\Sigma$, $q.A$ is set to $a$.
+$(2)$ For every $X\to A$ in $FD$ such that $q.X=t.X$, $q.A=a$ and $t$ is not defined over $A$, in $\Sigma$, $t.A$ is set to $a$.
+Since in a star schema, for every attribute $A$ in $U$ there is at most one functional dependency whose right hand-side is $A$, the construction of $\Sigma$ cannot generate conflicts, thus entailing that $\Sigma$ contains no conflicts.
+We now show that $m\_Chase(\Sigma)=\Sigma$.
+Indeed, let $q_{1}$ and $q_{2}$ be in $\Sigma$ and $X\to A$ be in $FD$.
+The only possibility for m_Chasem\_Chase to change $\Sigma$ is that $q_{1}.X=q_{2}.X$, $q_{1}.A=a$ and $q_{2}$ is not defined over $A$.
+As this does not happen in $S^{*}$ (because $m\_Chase(S^{*})=S^{*}$), and as this does not happen in $\Sigma$ (by definition of $\Sigma$), we obtain that $m\_Chase(\Sigma)=\Sigma$.
+It follows that $m\_Chase(S^{*}_{t})=\Sigma$, and since we have seen that no conflicts occur in $\Sigma$, $S^{*}_{t}\models FD$ holds.
+The proof is therefore complete.
+$\Box$
+Appendix C Proof of Proposition 6
+Proposition 6  Let $T$ be a star-table over universe $U$.
+For a given $X\to A$ in $FD$ and a given tuple $x$ over $X$, let $\{xa_{1},\ldots,xa_{q}\}$ be the set of all tuples over $XA$ that belong to ${\sf True}(\mathcal{T})$.
+Then, for every $R$ in ${\sf Rep}(T)$, there exists $q_{0}$ in $\{1,\ldots,q\}$ such that $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R})$.
+Proof.
+It is easy to show that the proposition holds for $q=1$.
+Indeed, in this case $xa_{1}$ is in ${\sf Cons}(\mathcal{T})$, and thus in ${\sf True}(\mathcal{R})$ for every $R$ in ${\sf Rep}(T)$.
+We now assume that $q>1$, which implies that for every $\sigma$ in m_Chase(T)m\_Chase(T) such that $XA\subseteq sch(\sigma)$ and $x$ is in ${\sf tuples}(\sigma(X))$, ${\sf tuples}(\sigma(XA))=(xa_{1},\ldots,xa_{q})$ with $q>1$.
+Moreover, given the form of functional dependencies in $FD$, either $(a)$ $X=\mathbb{K}$ and $A$ is in $\mathbb{M}$, or $(b)$ there exists $i_{0}$ in $\{1,\ldots,n\}$ such that $X=K_{i_{0}}$ and $A$ is in $sch^{*}(D_{i_{0}})$.
+Let $R$ be in ${\sf Rep}(T)$ such that ${\sf True}(\mathcal{R})$ contains no $xa$ from $\{xa_{1},\ldots,xa_{q}\}$, and let $q_{0}$ in $\{1,\ldots,q\}$.
+Denoting by $R^{\prime}$ the table $R\cup\{xa_{q_{0}}\}$, we show that in either case $(a)$ or $(b)$ above, $R^{\prime}\models FD$.
+To this end, we first notice that, since for every $\mathbb{K}$-value $k$ and for any $K_{i}$-value $_{i}, kk$ and $k_{i}$ are in ${\sf Cons}(\mathcal{T})$, the tuple $x$ is in ${\sf True}(\mathcal{R})$.
+Hence, ${\sf True}(\mathcal{R})$ contains tuples whose $X$-value is $x$ and for every such tuple, the $A$-value is missing.
+The cases $(a)$ and $(b)$ are then handled as follows when computing $m\_Chase(R^{\prime})$: every $t$ in ${\sf True}(\mathcal{R}^{\prime})$ such that $t.X=x$ is considered and its $A$-value is set to $a_{q_{0}}$.
+Denoting by $\Sigma$ the resulting m-table, let $X^{\prime}\to A^{\prime}$ be a dependency in $FD$, and let $\sigma_{1}$ and $\sigma_{2}$ be two m-tuples in $\Sigma$ such that $\sigma_{1}(X^{\prime})\cap\sigma_{2}(X^{\prime})\neq\emptyset$ where $X^{\prime}$ is either $\mathbb{K}$ or $K_{j}$.
+Since $R$ is in ${\sf Rep}(T)$, $R$ satisfies $FD$ and thus, all components of all m-tuples in m_Chase(T)m\_Chase(T) are singletons.
+Since this also holds in $\Sigma$, it follows that $\sigma_{1}(X^{\prime})\cap\sigma_{2}(X^{\prime})\neq\emptyset$ implies that $\sigma_{1}(X^{\prime})=\sigma_{2}(X^{\prime})$.
+Then, the following holds:
+- If $X^{\prime}=X$ then by definition of $\Sigma$, $\sigma_{1}(A)=\sigma_{2}(A)=(a_{q_{0}})$. Thus $\Sigma$ is not changed.
+- If $X^{\prime}=K_{j}$, $X=K_{i}$ and $i\neq j$, $\sigma_{1}$ and $\sigma_{2}$ are both in $R^{\prime}$ and their $A^{\prime}$-components have not been changed when building $\Sigma$ (because $K_{j}\neq K_{i}$ implies $A^{\prime}\neq A$).
+Since $m\_Chase({\sf True}(\mathcal{R}^{\prime}))={\sf True}(\mathcal{R}^{\prime})$ and $R^{\prime}\models FD$, either $A^{\prime}$ is in $sch(\sigma_{1})\cap sch(\sigma_{2})$ and $\sigma_{1}(A^{\prime})=\sigma_{2}(A^{\prime})$, or $A^{\prime}$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$.
+In either case, $\Sigma$ is not changed.
+- If $X^{\prime}=K_{j}$ and $X=\mathbb{K}$, since as above we have $A^{\prime}\neq A$, m_Chasem\_Chase does not change $\Sigma$.
+- If $X^{\prime}=\mathbb{K}$ and $X=K_{i}$, then in ${\sf True}(\mathcal{R}^{\prime})$, we have $\sigma_{1}(X)=\sigma_{2}(X)=x$, and $A$ is neither in $sch(\sigma_{1})$ nor in $sch(\sigma_{2})$.
+By definition of $\Sigma$, $\sigma_{1}(A)$ and $\sigma_{2}(A)$ are set to $(a_{q_{0}})$, and then here again, m_Chasem\_Chase does not change $\Sigma$.
+We thus obtain that $\Sigma=m\_Chase(\mathcal{R}^{\prime})$ and that $R^{\prime}\models FD$, because $R\models FD$ and no conflicts occur in $\Sigma$.
+Moreover, as ${\sf True}(\mathcal{R})\subseteq{\sf True}(\mathcal{R}^{\prime})$, we have ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}^{\prime})$.
+It follows from Definition 2 that $R$ can not be in ${\sf Rep}(T)$ since the inclusion ${\sf True}(\mathcal{R})\subset{\sf True}(\mathcal{R}^{\prime})$ is strict (because $xa_{q_{0}}$ is in ${\sf True}(\mathcal{R}^{\prime})$ but not in ${\sf True}(\mathcal{R})$).
+As this is a contradiction with our hypothesis that $R$ is in ${\sf Rep}(T)$, the proof is complete.
+$\Box$
+Appendix D Proof of Proposition 7
 Proposition 7  Let $T$ be a star-table over $U$. $R$ is a repair of $T$ if and only if there is a $\varphi$ as defined above such that ${\sf True}(\mathcal{R})={\sf True}(\mathcal{R}_{\varphi})$.
@@
 Indeed, let $X\to A$ in $FD$, $t_{\varphi}(\sigma)$ and $t_{\varphi}(\sigma^{\prime})$ in $T_{\varphi}$ such that $t_{\varphi}(\sigma).X=t_{\varphi}(\sigma^{\prime}).X$.
-If $\sigma$ and $\sigma^{\prime}$ are defined over $A$, then, by construction of $T_{\varphi}$, we have $t_{\varphi}(\sigma).A=t_{\varphi}(\sigma^{\prime}).A$, in which case $m\_Chase$ does not change $T_{\varphi}$.
-The case whereby $\sigma$ is defined over $A$ and $\sigma^{\prime}$ is not defined over $A$ is not possible in $m\_Chase(T)$, and thus it is not possible that $t_{\varphi}(\sigma)$ is defined over $A$ while $t_{\varphi}(\sigma^{\prime})$ is not.
+If $\sigma$ and $\sigma^{\prime}$ are defined over $A$, then, by construction of $T_{\varphi}$, we have $t_{\varphi}(\sigma).A=t_{\varphi}(\sigma^{\prime}).A$, in which case m_Chasem\_Chase does not change $T_{\varphi}$.
+The case whereby $\sigma$ is defined over $A$ and $\sigma^{\prime}$ is not defined over $A$ is not possible in m_Chase(T)m\_Chase(T), and thus it is not possible that $t_{\varphi}(\sigma)$ is defined over $A$ while $t_{\varphi}(\sigma^{\prime})$ is not.
 Therefore, we have $m\_Chase(T_{\varphi})=T_{\varphi}$, and this implies that $T_{\varphi}\models FD$, because no conflicts can occur in $T_{\varphi}$.
@@
 More formally: $\tau(\Sigma)=\{q\in\mathcal{T}~|~(\exists\sigma\in\Sigma)(\exists t\in{\sf tuples}(\sigma))(q\sqsubseteq t)\}$.
-Recalling from LauS25 that, when $FD$ is acyclic, we also have $m\_Chase({\sf Cons}(\mathcal{T}))={\sf Cons}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})\models FD$.
-We prove by induction that, at each step $k$ of the computation of $m\_Chase(R_{\varphi})$ the obtained m-table $\Sigma^{k}$ is such that $(1)$ $\Sigma^{k}$ contains no conflict and $(2)$ $\tau(\Sigma^{k})=\tau(R_{\varphi})$.
+Recalling from LauS25 that, when $FD$ is acyclic, we also have $m\_Chase({\sf Cons}(\mathcal{T}))={\sf Cons}(\mathcal{T})$ and ${\sf Cons}(\mathcal{T})\models FD$, we prove by induction that, at each step $k$ of the computation of $m\_Chase(R_{\varphi})$ the obtained m-table $\Sigma^{k}$ is such that $(1)$ $\Sigma^{k}$ contains no conflict and $(2)$ $\tau(\Sigma^{k})=\tau(R_{\varphi})$.
 $\bullet$ For $k=0$, i.e., $\Sigma^{0}=R_{\varphi}$, $(2)$ is obvious.
@@
 To show that $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$, we successively consider the following cases:
--- If $q$ and $q^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$): as these sets are ‘closed’ under m_Chase, $t$ and $t^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$), and so $t^{\prime}_{A}$ is in $T_{\varphi}$ (respectively in ${\sf Cons}(\mathcal{T})$), because $t^{\prime}_{A}$ is obtained through the m-Chase process.
+- If $q$ and $q^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$): as these sets are closed under the m_Chase rule, $t$ and $t^{\prime}$ are both in $T_{\varphi}$ (respectively both in ${\sf Cons}(\mathcal{T})$), and so $t^{\prime}_{A}$ is in $T_{\varphi}$ (respectively in ${\sf Cons}(\mathcal{T})$), because $t^{\prime}_{A}$ is obtained through the m_Chase rule.
 - If $q$ is in ${\sf Cons}(\mathcal{T})$ and $q^{\prime}$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$: in this case $xa$ is in ${\sf Cons}(\mathcal{T})$ (because $xa\sqsubseteq q$).
-If $\sigma^{\prime}$ is the m-tuple in $m\_Chase(T)$ such that $q^{\prime}=t_{\varphi}(\sigma^{\prime})$, $XA\subseteq sch(\sigma^{\prime})$ and $xa$ is in ${\sf tuples}(\sigma^{\prime}(XA))$.
+If $\sigma^{\prime}$ is the m-tuple in m_Chase(T)m\_Chase(T) such that $q^{\prime}=t_{\varphi}(\sigma^{\prime})$, $XA\subseteq sch(\sigma^{\prime})$ and $xa$ is in ${\sf tuples}(\sigma^{\prime}(XA))$.
 Hence $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$, and thus $t^{\prime}_{A}$ is in ${\sf tuples}(R_{\varphi})$.
--- If $t$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$ and $t^{\prime}$ in ${\sf Cons}(\mathcal{T})$: in this case, assume first that $xa$ is in ${\sf Cons}(\mathcal{T})$.
+- If $q$ is in $T_{\varphi}\cap{\sf Confl}(\mathcal{T})$ and $q^{\prime}$ in ${\sf Cons}(\mathcal{T})$: in this case, assume first that $xa$ is in ${\sf Cons}(\mathcal{T})$.
 Then, $t^{\prime}_{A}$ is in ${\sf Cons}(\mathcal{T})$, because for every $Y\to B$ in $FD$ other than $X\to A$, if $YB\subseteq sch(t^{\prime}_{A})$ then $YB\subseteq sch(t^{\prime})$.
-Assuming now that $xa$ is in ${\sf Confl}(\mathcal{T})$, we have $a=\varphi(x)$ and so, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime}_{A})$, $t^{\prime}_{A}.B=\varphi(t^{\prime}_{A}.Y)$ (because as $t^{\prime}$ is in ${\sf Cons}(\mathcal{T})$, we also have that for every $Y\to B$ such that $YB\subseteq sch(t^{\prime})$, $t^{\prime}.B=\varphi(t^{\prime}.Y)$).
-Hence, there exists $\sigma^{\prime}$ in $m\_Chase(T)$ such that $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$.
+Assuming now that $xa$ is in ${\sf Confl}(\mathcal{T})$, we have $a=\varphi(x)$ and so, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime}_{A})$, $t^{\prime}_{A}.B=\varphi(t^{\prime}_{A}.Y)$ (because as $t^{\prime}$ is in ${\sf Cons}(\mathcal{T})$, for every $Y\to B$ such that $YB\subseteq sch(t^{\prime})$, $t^{\prime}.B=\varphi(t^{\prime}.Y)$).
+Hence, there exists $\sigma^{\prime}$ in m_Chase(T)m\_Chase(T) such that $t^{\prime}_{A}\sqsubseteq t_{\varphi}(\sigma^{\prime})$.
 Therefore $t^{\prime}_{A}$ is in $\tau(R_{\varphi})$.
@@
 By Definition 2, in order to show that $R_{\varphi}$ is in ${\sf Rep}(T)$, we have to show that ${\sf True}(\mathcal{R}_{\varphi})$ is maximal.
-To this end, let $S$ be such that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{S})\subseteq{\sf True}(\mathcal{T})$ and $S\models FD$, and we consider a tuple $q$ in ${\sf True}(\mathcal{S})\setminus{\sf True}(\mathcal{R}_{\varphi})$.
-Then, as $q$ is in ${\sf True}(\mathcal{T})$, there exist $\sigma$ in $m\_Chase(T)$ and $t$ in ${\sf tuples}(\sigma)$ such that $q\sqsubseteq t$.
+To this end, we let $S$ be such that ${\sf True}(\mathcal{R}_{\varphi})\subseteq{\sf True}(\mathcal{S})\subseteq{\sf True}(\mathcal{T})$ and $S\models FD$, and we let $q$ be a tuple in ${\sf True}(\mathcal{S})\setminus{\sf True}(\mathcal{R}_{\varphi})$.
+As $q$ is in ${\sf True}(\mathcal{T})$, there exist $\sigma$ in m_Chase(T)m\_Chase(T) and $t$ in ${\sf tuples}(\sigma)$ such that $q\sqsubseteq t$.
 Since ${\sf Cons}(\mathcal{T})\subseteq{\sf True}(\mathcal{R}_{\varphi})$, it follows that $q$ is in ${\sf Confl}(\mathcal{T})$, implying that there exists $X\to A$ in $FD$ such that $XA\subseteq sch(q)$ and ${\sf True}(\mathcal{T})$ contains $xa^{\prime}$ such that $x=q.X$ and $a^{\prime}\neq q.A$.
@@
 $\Box$
-Appendix B Proof of Proposition 8
-Proposition 8  Let $T$ be a star-table over universe $U$, and $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ a query, such that $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $\Gamma$ is an independent selection condition.
-Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in $m\_Chase(T)$ such that
+Appendix E Proof of Proposition 9
+Proposition 9  Let $T$ be a star-table over universe $U$, and let $Q:{\tt select}$ $X$ from $T$ where $\Gamma$ be a query such that $\Gamma$ is an independent selection condition.
+Then ${\sf C\_ans}(Q)$ is the set of all tuples $x$ over $X$ for which there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that
   1. 1.
@@
   2. 2.
-    for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$,
-  3. 3.
-    for every $Y\to B$ in $FD$ such that $Y\subseteq X$ and $B\in sch(\Gamma)$, ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
+    for every $Y\to B$ in $FD$ such that $YB\subseteq sch(Q)$:
+    1. (a)
+      if $B\in X$, then $|{\sf tuples}(\sigma(B))|=1$,
+    2. (b)
+      if $B\in sch(\Gamma)$, then ${\sf tuples}(\sigma(B))\subseteq Sat(\Gamma(B))$.
 Proof.
-Assume first that $\sigma$ in $m\_Chase(T)$ satisfies the items of the proposition.
-To show that $x$ is ${\sf C\_ans}(Q)$, we consider a repair $R$ in ${\sf Rep}(T)$ and we show that ${\sf True}(\mathcal{R})$ contains a tuple $x\gamma$ where $\gamma$ is in $Sat(\Gamma)$.
-To this end, we first notice that by Proposition 1, the first two items imply that $x$ is in ${\sf Cons}(\mathcal{T})$ and thus, by Theorem 4.1, $x$ is in ${\sf True}(\mathcal{R})$.
+Assume first that $\sigma$ in m_Chase(T)m\_Chase(T) satisfies the items of the proposition.
+To show that $x=\sigma(X)$ is in ${\sf C\_ans}(Q)$, we consider a repair $R$ in ${\sf Rep}(T)$ and we show that ${\sf True}(\mathcal{R})$ contains a tuple $x\gamma$ where $\gamma$ is in $Sat(\Gamma)$.
+By Proposition 7, we can assume that $R$ has been obtained by applying process (P), which implies that ${\sf True}(\mathcal{R})$ contains a tuple $t_{\sigma}$ in ${\sf tuples}(\sigma)$.
+Moreover, by Proposition 1, item 1 and item 2(b) imply that $|\sigma(X)|=1$, thus $t_{\sigma}.X=x$ and $x$ is in ${\sf Cons}(\mathcal{T})$.
+By Theorem 4.1, it follows that $x$ is in ${\sf True}(\mathcal{R})$.
 Moreover, since $\Gamma$ is independent, item 1 implies that for every $B$ in $sch(\Gamma)$, there exits $b$ in $\sigma(B)$ such that $b$ is in $Sat(\Gamma(B))$.
-If $S_{1}$ is the set of all $B$ in $sch(\Gamma)$ such that there exists $Y\to B$ in $FD$ where $YB\subseteq sch(Q)$, let $S_{2}=sch(\Gamma)\setminus S_{1}$ and $\gamma_{2}$ a tuple in ${\sf tuples}(\sigma(S_{2}))$ such that for every $B$ in $S_{2}$, $\gamma_{2}.B\in Sat(\Gamma(B))$ (such a tuple exists because by item 1, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$).
-We show that $x\gamma_{2}$, which belongs to ${\sf tuples}(\sigma(XS_{2}))$, is in ${\sf True}(\mathcal{R})$.
-Indeed, by definition of $S_{2}$, if $Y\to B$ is in $FD$ and such that $YB\subseteq XS_{2}$, then $B\not\in S_{2}$.
-Thus, as $\mathbf{K}\cap S_{2}=\emptyset$ (because $\mathbf{K}\cap sch(\Gamma)=\emptyset$ and $S_{2}\subseteq sch(\Gamma)$), $YB\subseteq XS_{2}$ holds if and only if $YB\subseteq X$.
-Consequently $x\gamma_{2}$ is in ${\sf Cons}(\mathcal{T})$ because so is $x$, showing by Theorem 4.1 that $x\gamma_{2}$ belongs to ${\sf True}(\mathcal{R})$.
-Now let $B$ be in $S_{1}$ and $Y\to B$ the corresponding dependency in $FD$ such that $Y\subseteq X$.
-By Proposition 6, $R$ can be defined using process (P) by which a tuple $t_{\varphi}(\sigma)$ from ${\sf tuples}(\sigma)$ is set to be in ${\sf True}(\mathcal{R})$.
-Since $sch(t_{\varphi}(\sigma))=sch(\sigma)$ and $sch(Q)\subseteq sch(\sigma)$, we write $t_{\varphi}(\sigma)$ as $x_{\sigma}\gamma_{\sigma}q$ where $x_{\sigma}=t_{\varphi}(\sigma).X$, $\gamma_{\sigma}=t_{\varphi}(\sigma).sch(\Gamma)$, and $q=t_{\varphi}(\sigma).(sch(\sigma)\setminus sch(Q))$.
-In this setting, we have $t_{\varphi}(\sigma).Y=t.Y$ (because, by Proposition 2(1), $\sigma(Y)$ contains one tuple), and thus, by the Lossless-join rule (which applies here because $R\models FD$) applied to $t_{\varphi}(\sigma)$, $x\gamma_{2}$ and $Y\to B$, we obtain that $x\gamma_{2}b$ is in ${\sf True}(\mathcal{R})$.
-Reproducing this reasoning for every $B$ in $S_{1}$, we obtain a tuple $x\gamma_{2}\gamma_{1}$ in $sch(Q)$ such that for every $B$ in $S_{1}$, $x\gamma_{2}\gamma_{1}.B$ is in $\sigma(B)$ and thus, by item 3, in $Sat(\Gamma(B))$.
-It follows that $x\gamma_{1}\gamma_{2}$ is in ${\sf True}(\mathcal{R})$ and that $\gamma_{1}\gamma_{2}$ is in $Sat(\Gamma)$.
-Thus $x$ is in ${\sf Ans}(Q^{[R]})$, and since this reasoning holds for every repair $R$ of $T$, we obtain that $x$ is in ${\sf C\_ans}(Q)$.
+Depending on whether $B$ occurs as a right hand-side of a dependency in $FD$ or not, we have the following:
+(1) If $B$ is the right hand-side of no dependency in $FD$ (which happens if $B$ is a key attribute in $\mathbb{K}$), then we have $|\sigma(B)|=1$ and thus $\sigma(B)=(b)$, implying that $t_{\sigma}.B$ is in $Sat(\Gamma(B))$.
+(2) Otherwise, if $B$ is the right hand-side of a dependency in $FD$, by item 2(b), $t_{\sigma}.B$ is obviously in $Sat(\Gamma(B))$.
+It follows that in any case, $t_{\sigma}.B$ is in $Sat(\Gamma(B))$, thus that $t_{\sigma}.sch(\Gamma)$ is in $Sat(\Gamma)$ since $\Gamma$ is independent.
+Hence, for $\gamma=t_{\sigma}.sch(\Gamma)$, $x\gamma$ is in ${\sf True}(\mathcal{R})$, showing that $x$ is in ${\sf C\_ans}(Q)$.
 Conversely, let $x$ be in ${\sf C\_ans}(Q)$.
 By Proposition 8, $x$ belongs to ${\sf C\_ans}^{+}(Q)$, meaning that there exists $t$ in ${\sf True}(\mathcal{T})$ such that $sch(Q)\subseteq sch(t)$, $t.X=x$, $x\in{\sf Cons}(\mathcal{T})$ and $t.sch(\Gamma)\in Sat(\Gamma)$.
-As for every $t\in{\sf True}(\mathcal{T})$, there exists $\sigma$ in $m\_Chase(T)$ such that $t$ is in ${\sf tuples}(\sigma(sch(t)))$, there must exist $\sigma$ in $m\_Chase(T)$ such that $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ and for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$ (because $x$ is in ${\sf Cons}(\mathcal{T})$).
-In other words, there exists $\sigma$ in $m\_Chase(T)$ satisfying the first two items of the proposition.
-Denoting by $\Sigma(Q)$ the set of all these m-tuples, let $\sigma$ be in $\Sigma(Q)$ for which item 3 is not satisfied.
-Then, let $Y\to B$ in $FD$ such that $Y\subseteq X$, $B\in sch(\Gamma)$ and $b$ in ${\sf tuples}(\sigma(B))\setminus Sat(\Gamma(B))$.
+As for every $t$ in ${\sf True}(\mathcal{T})$, there exists $\sigma$ in m_Chase(T)m\_Chase(T) such that $t$ is in ${\sf tuples}(\sigma(sch(t)))$, there must exist $\sigma$ in m_Chase(T)m\_Chase(T) such that $sch(Q)\subseteq sch(\sigma)$, $x\in\sigma(X)$, $\sigma(sch(\Gamma))\cap Sat(\Gamma)\neq\emptyset$ and for every $Y\to B$ in $FD$ such that $YB\subseteq X$, $|{\sf tuples}(\sigma(B))|=1$ (because $x$ is in ${\sf Cons}(\mathcal{T})$).
+In other words, there exists $\sigma$ in m_Chase(T)m\_Chase(T) satisfying tem 1 and item 2(a) of the proposition.
+Denoting by $\Sigma(Q)$ the set of all these m-tuples, let $\sigma$ be in $\Sigma(Q)$ for which item 2(b) is not satisfied.
+Then, let $Y\to B$ in $FD$ such that $B\in sch(\Gamma)$ and $b$ in ${\sf tuples}(\sigma(B))\setminus Sat(\Gamma(B))$.
 By item 1, ${\sf tuples}(\sigma(sch(Q)))$ contains $t$ such that $t.X=x$ and $t.B=b$.
@@
 Consequently, $t$ and $x\gamma$ can not both belong to ${\sf True}(\mathcal{R})$ (because $R\models FD$).
-Therefore, if $\Sigma(Q)$ contains one m-tuple not satisfying item 3, then $x$ cannot belong to ${\sf C\_ans}(Q)$.
-We moreover notice that if one $\sigma$ in $\Sigma(Q)$ does not satisfy item 3 then for any other $\sigma^{\prime}$ in $\Sigma(Q)$, we have $\sigma^{\prime}(Y)=\sigma(Y)$ and thus $\sigma^{\prime}(B)=\sigma(B)$.
-It thus follows that $\sigma^{\prime}$ does not satisfy item 3 either.
-As a consequence, we obtain that if $x$ is in ${\sf C\_ans}(Q)$ then $m\_Chase(T)$ contains an m-tuple satisfying all three items in the proposition.
+Therefore, if $\Sigma(Q)$ contains one m-tuple not satisfying item 2(b), then $x$ cannot belong to ${\sf C\_ans}(Q)$.
+We moreover notice that if one $\sigma$ in $\Sigma(Q)$ does not satisfy item 2(b) then for any other $\sigma^{\prime}$ in $\Sigma(Q)$, we have $\sigma^{\prime}(Y)=\sigma(Y)$ and thus $\sigma^{\prime}(B)=\sigma(B)$.
+It thus follows that $\sigma^{\prime}$ does not satisfy item 2(b) either.
+As a consequence, we obtain that if $x$ is in ${\sf C\_ans}(Q)$ then m_Chase(T)m\_Chase(T) contains an m-tuple satisfying all three items in the proposition.
 The proof is therefore complete.
 $\Box$
-Appendix C Proof of Proposition 10
+Appendix F Proof of Proposition 10
+[…]
+  1. 1.
+    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{1}))(t.sch(\Gamma)\not\in Sat(\Gamma))$,
+  2. 2.
+    $(\forall\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ}))(\exists t\in{\sf tuples}(\sigma)\cap{\sf True}(\mathcal{R}_{2}))(t.sch(\Gamma)\in Sat(\Gamma))$.
+Proof.
+The proof relies on the fact that for every repair $R$ of $T$ and every $\sigma$ in $\Sigma(\mathcal{AQ})$, ${\sf True}(\mathcal{R})$ contains exactly one tuple $t$ such that $t\in{\sf tuples}(\sigma)$ and $t.K\in\sigma(\mathbb{K})$ (the existence is a consequence of Proposition 5, unicity follows from Proposition 2(1) because $|\sigma(\mathbb{K})|=1$, $\mathbb{K}$ is a key of $R$ and $R\models FD$).
+As a consequence of Proposition 7, the existence of repairs $R_{1}$ and $R_{2}$ is explicitly shown using the process (P).
+1. Regarding $R_{1}$, we build $\varphi$ as follows: For every $\sigma$ in $\Sigma(\mathcal{AQ})$:
+Step P1. For every $A_{p}^{k}$ in $sch^{*}(D_{p})\cap sch(\sigma)$ such that $\sigma(K_{p})=k_{p}$ and $\varphi_{p}^{k}$ has not been assigned:
+$(a)$ If $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then choose an $A_{p}^{k}$-value in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
+$(b)$ If $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ and ${\sf tuples}(\sigma(A_{p}^{k}))\not\subseteq Sat(\Gamma(A_{k}^{p}))$, choose an $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})\setminus Sat(\Gamma(A_{k}^{p}))$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
+If ${\sf tuples}(\sigma(A_{p}^{k}))\subseteq Sat(\Gamma(A_{k}^{p}))$, choose any $A_{p}^{k}$-value $a^{k}_{p}$ in $\sigma(K_{p}A_{p}^{k})$ and let $\varphi_{p}^{k}(k_{p})=a_{p}^{k}$.
+Step P2. If $\mathbb{K}\subseteq sch(\sigma)$, for every $M_{i}$ in $\mathbb{M}\cap sch(\sigma)$, if $\varphi_{i}(k)$ (where $k=\sigma(\mathbb{K})$) has not been assigned, then
+$(a)$ If $M_{i}\not\in sch(\Gamma)$ or if $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma^{+}(\mathcal{AQ})$, choose an $M_{i}$-value $m$ in $\sigma(M_{i})$ and let $\varphi_{i}(k)=m$.
+$(b)$ Otherwise (i.e., $M_{i}\in sch(\Gamma)$ and $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$), choose an $M_{i}$-value $m$ in $\sigma(M_{i})\setminus Sat(\Gamma(M_{i}))$ for $\varphi_{i}(k)$.
+Once all m-tuples in $\Sigma(\mathcal{AQ})$ have been considered, the two steps above are completed by considering all non-processed $K_{i}$- or $\mathbb{K}$-values as done in the generic description of process (P).
+Consequently, the corresponding set $\varphi$ in process (P) is properly defined.
+Step P3.
+Build up the tuples $t_{\varphi}(\sigma)$ for every $\sigma$ in m_Chase(T)m\_Chase(T).
+We notice that, according to the definitions of $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$, the following holds:
+- if $\sigma\in\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\in Sat(\Gamma)$,
+- if $\sigma\in\Sigma(\mathcal{AQ})\setminus\Sigma^{+}(\mathcal{AQ})$ then $t_{\varphi}(\sigma).sch(\Gamma)\not\in Sat(\Gamma)$.
+By Proposition 7, we therefore obtain a repair $R_{1}$ satisfying item 1 in the proposition.
+2.
+Regarding the item 2, the repair $R_{2}$ is obtained as above, by changing $(b)$ in Step P1 and Step P2 as follows: instead of choosing a value not in the corresponding set $Sat(\Gamma(A))$, choose a value that belongs to $Sat(\Gamma(A))$.
+We notice that, according to the definitions of $\Sigma(\mathcal{AQ})$ and $\Sigma^{+}(\mathcal{AQ})$, such a choice is always possible.
+Then, as above, these two steps are completed by considering all non-processed $K_{i}$- or $\mathbb{K}$-values, and Step P3 is considered.
+In this case, by Proposition 7, we obtain a repair $R_{2}$ such that for every $\sigma$ in $\Sigma(\mathcal{AQ})$, $t_{\varphi}(\sigma).sch(\Gamma)$ is in $Sat(\Gamma)$.
+Hence $R_{2}$ satisfies item 2 in the proposition, and the proof is complete.
+$\Box$
+Appendix G Proof of Proposition 11
 Proposition 11  Let $T$ be a star-table over universe $U$ and $\mathcal{AQ}:$~select $aggr(M_{i})$ from $T$ where $\Gamma$ an analytic query with no group-by clause.
-  • If $aggr$ is $min$, $max$ or $count$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where $min\_ans$ and $max\_ans$ are returned by Algorithm 3.
-  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then $min\_ans$ and $max\_ans$ as returned by Algorithm 3 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
+  • If aggraggr is minmin, maxmax or countcount, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$ where min_ansmin\_ans and max_ansmax\_ans are returned by Algorithm 4.
+  • If $aggr=sum$ and if ${\sf C\_ans}(\mathcal{AQ})=[glb,lub]$, then min_ansmin\_ans and max_ansmax\_ans as returned by Algorithm 4 satisfy that $min\_ans\leq glb$ and $max\_ans\geq lub$.
     Moreover, if for every $m\in adom(M_{i})$, $m\geq 0$, then ${\sf C\_ans}(\mathcal{AQ})=[min\_ans,max\_ans]$.
-  • For every aggregate function and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 3 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
+  • For every aggregate operator and every selection condition $\Gamma$, $ans^{*}$ as returned by Algorithm 4 is equal to ${\sf C\_ans}^{*}(\mathcal{AQ})$.
 Proof.
 We separately consider two cases, depending on whether the test line 2 in Algorithm 4 succeeds or not.
-If the test fails, i.e., if $change\_min\_max$ as returned by the call of Compute_aggregate has value false, this means that $m\_Chase(T)$ contains no $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
-By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is empty.
+If the test fails, i.e., if change_min_maxchange\_min\_max as returned by the call of Compute_aggregate has value false, this means that m_Chase(T)m\_Chase(T) contains no $\sigma$ such that $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$, $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
+By Corollary 1, this holds if and only if the consistent answer to Q:Q:~select $\mathbb{K},X$ from $T$ where $\Gamma$ is empty.
 Thus, the test fails if and only if there exists a repair $R$ of $T$ for which ${\sf Ans}(Q^{[R]})=\emptyset$.
-In this case ${\sf C\_ans}(\mathcal{AQ})$ is NULL, which is expressed by the fact that the values of $min\_ans$, $max\_ans$ and $ans^{*}$ returned by the call of Compute_aggregate are respectively -dummy, +dummy and -dummy.
+In this case ${\sf C\_ans}(\mathcal{AQ})$ is NULL, which is expressed by the fact that the values of min_ansmin\_ans, max_ansmax\_ans and $ans^{*}$ returned by the call of Compute_aggregate are respectively -dummy, +dummy and -dummy.
 Hence, in this case, Algorithm 4 provides the correct answer.
-Suppose now that the test line 2 succeeds, i.e., that the value of $change\_min\_max$ returned by the call of Compute_aggregate is true.
-The statement line 16 of Algorithm 2 shows that $m\_Chase(T)$ contains at least one $\sigma$ such that $\mathbf{K}\cup sch(Q)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
-By Corollary 1, this holds if and only if the consistent answer to $Q:~{}$select $K,X$ from $T$ where $\Gamma$ is not empty.
-Thus, the test succeeds if and only if for every repair $R$ of $T$, ${\sf Ans}(Q^{[R]})\neq\emptyset$, in which case, $min\_ans$, $max\_ans$ and $ans^{*}$ are proper values, either values from $adom(M_{i})$ if $aggr\neq count$, or positive integers if $aggr=count$.
-In this case, it can be seen that the m-tuples $\sigma$ in $m\_Chase(T)$ that contribute to the construction of the interval are such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$.
+Suppose now that the test line 2 succeeds, i.e., that the value of change_min_maxchange\_min\_max returned by the call of Compute_aggregate is true.
+The statement line 14 of Algorithm 3 shows that m_Chase(T)m\_Chase(T) contains at least one $\sigma$ such that $\mathbb{K}\cup sch(Q)\subseteq sch(\sigma)$, $|{\sf tuples}(\sigma(X))|=1$ and ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$.
+By Corollary 1, this holds if and only if the consistent answer to Q:Q:~select $\mathbb{K},X$ from $T$ where $\Gamma$ is not empty.
+Thus, the test succeeds if and only if for every repair $R$ of $T$, ${\sf Ans}(Q^{[R]})\neq\emptyset$, in which case, min_ansmin\_ans, max_ansmax\_ans and $ans^{*}$ are proper values, either values from $adom(M_{i})$ if $aggr\neq count$, or positive integers if $aggr=count$.
+In this case, it can be seen that the m-tuples $\sigma$ in m_Chase(T)m\_Chase(T) that contribute to the construction of the interval are such that $\mathbb{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$.
 This is so, because of our assumption on missing values in the fact table $F$ (every tuple with an $M_{i}$-value must be defined over all key attributes); and because any $\sigma$ not defined over some attributes in $X\cup sch(\Gamma)$ or such that ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)=\emptyset$ cannot contribute in the consistent answer to $\mathcal{AQ}$.
 This explains why, when the test line 7 of Algorithm 3 fails, no action is taken.
-On the other hand, the m-tuples $\sigma$ such that $\mathbf{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ are precisely those in $\Sigma(\mathcal{AQ})$, and recalling that $\Sigma^{+}(\mathcal{AQ})$ is the set of m-tuples in $\Sigma(\mathcal{AQ})$ such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$, we have the following for every $\sigma$ in $\Sigma(\mathcal{AQ})$:
+On the other hand, the m-tuples $\sigma$ such that $\mathbb{K}\cup X\cup sch(\Gamma)\subseteq sch(\sigma)$ and ${\sf tuples}(\sigma(sch(\Gamma)))\cap Sat(\Gamma)\neq\emptyset$ are precisely those in $\Sigma(\mathcal{AQ})$, and recalling that $\Sigma^{+}(\mathcal{AQ})$ is the set of m-tuples in $\Sigma(\mathcal{AQ})$ such that ${\sf tuples}(\sigma(sch(\Gamma)))\subseteq Sat(\Gamma)$, we have the following for every $\sigma$ in $\Sigma(\mathcal{AQ})$:
 $(1)$ By Proposition 5, for every $t\in{\sf tuples}(\sigma)$, there exists a repair $R$ such that $t\in{\sf True}(\mathcal{R})$.
-Moreover, $t$ is the unique tuple over $sch(\sigma)$ in ${\sf True}(\mathcal{R})$ having $t.K$ as a $\mathbf{K}$-value.
+Moreover, $t$ is the unique tuple over $sch(\sigma)$ in ${\sf True}(\mathcal{R})$ having $t.\mathbb{K}$ as a $\mathbb{K}$-value.
 Notice that since $t.sch(\Gamma)$ may not be in $Sat(\Gamma)$, it is possible that $t.M_{i}$ does not contribute in the computation of the aggregate in $R$.
@@
 In this case, since $t.sch(\Gamma)$ is in $Sat(\Gamma)$, $t.M_{i}$ does contribute in the computation of the aggregate in $R$.
-Before showing that the values returned by Algorithm 3 are as stated in Definition 5, we mention that the aggregate operators $min$, $max$ and $sum$ are not defined when their argument is empty, which we write as $aggr(\emptyset)={\tt NULL}$.
-Otherwise, if $v_{1}$, $v_{2}$ and $v_{3}$ are values to which $aggr$ applies, then
+Before showing that the values returned by Algorithm 4 are as stated in Definition 5, we mention that the aggregate operators minmin, maxmax and sumsum are not defined when their argument is empty, which we write as $aggr(\emptyset)={\tt NULL}$.
+Otherwise, if $v_{1}$, $v_{2}$ and $v_{3}$ are values to which aggraggr applies, then
 Commutativity: $aggr(\{v_{1},v_{2}\})=aggr(\{v_{2},v_{1}\})$.
@@
 Moreover, the third property shows that, if $aggr\neq count$, the higher the values, the higher the aggregate values.
-In our context, recalling that elementary values are values in $\sigma(M_{i})$ for $\sigma$ in $\Sigma(\mathcal{AQ})$, this last property shows that, when the aggregate is different than $count$, for a fixed $\sigma$, the least, respectively the highest, aggregate value is obtained by considering the least, respectively the highest, possible $M_{i}$-value.
-These values, respectively denoted by $min_{\sigma}$ and $max_{\sigma}$ are computed lines 8-14 of Algorithm 2.
+In our context, recalling that elementary values are values in $\sigma(M_{i})$ for $\sigma$ in $\Sigma(\mathcal{AQ})$, this last property shows that, when the aggregate is different than countcount, for a fixed $\sigma$, the least, respectively the highest, aggregate value is obtained by considering the least, respectively the highest, possible $M_{i}$-value.
+These values, respectively denoted by $min_{\sigma}$ and $max_{\sigma}$ are computed lines 8-12 of Algorithm 3.
 Since the second property shows that aggregates can be seen as operating over a set of values we recall the following standard properties that will be used in this proof.
-Let $S_{1}$ and $S_{2}$ be sets of values to which $aggr$ applies, and such that $S_{1}\subseteq S_{2}$, then: $\min(S_{1})\geq\min(S_{2})$, $\max(S_{1})\leq\max(S_{2})$, $count(S_{1})\geq count(S_{2})$, and if all values in $S_{1}$ or $S_{2}$ are positive, $sum(S_{1})\leq sum(S_{2})$.
+Let $S_{1}$ and $S_{2}$ be sets of values to which aggraggr applies, and such that $S_{1}\subseteq S_{2}$, then: $\min(S_{1})\geq\min(S_{2})$, $\max(S_{1})\leq\max(S_{2})$, $count(S_{1})\geq count(S_{2})$, and if all values in $S_{1}$ or $S_{2}$ are positive, $sum(S_{1})\leq sum(S_{2})$.
 We notice that regarding the last property, if values in $S_{1}$ or $S_{2}$ can be positive or negative, no generic comparison can be stated.
@@
 First, as for every $\sigma\in\Sigma^{+}(\mathcal{AQ})$, every repair contains exactly one tuple in ${\sf tuples}(\sigma)$, the $M_{i}$-values in $\sigma(M_{i})$ if $M_{i}\not\in sch(\Gamma)$, or in $\sigma(M_{i})\cap Sat(sch(\Gamma))$ otherwise, contribute in the computation of the aggregate.
-Moreover, since by monotonicity, $min_{\sigma}$ contributes in the computation of $min\_ans$ and $max_{\sigma}$ contributes in the computation of $max\_ans$, we have the following, given that $\Sigma^{+}(\mathcal{AQ})\subseteq\Sigma(\mathcal{AQ})$:
+Moreover, since by monotonicity, $min_{\sigma}$ contributes in the computation of min_ansmin\_ans and $max_{\sigma}$ contributes in the computation of max_ansmax\_ans, we have the following, given that $\Sigma^{+}(\mathcal{AQ})\subseteq\Sigma(\mathcal{AQ})$:
   • For $aggr=min$, $min\_ans=\min\{min_{\sigma}~|~\sigma\in\Sigma(\mathcal{AQ})\}$ and $max\_ans=\min\{max_{\sigma}~|~\sigma\in\Sigma^{+}(\mathcal{AQ})\}$.
-These values are computed respectively in lines 23 and 17 of Algorithm 2.
+These values are computed respectively in lines 23 and 18 of Algorithm 3.
   • For $aggr=max$, $min\_ans=\max\{min_{\sigma}~|~\sigma\in\Sigma^{+}(\mathcal{AQ})\}$ and $max\_ans=\max\{max_{\sigma}~|~\sigma\in\Sigma(\mathcal{AQ})\}$.
-These values are computed respectively in lines 17 and 25 of Algorithm 2.
-  • For $aggr=count$, $min\_ans$, respectively $max\_ans$ is the cardinality of $\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$.
-These values are computed respectively in lines 17 and 27 of Algorithm 2.
-  • For $aggr=sum$, $min\_ans$, respectively $max\_ans$, is the minimal, respectively maximal, sum that can be obtained by adding one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$ and then possibly one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$.
-These values are computed in lines 17 and 30, respectively in lines 18 and 32 of Algorithm 2.
-Notice that if all $M_{i}$-values are positive, knowing that adding 00 is neutral, then the test line 29 always fails and thus, $proc\_min$ is the sum of all minimal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$.
-Similarly, in this case, the test line 31 always succeeds, except for 00, and thus, $proc\_max$ is the sum of all maximal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma(\mathcal{AQ})$.
-To complete the proof that the returned values $min\_ans$ and $max\_ans$ are as stated in the proposition, we have to show that ${\sf Rep}(T)$ contains repairs $R_{min}$ and $R_{max}$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{min}]})=min\_ans$ and ${\sf Ans}(\mathcal{AQ}^{[R_{max}]})=max\_ans$.
-These results are consequences of Proposition 9 by considering $R_{min}=R_{2}$ and $R_{max}=R_{1}$ if $aggr=min$, and $R_{min}=R_{1}$ and $R_{max}=R_{2}$ if $aggr$ is $max$ or $count$ or if $M_{i}$-values are positive and $aggr=sum$.
-We also mention that if $aggr=sum$ with no restriction on the $M_{i}$-values, $min\_ans$, respectively $max\_ans$, is the least, respectively the largest, sum that can be obtained using all relevant sets $\sigma(M_{i})$.
+These values are computed respectively in lines 18 and 25 of Algorithm 3.
+  • For $aggr=count$, min_ansmin\_ans, respectively max_ansmax\_ans is the cardinality of $\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$.
+These values are computed respectively in lines 18 and 27 of Algorithm 3.
+  • For $aggr=sum$, min_ansmin\_ans, respectively max_ansmax\_ans, is the minimal, respectively maximal, sum that can be obtained by adding one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$ and then possibly one $M_{i}$-value in every $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$, respectively of $\Sigma(\mathcal{AQ})$.
+These values are computed in lines 18 and 30, respectively in lines 18 and 32 of Algorithm 3.
+Notice that if all $M_{i}$-values are positive, knowing that adding 0 is neutral, then the test line 29 always fails and thus, proc_minproc\_min is the sum of all minimal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma^{+}(\mathcal{AQ})$.
+Similarly, in this case, the test line 31 always succeeds, except for 0, and thus, proc_maxproc\_max is the sum of all maximal $M_{i}$-values in $\sigma(M_{i})$ for $\sigma\in\Sigma(\mathcal{AQ})$.
+To complete the proof that the returned values min_ansmin\_ans and max_ansmax\_ans are as stated in the proposition, we have to show that ${\sf Rep}(T)$ contains repairs $R_{min}$ and $R_{max}$ such that ${\sf Ans}(\mathcal{AQ}^{[R_{min}]})=min\_ans$ and ${\sf Ans}(\mathcal{AQ}^{[R_{max}]})=max\_ans$.
+These results are consequences of Proposition 10 by considering $R_{min}=R_{2}$ and $R_{max}=R_{1}$ if $aggr=min$, and $R_{min}=R_{1}$ and $R_{max}=R_{2}$ if aggraggr is maxmax or countcount or if $M_{i}$-values are positive and $aggr=sum$.
+We also mention that if $aggr=sum$ with no restriction on the $M_{i}$-values, min_ansmin\_ans, respectively max_ansmax\_ans, is the least, respectively the largest, sum that can be obtained using all relevant sets $\sigma(M_{i})$.
 We thus have $min\_ans\leq glb$ and $max\_ans\geq lub$, which concludes this part of the proof.